A Bibliography of Publications about Typographic Fonts

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

03 May 2025 Version 1.130

Title word cross-reference

#1 [Ano89a]. #2 [Ano89b]. #6 [Ano89c].

{ [Hoe94, Hoe94]. 5 × 7 [MBG77]. Ω [Har96, HP98].

-the [PBC91].

1 [Mal95]. 1-2-3 [Ano91a]. 1.3— [Ber96b]. 100-font [BN94]. 1000C [Sta87]. 106 [FN85a, FN85b]. 11#4 [Mit92]. 11th [IEE92]. 13 [MAYB02]. 1986 [Van86]. 1988 [ICC88]. 1989 [AH89]. 1990 [Fur90]. 1990’s [Rie89]. 1991 [MA91]. 1992 [DC93]. 1993 [SK93]. 1995 [AOM95]. 1st [Ano93a].

2 [Orr94, Tur00a, Tur03]. 2-D [Orr94]. 2.1.7 [The03]. 202 [CKT80]. 208pp [Tho19]. 24-Wire [BM87]. 2nd [Ano94a].

3 [Orr94, Tho19]. 3-D [Orr94]. 300-dpi [Mac91a]. 3800 [BS84, FLS78]. 3D [Emm95, SFO88].

4.2 [Ano93f].

5.25in [Bec91b].

6670 [Int82].

7th [JN93].

8-bit [Har92]. 855 [Pet86]. ’86 [CVP86]. 8700/9700 [Xer82a, Xer82b]. 8th [BDV95].

’93 [SK93]. ’95

[Ano95a, AOM95, BDV95, Ano95a]. 9541 [SBSG89]. 978 [Tho19]. 978-1-61689-572-3 [Tho19].

a2ac— [Ols95]. abacus [MVM09b]. abcdefg [Ada89]. ABF [Pos89a]. Abjad [MVM09a]. Abstracts [BWT85]. accented [Zla91]. Access [PRB87]. According [Ano86b]. accounting [Ano94s]. accuracy [AIDG95]. ACM [Rie89]. Acorn [RSSW89]. Acrobat [Ano93b]. across [Van93a, Van93b]. act [Ano93b]. Adaptation [BBG85]. Adaptive [CW85, Ano96c, Coh93]. Addressable [Ano86a]. Addressing [Ano86d]. adjunct [MW92]. Adjust [Ano85e]. Administrator [MO92, Pea92]. Adobe [Ado92, Goo95, Pos89a, Pos89b, Pos90, Pos92a, Pos92b, Ano92a, Ano92b, Ano93b, Ano94t, Dea92c, Har93, Kay96, Pon89, Rig14]. adorns [Bur18]. Advanced [Her81, RS81, Ano91a, Flo89]. Advances [Grä93, Pie91]. Aesthetics [Gai85, BS81, Ben93a, Tay91]. Affair [Dak90]. Africa [Kna93]. against [Flo94]. Age [Ber02, BS81]. Aided [Car84, LS84, Sch98b]. al [CP91b]. Aldus [Ano85f]. algebraic [PMZS97]. Algorithm [BDN83, HWY92, HYH89, HYH91, IO93, MP94, SR02, SZ95]. Algorithmic [DD15]. Algorithms [OO89]. Aliasing [NN89]. Alice [Gar84]. All-Points-Addressable [Ano86a]. aller [Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00, Fau04a, Fau04b]. allocation [Mil81]. ALLText [Ben93b]. Alphabet [HP98, MVM09a, Zap07a, Zap07b]. Alphabete [Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00, Fau04a, Fau04b]. Alphabetgeschichten [Zap07c]. Alphabetic [DMK69]. alphabetics [Glu69]. Alphabets [ST85a, ST85b, Sol83, Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00, Fau04a, Fau04b, Kel11]. Also [Sha94]. Alternate [TX92]. alternative [Har93]. am [SV73]. ambitious [Van93a, Van93b]. America [Ano97a]. American [Ano94d, Ano97a, Ara95]. AMSFonts [Raf93]. Analysis [CBE74, VK93, WCW82, BDV95, MVM09b, SO94]. Analyzing [SFO88]. Andrew [Gla04]. Anecdotes [TBLJ90]. animated [Emm95]. animation [Van93a, Van93b]. Anniversary [Rig14]. annual [Rie89, Mor93]. anomalies [Mid95]. Answers [Ano91c]. Anti [NN89]. Anti-Aliasing [NN89]. Antipolis [Ano87d]. Antiqua [SA70]. Antoine [Ano85g]. Antwerp [Ano85h]. Any [KPB87a, KPB87b]. Anything [Per88]. ao [Jon95]. API [The03, Orr94]. APL} [HK87, GSMN93, How86]. APL2TEX [GSMN93]. ApOFIS [ZI95]. appearance [Bro91]. Apple [Ano93c, Ano93d, Ano93s, Ano94b, BH91, Fox16, Lin91, Sha94, Wag85]. Application [CBE74, KNT^{+}01, cYL88, MCL92]. Applications [Che23, Cla90, Ara95, MCL92, SK93]. Applied [CKM86, CK86]. Approach [CCH88, DL91, Due93b, Dur93b, KM83, LZ86, NNGP89, Ram87, Sch00, Yi92, Zap85, BS96, Krt94, Sha03]. Approaches [Har98, MRT89]. appropriateness [Sha07]. Apr [Ano94a, JN93]. April [Van86]. Arabic [MVM09b, Adi90, All94, AM86, Ano96b, EEN90, Har98, Mac90, MVM09a, Niz90, SB92, Str93a, TA82]. Arabic/English [Adi90]. Arbitrary [MN77]. Architectural [Tho19]. Architecture [SFO88, Ano96c, MCL92, MNSJ99, NNMS94]. archived [PB03]. Area [WH86]. Arena [Ano93b]. arithmetic [MVM09b]. Army [Van92a]. array [Emm95]. arrays [War80]. Arrighi [Mac35]. ART [Dak90, Bix92, Got85, Jan90, Van93a, Van93b, Ano97a]. artifact [AOM95]. Artist [Wri83]. Arts [Car10, Pre57]. ASCII.sty [RNR94]. ascp0.8.2 [Ano98]. Asian [TL92b]. Aspect [Gai85, Som90]. ASQA [Ano96c]. assembly [Hol87, Pet86]. assessing [UB95]. Assessment [AS89, Gib90]. Assumptions [BM87]. Atari [Min83]. Atlanta [Ara95]. Atlante [TP90a]. Atlas [Ham90, TP90b, TP91]. ATM [Ano93u, Ano94c, Kay96]. Attribute [RN83, Lin95]. ATU [TA82]. ATypI [Ano95a]. Aug [SK93]. August [ICC88, IEE92]. Authoring [Sta98]. Authors [Ano85a]. auto [HH94]. auto-hinting [HH94]. Autologic [Con89]. Automatic [AYO91, And90, Ano86b, Bro98, CFW82, CBE74, HHGH98, HZ93, Kar89, Kok94, MP94, MPZ95, Pet86, TCLM82, Zla91, Mil81, PM93, SR02, Sha03, Mal94]. Automatically [Ano85e, BS96, Hob93]. Avant [Ano93l]. award [Ano95a]. awareness [Man93]. Away [Adl94]. axcess [Ano93h]. Axioms [Knu92].

B [Car10, BMU95]. B1 [BMU94]. Babylon [Ano98]. Back [HYH89, HYH91]. Back-propagataion [HYH89]. Back-propagation [HYH91]. balancing [Ano93b]. Bangla [PC94]. Bangladeshi [PC94]. Banner [EMP90]. Barry [McC16]. bars [Flo94]. Base [Ao90]. Based [CJ83, Due93b, Dur93b, HHGH98, HZ93, HH01, Moh98, NN89, Nel85a, SFO88, Sch00, SR98, VK93, Ano90a, BBK^{+}19, BS96, BL94, HB91, Van93a, Kok94, MPZ95, PMZS97, Sch98a, Sha03, Ten90, TC92, Van93b]. Basic [Gho84, MS89, Her95]. basis [SA70]. Batch [Ano98]. Batty [Ano94u]. BBN [Ano96b]. Be [War09, Bar95]. Beach [CVP86]. Beauty [Kad11]. Beginning [SP87, Web15a]. behind [BH18]. Beispiel [SV73]. Berkeley [Anoxx]. Best [Che22, Tre94, Adv89, Ano91e, CGP95]. Beta [LC98, LH91]. beta-spline [LH91]. Beta-Velocity [LC98]. Better [Ada89, Pad92, UBR90, Van92b]. Beyond [Gos86, Orr94]. Bézier [Ell89, FB85, MP94]. Bibliographic [AKWC82]. Bidirectional [BBG85]. Bigelow [Lov24]. Binary [CY92, PS84, Pos89a]. bit [Har92]. Bitmap [Cas93, Alt86a, Alt86b, Par90, PB03, VCA87]. Bitmap-free [Cas93]. Bitmaps [Adl91, Ros91, Hob93, Sow91]. Bitstream [Ano95a, Ano92a]. Bleeding [Rey93]. block [SO94]. block-class-specific [SO94]. Blog [TBP16]. Blossoming [Ram87]. BM2FONT [Sow91]. Bodoni [Les15]. bonanza [EMP90]. Book [Ano88f, Ano90e, Ano94d, Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00, Tho19, AH92, Fen89, Fen91, Fen96, Gar10, Pan87, PWH90, RM87, Sea93, Wil04, Wil06b, WT06b, Che23]. Books [PRB87, Bur18]. borders [Mac78]. Borland [Van92b, Wil91]. boss [Wil94]. Boston [MA91]. Both [LZ86, Ano91c]. Bounds [Ano93b]. Box [Krt94, Ben93b]. Braille [Lor86]. Break [Pon89]. Brief [Sto16, Ste14, Ste18]. Briefly [Ano94d, Ano95a]. Brilliant [Sto16]. bring [Ano94c, Ano94v, Dea85]. British [Pou95]. Broken [KW93]. browsers [Phe93, Phe94]. Bruce [KB16]. Brush [ZST84]. Brush-Written [ZST84]. Brushless [JSG78]. Buch [Bar80, Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00]. Buchstaben [Poh15]. budget [LS91]. Buffalo [SA86, Sic85]. Buffer [Ano86i]. Buffers [Buy85]. bugs [Mid95]. BUILDER [Dak90]. Building [Sau86]. builds [Van92b]. Built [Wil06a, Gun85]. Built-in [Wil06a, Gun85]. bulletin [Xer80]. bundle [Van93a, Van93b]. Bureau [Ano95a]. Business [TBLJ90, Rei92]. buyer [DW90]. Buying [Ano91b]. Byblos [Ano96b]. Byrne [Ano94u].

C

[Jap78, Ste14, Ste18, Ano88a, Cha88, Jae91]. C. [Bar80]. CA [USE89]. Cache [KM89]. Caching [FK85, FK82]. calculated [Dea85]. California [Uni00, VP94b]. Callifonts [Bar95]. Calligraphic [Moh98, IWM94, Kim81, Par93b]. calligraphicum [Pan16]. Calligraphy [Her67, SK11, Cus11, MVM09a, Pan16]. Cambridge [USE87, AKWC82]. Can [Ano91c, Ano91e, Bar95]. Canada [ICC88]. Capabilities [DM92, Mey94]. capturing [SR02]. Cards [Cus11]. Career [Sto16]. Carl [Ano95a]. Carol [Sto16]. Carter [F.10, Lan94]. Cartography [Her69, Her95]. Cartridge [Ano86h, Egl90, Hew89]. cartwheels [Kim81]. Cary [Car10]. Case [Per88]. Casting [HH86, Nel85b]. Catalog [Ano94d, Anoxx, Hew88, Hyp90, Kim81, Pou95]. Catalogs [Ano95a]. catalogue [Sic85, SA86]. cause [Mid95]. CCIGS [CY92]. CD [Ano93d, Ano93e, Ano94b, Hav93]. CD-ROM [Hav93]. CDs [Ano94d]. Cedar [Tei85]. Centaur [KB16, Mac35]. Centennial [Gei15]. Center [Ben93a, Rie89, AKWC82]. centroid [AIDG95]. cents [Gab93]. century [Sou05]. CEO [Ano97b]. CGM [Mum89]. Chain [TKE83]. chairman [Ano97b]. Challenges [BH25]. Chameleon [Sey94]. Chancery [Hol80]. change [Ano97b, Flo94]. Changeable [BG73]. Changed [Wal13, Dam16]. Changes [BM87, Mey94, Knu89, Zap00]. Changing [Ano86b]. Channel [Ano93o]. Chapter [Kin05]. Character [Ano86e, BG73, CKM86, CK86, CC83, Che94, CBE74, DH83, Due93b, Due94, Dur93b, Gon98, Her87, HD88a, HWY92, Jap78, KM87, KM83, LB86, McC91, MRS83, Mei80, Mei81a, Mei81b, MS84, NSA90, Rei66, SP87, Som90, Ste87, TX92, Tri69, UKKS84, UM85, WW82, Wol86, Wol87, Ada91, All94, Ano96c, AIDG95, Bow68, DeC83, ELS94, HHG97, HD88b, Woo87, IO93, KD94, Krt94, Lam80, LK95, Lin95, LCO94, NNMS94, SZ95, UBR90]. characteristic [Glu69]. characteristics [BL94]. Characterize [RN83]. Characters [Ano85d, Ano85k, CCH88, Chu84, Ell89, GWW84, HDK68, KPB87a, KB86, KPB87b, LZ86, MBG77, NF87, OO89, TL92b, Ume82, ZST84, AV94, Ano96a, Cro78b, Cro78a, Dea85, Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00, Hus95, PM93, Pow90, RNR94, TC92, War80, Zap00, Fau04a, Fau04b]. Charles [GS21]. charting [Pat92]. chat [MVM09a]. Chel [Ano86f]. ChemText [Mol89a, Mol89b]. chess [Lar89, Tof90, Tut92]. chi [Jon95]. chieh [Jon95]. Chinese [HG84, ICC88, TL92a, Won93, Ano94e, Ano96a, Ano96c, BS96, CCL88, Cha83, CY92, CCH88, CM94, CC83, DL91, Ell89, GWW84, HG83, HZ93, HWY92, IBC87, IWM94, Lam80, LZ86, Lin95, MPZ95, Mei80, Mei81a, Mei81b, MS90c, PMZS97, PM93, SZC94, TC92, Ume82, UM85, WH95, Yi92, ZST84]. Chinese/Japanese [IBC87]. Chip [CP91a, Ano96c]. choice [Ano91e, Ano93f]. Choosing [Ano95d]. Christopher [Dur22]. Chronicle [Zap07a, Zap07b]. Chronik [Zap07c]. Chuck [Ano94u, Lov24]. Cie [CP91b]. class [SO94]. Classic [Sol83, Ano94g, Ano94h, CGP95]. classics [Lat95]. Classification [Glu69, CY92, ELS94, LCO94]. classifier [BN94, CDV95]. Classifiers [SV73]. clears [Swz93]. Cleary [Ano94d]. Clinic [Cam92]. clip [Van93a, Van93b]. cluster [UB95]. clutter [Swz93]. CM [Knu96]. Coast [Kin05]. Cockpit [DB93, UB95]. Code [Ano88a, Jap78, Lin95, Mil81]. Coded [TKE83]. codes [Jac82]. Coding [CCH88, IWM94]. coexistence [Ano91e]. Coherent [HHG97]. collaboration [Zap01]. Collection [Ano93t, Ano24, Ado87, Ano93d, Ano94g, Ano94h, Ano94i, Car10, Fon92, WT06a]. Colophon [Ano85b]. Color [Ano93a, Dun92, RRW93, Seb91]. Colorado [Sou98]. colorgraphics [Chi84]. Columbus [JN93]. Column [Ano85e]. Come [Kel19b]. comes [Van93a, Van93b]. Command [Ano86i, Mid95]. commands [Mid95]. Comments [TBLJ90, Mit92]. Committee [Doh81]. Common [Dun92, Dea85, MCL92, Orr94]. Commonwealth [Rie89]. communication [IEE93]. Communications [Gai85, MCL92]. Como [Luc85]. Compact [LB90]. Companies [Pon89]. Companion [MFW23, MFB^{+}23, Fon92, Kna95]. Comparative [VK93, Bau88]. Compared [GM85]. Comparison [SV73, MRT89]. Comparisons [Dea92b, MBG77, LCO94]. compass [Dea85]. compatibility [Egl90, NW83]. compendium [CSS93, K^{+}02, KGR03, KGR07]. Complete [Sol83, Cen88, Lou91, Lou91]. completeness [Kna96]. Complex [Cor00, CS87]. Complexity [Ros91]. Components [HH01, WH95]. composite [Ado89]. compositing [Emm95]. Composition [Sch98b, Sch00, Ken86, Lam80, LK95, PMZS97]. composizione [Bec91b]. Compressing [Ros91]. Compression [Ano86c, BHW83, CY92, IKM82, Won85b, SO94, Wan80, Won85a]. Compression/Decompression [Ano86c]. Computation [RN83, NC89]. Computational [All91, Par93a]. Computations [INT91]. Computer [AKWC82, Ano98, AOM95, Car84, Cat83, DMK69, Gai85, Gos86, Hew85, Her72, Her81, ICC88, CVP86, LS84, McC91, MNSJ99, Ost98, Rie89, Sau86, Sau87, Sch98b, SK93, TBLJ90, USE87, USE89, Wag85, YWL89, ZST84, Ano97b, Bie90, Sea93, TL92a, Ueh90, WH76, Ber96a, Knu86d, SK93]. Computer-Aided [Car84, LS84, Sch98b]. Computer-Generated [Ost98]. computerisation [Hus95]. Computerized [Chu84, ST85a, ST85b]. Computerphile [Bra13]. Computers [Her67, Knu89]. computing [Ano94e, Ano97b, Hol87, Rie89]. Con [Bec91b]. Concept [Knu81, TKMM95, Knu82]. Condensed [MHC98, Cen88]. Cone [Lan94]. Conference [AOM95, BDV95, Dés86, Fur90, ICC88, IEE88, CVP86, IEE86, IEE92, Luc85, Rie89, SK93, Uni00, Ano93a, Ano94e, Ano94a, Ano94q, TL92a, Van86, AGS93a, AGS93b]. Conferences [AH89]. Congeniality [Wen67, Wen68]. Congress [Pie91]. conjuring [Emm95]. Connect [Ram87]. Connect-the-Dots [Ram87]. Connection [Uni00]. Connectionist [GSKM92]. connectivity [Krt94]. Considerations [Bow68, PRB87]. Consortium [Flo94]. Constraint [Ada89, Nel85a, Sha03]. Constraint-Based [Nel85a, Sha03]. Constraints [Her87, SR98]. Construction [UM85]. containing [Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00]. contenente [Bec91b]. Context [Raf93, KH94, Ten90]. contextually [Coh93]. Continued [Ano97a]. Continuity [FB85]. contour [NSA90, SZ95]. Contoured [HWY92]. contribution [Cus11, WH76]. Control [BS84, FLS78, Mac91b, May85, RS81, SP87, Ano92f, Ben93b, Bro91, Pad92, RNR94]. controller [Ano90a, Nas94]. convenience [Dea85]. Convention [McV10, Rie89]. Conventions [Flo94, Tur00b]. Conversion [Ano96a, BDN83, LL94, Mal94, Mal95]. Convolution [NF87]. cool [Sey94]. Coordinate [Due93a, Dür93a]. Coordinate-independent [Due93a, Dür93a]. coordinates [WH76]. Corel [Ano98, Ano93f, Van93a, Ros89, Van93b]. CorelDraw [Van93a, Van92a, Van93b]. Corp [Ano98]. corral [Ano91g]. correcting [BN94]. corrections [Bar80]. Cost [BG73, Buy85, Col86, HH83, Gun85, Joc90, Lin91, Bee83]. Counterpoint [Hal07]. Course [Kre90]. CPU [KKM^{+}93, KKM^{+}94]. CRAFT [RRW93]. Create [Buy85]. creating [Wil06b, WT06b]. Creation [GWW84, Ano94v, McC87]. Creative [Che23]. Criterion [Yi92]. Croatian [Žub96]. Cross [Bro94]. Cross-platform [Bro94]. CRT [MRS83, Seb91]. Crunchers [USK^{+}17]. Cruz [Ano94d]. Crystal [Ano86d, McV10, War09]. Crystallographic [AKWC82]. CTEX [Lee94]. Cubic [KM83, LS84]. Cubics [Kla94]. CUG389 [Gag93]. Culture [Ber02]. curator [Sol87a]. Current [Ueh90]. Cursor [Ano85k]. Curve [HWY92, IO93, Kok94, LH91, MP94]. Curve-Contoured [HWY92]. Curves [FB85, GM83a, Ell89]. CurveSoft [Ano98]. Custom [RS81, Ano92f]. Customizing [RRW93]. cut [Zap00]. Cutters [USK^{+}17]. cutting [Kar91]. CVPR [CVP86]. Cwpr [IBC87]. Cyrillic [ZZ96, Bee85a, Bee85b, Str93a].

D [Orr94]. Dànielle [Dur22]. Darmstadta [Zapxx]. Darstellung [Kar92]. Data [AKWC82, CY92, IKM82, SR80, SO94, Egl90]. database [Bro94, Van93a, Van93b]. Databases [Kim90]. Datasouth [Bry84]. DaType [Sch98a]. David [KC87]. Day [Adl94]. days [Str93a]. DC [JSG78, Kna96, Kna95]. dc-fonts [Kna95]. Dec [TL92a]. Decades [Kar98]. Decisions [Wag85]. Declarative [NNGP89]. Declares [Ano90d]. Decompression [Ano86c]. Decoration [Bro98]. Decorative [Ano87b]. Dedicated [CP91a]. default [Ano91c]. defects [Ada91, UB95]. deformations [Par93a]. del [TP90a]. delivering [Ano93b]. Delivers [RS81]. delivery [Pat94]. demands [Nas94]. Demonstration [Dro85]. density [Lam80]. Dental [McC91]. Depressed [Wal05]. Derivation [WW82]. Description [Due93a, Due93b, Due94, Dur93b, Flo94, Kar94b, Dür93a, Kar94a, Kom91, Ste14, Ste18]. descriptions [CDV95, Rul87]. Design [Ano87d, Ber02, Big85, Car84, Chu84, Cyr88, DL91, Dre85, Due93b, Due94, Dur93b, Eng87, GM83a, Got85, GM85, HP98, Kad11, KM89, Koh89, LS84, May86, McC16, McC19, Mei80, Mei81a, Mei81b, Moh98, Moy95, NKMO89, NNGP89, PRB87, Rug83, Sch98b, SR98, Sto16, TX92, Tri69, Ung18, Wag85, Wit88, Zap85, Ada86, Pos89c, Bar86, Ben93a, Big86a, Big86b, BH93, BH18, Bow68, Cha83, Cus11, Gar84, Gho84, LK95, Par93a, Par93b, Rul87, Sch98a, Som90, Sou05, Tur00a, Wil04, ZL87, Zap01, McC16]. Designed [BH14, Kel19a]. Designer [Web15b, Bur18, WT06a, Wil06b, WT06b]. Designers [Pou12, Wil04]. Designing [Con98, Ear02, Sou85, Sou86, WH95, WT06b]. Designs [F.10]. Desk [Yuz89]. Desktop [Ano94f, Hol87, RM87]. determination [DeC83]. developed [Ano96c]. developers [Bro94]. Developing [Man93]. Development [SP87, Wag85, Bro94, Har96, LeV81, MS91]. Developments [GM85, Zap07a, Zap07b]. Device [Bro85, IBC87, RAK83, Rul87]. Device-Independent [Bro85, IBC87]. Devices [Con98, DS88, Gun85, Hug88, Bee83]. DeVoss [Dur22]. diacritics [MVM09a]. Diagnostics [EM84]. Dialogs [Dun92]. dialogue [Ada86]. Dictionary [Jae91]. Did [Kel19b]. Dies [Fox16, Web15b]. Differencing [Kla94]. Different [Har98]. Differentials [Wen68]. Digital [AH89, AGS93a, AGS93b, BBK13a, BBK13b, BH25, CFW82, Col86, Con98, DM89, Gar84, Kar91, Kar92, Kar94a, Lan94, McC20, MA91, Sta85, Wad84, Ada86, Ano96c, BS82a, BS82b, CGP95, CSS93, GC88, K^{+}02, KGR03, KGR07, Knu87, KMU94, Pat94, SR97, Sha03, Tay85, Tho90, Wuj88]. Digitale [Kar92]. Digitalisation [Tam89]. Digitalisierung [Zap00]. Digitization [MS90c, Zap00]. Digits [Ast72]. Dimensions [AKWC82, Mid95]. DIPS [FN85a, FN85b]. DIPS-106 [FN85a, FN85b]. Direction [Ano93y]. Directions [Knu79]. Directory [Gre95, GMP92a, Tay85, GMP92b, Lat95]. dirty [Ano95d]. disco [Bec91b]. Discrete [WH86]. Discrete-Element [WH86]. disk [Ano88b, Ano88c, Ano88d, Ano89a, Ano89b, Ano89c, War84]. Displacement [AYO91]. Display [AMS83, Ano86i, Ano91h, Con98, Cro78b, Cro78a, Egl91, Fon92, Mod93, UB95, War80, Bee82]. Displays [MRS83, WH86, Niz90, Tho90, Wil80]. distance [Wil80]. distinction [Way92c]. Distributed [KM89, dGH89, MCL92]. distribution [Bee85b, ELS94, Bee85a]. distributor [Int82]. Districts [MVM09b]. dithering [AIDG95]. DITROFF [BBG85, Pow90, SB92]. DITROFF/FFORTID [BBG85, SB92]. Do [Bry84, Ano95d, Swz93]. Do-It-All [Bry84]. doable [Min91]. Document [Fur90, KH96, SO94, SP87, AJ92, Big86a, But94, Kom91, Sha07, Van86, VP94a, VP94b, WCW82]. Documentary [Hus07]. Documentation [DB93, Dés86, Luc85, Mid95]. Documents [Cam91, KM89, TCLM82, ACG94, Ano93b, Les95, Tay91]. Doing [McC19, And89a]. Domain [Phe93, Phe94]. Domain-specific [Phe93, Phe94]. Don [Zap01]. donated [DM92]. don’t [Ano92b, MAYB02]. Dot [BM87, DiV84, MBG77, DeC83, Ken86, Niz90, UB95, Wil80]. Dot-Matrix [DiV84, Ken86, UB95, Wil80]. Dots [Ram87]. Douglas [Tho19]. down [War84]. download [Ano88b, Ano88c, Ano88d]. downloadable [Pos89d, Bee83]. downsample [ACG94]. DP [Cat83]. dpi [Mac91b, Mac91a]. Dr. [TBLJ90]. Draft [BDN83, Flo94, Kit84]. Draft-To-Text [BDN83]. Draw [Ros89]. drawing [Ano94v]. dream [TS94]. Drive [JSG78]. Driven [Ada89]. Driver [Ano98, Gun85]. Drives [Ano93e]. DTP [Mur90]. Dual [TA82]. Dual-Font [TA82]. DXF [Gag93]. Dynamic [AB89, AV94, KG77, Sta93a, TKMM95, Coh93]. Dynamo [Bry84]. Dyslexia [RBY16].

Earth [Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00]. eases [Nas94]. Eastern [Rei93b]. Eastern-language [Rei93b]. Easy [Cam91, Egl90, Emm95, WT06b]. Ecole [AH89]. Economist} [Nic91, Nic91]. Ed [Ano94d]. Edge [Rey93]. Edged [WW82]. Edges [Ano91h, Egl91]. edited [Dur22]. editing [Ano93k, Int82, Int83]. Editor [Chi84, Cow87, Sim88, Sym81, TL92b, Alt85, Alt86a, Alt86b, Cyr88, Min83, War84, TS85, VCA87, VCA88, VCG^{+}89, VCGA90]. editoria [Bec91a]. Editorial [AGS93a, AGS93b, And89a]. editors [Her79]. educators [Ano93d]. Effect [RBY16, BSS92, Ken86]. Effects [AMS83, Dea93a, San92, Seb91, DeC83, Ste92, Ten90]. Efficient [IKM82, KB86, Koh89, Lam80, TC92]. EGA [Cha88]. Egyptian [Hug88]. EHE [Yi92]. Eighth [IEE86]. Einstein [Gei15]. Electric [TBLJ90]. Electronic [Fur90, KM89, Mar87, SP87, Xer79, Xer80, Ano93b, Big86a, Xer82a, Xer82b]. electronics [BS81]. Electrophotographic [BS84]. Element [WH86]. Elements [Con98, SR96, LK95]. elettronica [Bec91a]. ElseWare [Ano93g]. Elumaze [Ano98]. Embedded [KKM^{+}94, KKM^{+}93]. embedding [MG95]. Embodied [BS84]. Emigre [Ano94d]. Empress [Ano98]. emulates [Ano93g]. Enable [Mey94]. encoded [RNR94]. Encoding [PS84, SP87, Lam80, ZZ96]. encodings [Jef93b]. Encryption [Pon89]. encyclopaedia [Hea05]. end [GSMN93, LS91]. Ending [Kur93]. Engine [Cor00, Orr94]. Engineering [Kad11]. engines [Bil87]. English [And89b, And89a, AO89, AB91, CP91b, Dar89, Fer90, Knu91, Lou91, Mac91a, Nic91, TL91, Adi90, Lan40]. Enhancements [DM92, SR97]. enhancing [Ste92]. enjoy [Ano91e]. enthaltend [Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00]. Entwicklungen [Zap07c]. envelope [Ano92f]. ENVI [Ano98]. Environment [Ada89, SD86, dGH89, Pos91c, Bar86, Dés84, Hol87, DP93]. Environments [Pet93]. Eólico [TP90b]. Éolien [TP91]. episode [Ten90]. episode-based [Ten90]. equalization [ACG94]. equipment [Kan84]. Era [Sta85]. Erdkreises [Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00]. Ergonomic [MRS83]. errata [Knu89]. Errors [AMS83, ELS94, WW93]. Erzeugung [Glö92]. especially [Dea85, Mac91b]. essentials [Way92b]. estimation [Kop93]. Etruscan [MVM09b]. Euler [Sie85]. Europäischer [TP90c]. European [Dés86, Fer90, Kna95, Luc85]. Européen [TP91]. européennes [Fer90]. Europeo [TP90a, TP90b]. Evaluating [Ada91]. Evaluation [VK93]. Evaluations [Tam89]. even [Van93a, Van93b]. Everyone [F.10]. everything [Dam16]. Everywhere [Web15b]. evidence [Kan84]. Evolutional [Tam89]. Evolving [Lee94]. Exact [Kla94]. Example [Bor92, Due93a, SV73, Dür93a, GSKM92]. examples [Pan16]. Exercising [Ano91d]. Exhibition [SK11]. EXP [MS91]. Expanding [Rig14]. Experience [CKT80, San92]. experimental [Cha83, DeC83, Sch90]. Expresiv [Ano94g, Ano94h, Ano94i]. extend [Hol87]. extended [Rah93]. extensibility [Sta93b]. Extensible [Cor00]. extensions [Ado89, Pos92b, Bie90]. Extraction [HHGH98, SR96, MPZ95, PM93, Par90, TC92]. eye [DeC83]. eyes [Lov24].

Face [Gla04, Ano91h, IEE93]. Faces [Kin05, Cen88, Pre57, Wuj88]. fading [Rei92]. Fail [Ast72]. Fail-Safe [Ast72]. failures [UBR90]. fairly [Emm95]. fait [And89a]. fallacies [Mid95]. Falsified [TCLM82]. Families [Ano87c]. Family [Dak90, DiV84, Lan40, Big86a, Hew88, Hoe95, MS90b, MS90a, Tob84, Dak90, Ano86f, Ano86g, Ano87b]. Fancy [Hof83, Sof87, Hoe95, Min91]. FAQ [The00]. Fast [Ano90a, Cra90, NF87, Rei92]. Faulmanm [Bar80]. fc [Kna93]. fc-fonts [Kna93]. Feature [SR98, TL92b, Par90]. Feature-Based [SR98]. Feature-Rich [TL92b]. Features [HD88a, Emm95, HD88b, Van93a, Ste14, Ste18, Van93b]. February [Rie89, VP94b]. Fédérale [AH89]. feedforward [MRT89]. Fellowship [Ano94d]. FFORTID [BBG85, SB92]. fichiers [TL91]. Fickle [Bee81]. Fiddler [Ano94s]. fiddling [Sey94]. Fifth [USE89]. Fifty [Wal13]. File [Ano91h, TL91, WDL87, Pos92a, Ano84b, Ano85c, Ano91f, Ano92j, Egl91, Phe93, Phe94, Rok85]. file-tree [Ano91f]. Filenames [Mit92, Ber90]. Files [LB90, LL94, Pos89a, Pos89b, Pos90, Ano95b, PB03, Fuc81]. fill [SZ95]. Film [Hus07, Ano94t]. Filter [Ano98]. Filtering [NF87]. Finally [Kay96, Ano91h]. Find [Ano91e, Ear76, SG02]. finds [Van93a, Van93b]. Fine [Con89]. First [Ano94j, Luc85, SK93, Kna95]. Fitting [Bro98, Nel85b, IO93]. Fixed [PZL94]. Fixed-Pitch [PZL94]. FL [CVP86, TL92a]. Flow [Pat92]. flue [Fonxx]. Flying [LM91, Emm95]. Focused [Yi92]. Font [AYO91, AMS83, Ada89, AS89, Adi90, Adl94, Ado87, Pos91a, Adoxx, AM86, And89b, And92, Ano84b, Ano85d, Ano85k, Ano86b, Ano86c, Ano86d, Ano86e, Ano86h, Ano86i, Ano87d, Ano89a, Ano89b, Ano89c, Ano90b, Ano90d, Ano90f, Ano91b, Ano91d, Ano92c, Ano92d, Ano92l, Ano93c, Ano93g, Ano93m, Ano93n, Ano93p, Ano93t, Fonxx, Ast72, Ben93a, BDN83, BG99, Big85, BG73, BU78, BHW83, Bro85, Bro98, CKM86, CK86, Cha88, CCH88, CM94, Chi84, Chu84, CP91a, CBE74, CFG92, DB93, Dea92b, Dea93a, Dea93b, DM92, Doh81, DL91, Due93a, Due93b, Due94, Dun92, Dur93b, Ear76, EM84, Eng87, Fen94, Flo94, Ano93h, Fow73, Fuc81, FK85, FN85a, FN85b, Gei15, Gla90, GL84, GMP92a, HP98, Her89, Hoe94, Hof83, HD88a, HD88b, HZ93, HWY92, JSG78, Jon94]. Font [KPB87a, Kar89, Kar94b, Kar94c, KM89, KKM^{+}94, KB86, KPB87b, KH94, KH96, Kit84, Knu81, Knu88b, KM87, Kur93, LB86, LO90, Lee94, LS84, LNN95, LNN00, LZ86, LH91, Mac78, MBG77, McC84, McC91, MO92, MB93, MRS83, Mey94, War84, MN77, MS84, NN89, NNGP89, OO89, PS84, Pan87, Pea92, PZL94, RSSW89, RS81, RG93, Rei66, Rei93b, RBY16, RRW93, Ros91, SV73, SZC94, Sha94, Sim88, Sla09, SP87, SBSG89, Sta98, Ste92, SSST83, Sym81, Tex86, TKE83, TKMM95, TX92, Tim87, Tim88, TCLM82, Tre94, Tri69, TL92b, TA82, UKKS84, Ume82, UM85, VK93, Wad84, WDL87, Wan80, Way92a, Way92b, Way92c, WT92, WB95, Wil80, WH86, Wit88, Wol86, Wol87, Won85a, Won85b, Yi92, ZL87, ZI95, dGH89, Ada86, Ada91, Pos89a, Pos89b, Ado89]. font [Pos89c, Pos90, Ado90a, Ado90b, Pos91b, Pos92a, Pos92b, Ado92, Ao90, Alt85, Alt86a, Alt86b, And89a, AB91, AH92, Glu69, Ano85c, Ano91c, Ano91e, Ano91a, Ano91g, Ano91h, Ano92a, Ano92b, Ano92e, Ano92f, Ano92h, Ano92j, Ano93b, Ano93f, Ano93k, Ano93u, Ano93w, Ano93x, Ano94c, Ano94g, Ano94h, Ano94i, Ano94k, Ano94t, Ano94v, Ano95a, Ano95b, Ano95d, Ano96a, Ano96b, Ano96c, Anoxx, Ate92, BMU94, BMU95, BN94, Bar86, Bar95, Bee83, Bee82, Ben93b, Ber96b, Bie90, BH93, Bit91, BL94, Bow68, BSS92, Bus92a, Bus92b, Cen88, Cha83, Cyr88, DeC83, Dea92a, Dea92c, Dub90, Dür93a, EEN90, Emm95, Fen89, Fen91, Fen96, Gar84, Gho84, GL83, GC88, GMP90, GMP92b, Hew88, Hew89, Har92, HFW^{+}96, HN64, Her79, HH86, Hug88, Hyp90]. font [Int82, Int83, Woo87, Jef93a, Jef93b, Jef94, KMNO89, KNT^{+}01, Kan84, KKM^{+}93, Kay96, Ken86, KKL89, Kla93, Kna86, KMU94, Kok94, Kop93, Lan40, Lar89, cYL88, LK95, Lin95, LCO94, Lor86, Lou91, Min83, MP94, MPZ95, Mac35, MG95, May86, McC87, Mic92b, Mic92a, Mid95, Mil81, MS89, MS90b, MS90a, Mod93, MRT89, MS91, NKMO89, NNMS94, NW83, Orr94, Pad92, PMZS97, Par93a, Pat94, Pet86, PWH90, Phe93, Phe94, RNR94, Rei92, Ris88, Rok85, Rul87, RM87, San92, SR02, Sea93, Seb91, Sic85, SA86, SB93, Sof87, Sta93b, Sta87, Sta89, Ste14, Ste18, Str93b, Tay85, Tay91, Ten90, TS85, TL91, TC92, Tut92, UBR90, VCA87, VCA88, VCG^{+}89, VCGA90, Whi27, Whi47, Whi91, Wil94, Xer79, Xer80, Xer82a, Xer82b, Zap01, ZZ96, ZI93]. font [di 95, ALS88, Ano94d, Ano94k, Ano94s, FK82, HFW^{+}96, HG83, Jac82, Knu82, Knu88a, Ols95, Sch92b, Sey94, Whi81, Wuj89, Coh93, HG84, HK87, McC83, Tob89, Tof90, LeV81]. Font- [HD88a, San92]. Font-and [HD88b]. font-delivery [Pat94]. font-matching [BL94]. font-names [Lan40]. font-scaler [Sta93b]. Font/DA [ALS88]. FONTastic [Alt85, Alt86a, Alt86b, VCA87]. FontBook [CSS93]. FontChameleon [Ano94l]. fonte [Lou91]. Fontes [Fer90, And89b, AB91, Dar89, Mac91a, TL91]. FontLab [Ano94v]. FontLib [Gag93]. FontMixer [Ano94m]. FontMonger [Ano93i]. Fontographer [Moy95, Dak90, Ano94n, Ano94o, VCA88, VCG^{+}89, VCGA90]. Fonts [Adl91, And90, AB89, Ano90a, Ano91f, Ano95b, Ano95d, BBK^{+}19, von92a, BH91, BB89, Cam91, CFW82, CY92, CM94, CC83, Col86, CP91b, DH83, Fox16, Gos86, Grä93, GMP92a, HBBG95, HH01, HWY92, Kar92, Knu87, Knu90b, Koh89, Mac91b, Mar87, McC07, Min91, Mit92, MS90c, MHC98, NN89, Par88, ST85a, ST85b, Sch92a, SR96, SR98, Sol83, Wil91, von92b, von93, Pos89d, Pos91c, Ano84a, Ano91e, Ano92b, Ano93g, Ano94g, Ano94h, Ano94i, Ash90, Bee85a, Bee85b, Bil87, BSS92, CGP95, Chi88, Cra90, Dar89, DD15, Dés84, DM95, DP93, Fer90, Gar10, Gen87a, Gen87b, Gla90, Goo95, GC88, GMP90, GMP92b, Gun85, Har93, Har96, Hea05, Hen89, HB91, Hoe95, Van93a, Hor93, HP93, IWM94, IO93, K^{+}02, KGR03, KGR07, Kna95]. fonts [Kna96, Knu86a, Knu90a, Kom91, Lam80, Lev88, Mac91a, Mal94, Mal95, McW87, Nic91, Par90, Pow90, PB03, RM87, Sau87, Sch88, SR97, Sta93a, Tam89, TBP16, Ueh90, Van93b, Wil94, Win90, WH76, YWL89, Zla91, Ano86f, Ano86g, Ano87b, Ano87c, Ber96a, Glö92, Har93, Kna93, LW88, Sau86, Bee81, Ber90, Con89, Knu96, Les95, MG95, MW92, Rah92, Rah93, Tob84, Van88, Won93, YWL89, Žub96, Par89]. FontScope [Ano98]. FontWINDOW [Ano89d]. Forensics [BBK13a, BBK13b]. Form [GM85, KW93, Cen88, Bee82]. Format [Hos24, PZL94, WDL87, Ado90a, Ado90b, Pos92a, Pos92b, Ado92, But94, Dea92c, Ano84b, Ano85c, Rok85]. Formate [Kar92]. Formats [Kar92, MO92, Kar94a]. Formatting [BBG85, SB92]. Forming [ZST84]. Forms [Upd22, Upd37, Upd62, Upd80, WD18, Ada86, HHG97, Van93a, Van93b]. Formulation [Moh98]. FORTRAN [Her69, Her70]. Forum [Hoe94]. Forward [Kar92, Kla94, Tre94]. Found [Web15b, Web15a]. Foundation [Woo87]. Foundries [Gre95, K^{+}02, KGR03, KGR07]. foundry [Ano93w, Ano93x, Lan94]. Fount [Fou91, Dea85]. Fourth [USE87]. FoxPro [Bro94]. Fractal [IWM94]. Fractured [Kur93]. Fraktur [KW93]. France [Ano87d, Dés86, IEE86]. Free [Ano91f, Cas93, KD94, Sey94]. FreeType [The00, The03, Tur00a, Tur00b, Tur03]. FreeType-2.1.7 [The03]. French [Dés84, BWT85]. frequency [Ten90]. front [GSMN93, Kom93]. Full [PRB87, Cen88]. Fun [DD15, Knu90b, Ash90, Knu90a, Way92a]. Function [KH96, Ado87, Adoxx].

Fundamental [GM85]. furniture [Mac78]. FUSE [Ano95a]. fuss [Joc90]. Futura [TL17, Tho19]. Future [Knu91, Zap85, Mur90].

GA [Ara95]. Gain [Pad92]. Gainesville [TL92a]. Gaithersburg [Fur90]. Galliard [Car85]. Garde [Ano93l]. gebrochenen [KW93]. Geheimnis [SG03]. Geleitwort [Kar92]. GENDRIV [Gun85]. General [Gei15, KNT^{+}01, Par93a, TBLJ90]. generalization [GSKM92]. Generalized [WDL87]. Generated [Ost98]. Generating [DH83, Hob93, HWY92, Sch92a, BS96, HFW^{+}96, HK87]. Generation [AYO91, Ada89, And90, CC83, Her87, HBBG95, Kit84, KM83, MB93, OO89, PS84, ST85a, ST85b, SZC94, TKMM95, UKKS84, Van93a, KKL89, Kir05, Kok94, LH91, MP94, MPZ95, Ris88, SZ95, Van93b, Win90, Zla91]. Generator [SSST83, May86, Mod93]. Generic [Ano84b, Ano85c]. Genius [Gei15]. Geometric [FB85, PMZS97]. Gerard [Bur18]. German [Ano88e, BWT85, Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00, Fau04a, Fau04b, KW93, Zap00]. Geschichte [KW93]. Geschke [GS21]. Gestaltungsbuch [Rud88]. Get [Ano88e, Gre91, Kay96, DW90]. Gets [Ano90a, Cra90]. Getting [Mac91a, Min91, Rei93a]. Ghost [Ano85k]. Giambattista [Les15]. Give [Kur93]. gives [Egl90]. Glance [Ano90b]. Glassner [Gla04]. Global [Uni00, KH94]. Glyph [Hud00, Tur00b, MCP91]. Glyphs [Wil06a]. Go [TBP16]. Goblet [McV10, War09]. goes [Ano94t]. Going [Kel19b]. good [Bar95]. Goodbye [Kir05]. Gothic [Ano93l]. Goudy [Ben93a]. grâce [Mac91a]. Grammars [RN83]. Grand [Ash90, Knu90a, Knu90b]. Granjon [Car85]. Graphic [Car10, Con98, GM83a, Jap78, Wil91, WH76]. Graphical [HDK68, Sta98]. Graphics [Ano93j, Ano95b, DL91, DMK69, Gos86, HM78, McC91, MNSJ99, Nel85a, RS81, USE87, USE89, Ike95, Mid95, RNR94, Ros89, SK93, Wil89]. Graphite [Cor00]. graver [Pan16]. Gray [AYO91, War80, Ano84a]. Gray-Scale [AYO91]. Grayscale [HBBG95, MHC98, Cro78b]. Great [Bra13, Cam91, MC93, Ros89]. Great-Looking [Cam91]. Greek [Lev88, MW92]. grey [Cro78a]. Greyvalue [HD88a, HD88b]. Grid [Her87]. Gridfitting [And90]. gridfont [GSKM92]. Group [Ara95]. GS [Sim88]. Gudrun [SK11]. Guida [Bec91a]. Guide [Des99, LNN95, LNN00, Raf93, Ado91a, Ado91b, DW90, GMP90, Mic92b, NW83, Tex86, WT06b, Xer79, Xer82a, Xer82b]. Guided [RRW93]. Guinta [Ano85h]. Gutenberg [Kir05]. guter [SG03]. GVIP [Ike95]. GX [Ano94d, Ano94t].

Hacker [Raf93]. Hackers [Kad11]. Hackiana [Ano85i]. Hacking [Sla09]. Hague [IEE92]. Halftone [Mod93]. halftones [Knu87, Sow91]. Hallmark [Cus11]. Hand [Egl90, Nel85b]. Handbook [GMP92a, GMP90, GMP92b, RM87]. Handbuch [Poh15]. Handler [RG93]. Handling [Ano86h, KM87, Mar87, Ols95]. handwriting [DM95]. Handwritten [LC98, LZ86]. handy [Tob89]. Hangul [Chu84, Kim90]. HanZon [Ano90a]. Hardcopy [DS88]. hardware [Orr94]. Harsy [Ano85g]. hassle [Ano93b]. Having [Sha94]. HBF [PZL94]. HCI [AOM95]. HDTV [DC93, Mod93]. HDTV’92 [DC93]. Headliners [Ano94d]. Hebrew [Str93a, Tam89]. Height [Ano85e]. held [MA91]. hello [Kir05]. Helvetica [Har96, Hus07]. Her [Sto16]. Hermann [Kar92, Cus11, Hal07, Hol15, Kar92, Kel11, Kel19a, Pan16, Rom16, SK11, Web15b, Web15a, Zapxx]. Hershey [DH83, Gos86, WH76]. Herzog [Ano95a].

Hewlett [Ano93n]. hidden

[CWB93, HYH89, HYH91]. Hierarchical [Sch98b]. hieroglyphs [Hug88]. High [AYO91, AIDG95, BHW83, DiV84, KMNO89, KKM^{+}94, LS91, ST85a, ST85b, UKKS84, Ano96a, BS82a, BS82b, Ike95, KKM^{+}93, Lam80, MP94, ZL87]. high-density [Lam80]. High-end [LS91]. high-performance [Ike95]. High-Quality [DiV84, ST85a, ST85b, UKKS84, MP94]. High-Speed [KKM^{+}94, KKM^{+}93, ZL87]. Highlights [Ano97a]. Highly [IKM82]. HiJaak [Ano87a]. Hint [KKM^{+}94, Pos91b, KKM^{+}93]. Hinting [HZ93, Kar89, HB91, HH94, PM93, Sha03]. Hints [And90, Dea92a]. History [Bri16b, KW93, KB16, McC16, McC19, MVM09a, Per88, Upd22, Upd37, Upd62, Upd80, BH18, Bri16a]. HomePublisher [Ald93]. Hood [BL94, Ano90h]. Hoops [Orr94]. hot [Ano90a]. house [Way92a]. HP [Bro90, DM92, Egl90, Hew89]. HP-donated [DM92]. hsing [Jon95]. Hulls [Knu92]. Human [Gai85, KB86, AOM95]. Human-Computer [AOM95, Gai85]. Hybrid [Kla94, SFO88, Sch00]. Hypermedia [Ano98]. Hypertext [LM91, Bro91]. Hyphenation [Dés84].

IAPA [IEE92]. IBM

[RNR94, Ado91a, Ano90d, BS84, FLS78, Int82, Int83, Pap89a, Pap89b]. ICIAP [BDV95]. Iconographic [McC91]. Identification [KH96, TCLM82, ZI95, Kan84, KH94]. IEC [SBSG89]. IEEE [CVP86]. II [Ano93s, Egl90, Gen87b, Int82, MFB^{+}23, MA91]. III [Bro90]. Illusion [Wal05]. illustrations [Van93a, Van93b]. illustrators [Van92a]. Image [Ano93k, HD88a, IKM82, Yi92, Ada91, Ano94p, Ano95c, HD88b, SO94, SZ95, Win90, BDV95]. Images [CY92, CWB93, Kop93]. Imaging [AH89, AGS93a, AGS93b, MA91, Ano93a].

Immunity [BB89]. IMP [Car84]. Impact [Che94, DiV84, Pie91, Sha07, UB95, ZL87]. Impaired [GL84, GL83]. implementation [TS85]. Implementations [Grä93, Har98]. Implementing [Rah93]. Importance [DM89]. Important [Knu96]. IMPORTER [Dak90]. Impressions [Ano97a]. impressive [Van93a, Van93b]. imprimantes [Mac91a]. Improve [Mey94, UBR90]. Improved [Buy85, Cro78b, Cro78a, Tof90]. Improvements [Kna95]. improves [Ano93f]. Improving [Tay91]. Inc. [Ano98, Lan94]. Including [Sch00]. Increase [Mey94]. Incredible [Ano94p, Ano95c]. Incremental [NNGP89]. Incubator [Ano93i]. Independence [Ano90d]. Independent [Bro85, HM78, IBC87, Due93a, Dür93a, K^{+}02, KGR03, KGR07, Rul87, Wil89]. Indexes [BK88]. India [MNSJ99]. Indian [Moh98, RAK83, SR80]. Indic [MNSJ99]. Indie [K^{+}02, KGR03, KGR07]. Indo [PC94]. Indo-Bangladeshi [PC94]. Industrial [SFO88]. industry [Ano97b]. Infinifont [MB93]. Info [Swz93]. INFORM [FG85]. Information [Ano98, Bro85, II88, Jap78, Ano95a, Int82, Kom93, Phe93, Phe94, SBSG89, Ste14, Ste18]. Initial [Ano86i]. Ink [di 95]. Innovations [GM85]. Innovative [UB95]. Input [HDK68, RAK83]. Input/Output [RAK83]. insights [Wil06b]. install [Bus92a, Bus92b]. installation [Jef93a, Lou91]. Institute [Car10]. Instruments [McC87, Pet86]. Integer [Kla94]. integrated [Big86a]. integration [Van93a, Van93b]. intellectual [McC07]. Intelligence [Sta98]. Intelligent [Kar89, Sta93a]. Interaction [AOM95]. Interactive [GM83b, GWW84, Par93b, Rul87]. Interchange [Jap78, SP87, Kom93, SBSG89]. Interface [GL84, Wil91, Cra90, GL83, MS90b, MS90a].

International

[AH89, Ano94d, Ano94e, Ano94a, AOM95, Ben93a, Ber02, BDV95, DC93, Fur90, ICC88, IEE88, IEE86, IEE92, JN93, Kom91, Kom93, Pie91, Uni00, Van86, Ano94q, TL92a]. internationalization [Mac90]. Interpolating [FB85]. Interval [INT91]. Interview [Ano97b, TBLJ90]. introduce [Ano92a]. introduces [Lin91]. Introduction [Her89, GC88]. invariant [HD88b]. Inversions [Kim81]. Investigations [Sha07]. Invisibility [Rah92]. Invisible [War09]. IPCC [IEE93]. ‘i’rab [MVM09a]. IS&T [Pie91]. ISBN [Tho19]. Islamic [MVM09a]. ISO [SBSG89]. ISO/IEC [SBSG89]. issue [AGS93a, AGS93b]. Issues [KM89, Gho84]. Italic [Cha88]. Italics [Fow73]. Italy [BDV95, IEE88, Luc85]. ITC [Ano93l, Hol80]. IUC17 [Uni00]. IV [DC93, Her69].

JACM [CCL88]. jaded [Emm95].

Jailbreak [Bra13]. Janson [Sta85]. Japan [AOM95, DC93, Yuz89]. Japanese [Ano90e, BWT85, IBC87, Jap78, KM89, ST85a, ST85b, ZST84]. Jargon [Ano90f]. Java [Ano98, Ano97b]. JDBC [Ano98]. Jim [Sey94]. JIS [Jap78]. JIS-C-6226 [Jap78]. Johnson [TBLJ90]. Johnston [Dam16]. Joint [Ano94r]. Jos [Ano98]. Jose [Uni00, VP94b]. Jr [Car10]. juggler [ALS88]. July [AOM95]. jumping [Mid95]. Jun [Ano94q]. June [Ano94e, Dés86, CVP86, Rom16]. Juno [Nel85a]. Just [Gar10, CCL88, Emm95].

Kalb [Gre92]. Kanji

[AYO91, Due93a, Dür93a, FN85a, FN85b, OO89, Ros91, SSST83, TKE83]. Kanji-Printers [FN85a]. Katz [McC16]. Kawasaki [DC93]. Keeps [Adl94]. Key [BM87]. Keys [Ano91f]. Kharosthi [MVM09b]. Khmer [Har94]. Kind [HWY92]. Kindersley [KC87, Lop85]. Kis [Sta85]. kit [MS91]. Klassifikatoren [SV73]. knife [Van92a]. Know [Rei93a, Rei92]. Knowledge [Tra89, TC92]. knowledge-based [TC92]. Knuth [Zap01]. Korea [SK93]. Korean [Chu84, Kim90, LC98]. KY [Rie89].

labels [SB93]. Laboratory [DH83, Sch90]. Laden [LW88]. l’Algérie [Ham90]. Language [Ber02, KM89, Mei80, Mei81a, Mei81b, Ado89, Hol87, PC94, Rei93b]. Languages [ICC88, SR80, Kom91, TL92a]. Large [Buy85, GL84, KM87, WH86, GL83, NNMS94]. Large-Area [WH86]. larger [MAYB02]. Laser [DW90, Mac91b, Nas94, Lin91, Mac91a, May86]. Laser-printer [Nas94]. LaserJet [Bro90, Egl90, Hew89, Ano90a, Hew88, Neb93, PWH90]. LaserJets [Min91]. LaserTools [Ano92e]. latest [Ano94s, Pad92]. LaTeX [MFW23, MFB^{+}23]. LaTEX2e [Jef94]. Latin [Fer90, TA82]. Latin/Arabic [TA82]. latines [Fer90]. laurea [Bec91b]. Lausanne [AH89]. L’avenir [Knu91]. lawsuit [Ano95a]. layout [Rah93]. LCCD [Mei80, Mei81a, Mei81b]. leads [Mac78]. Learned [Knu85]. learners [Dea85]. Learning [Adl91, BBK^{+}19, CKM86, CK86, NNGP89]. Least [Kop93]. Least-squares [Kop93]. Legibility [BB89, MHC98]. Lemmens [Ano98]. less [Ano92f]. Lessons [Knu85]. Let [Ano95d, Tre94]. LetraStudio [Ano92f]. lets [Ano92b]. Letter [Wal05, WD18, Ada86, Ken86, San92]. letter-specific [San92]. Letterfontäne [Poh15]. Letterform [Rug83, Sch00]. Letterforms [Sch98b, Zap85, Man93, Dur22]. Lettering [BU78, Cus11]. Letters [Web15b, Kna95, San92, Zla91]. level [War80, Wil06b, Ols95]. LEXX [Cow87]. Leyden [Ano85i]. libraries [Ano92f, Orr94, Tay85]. Library [PRB87, Ado87, Cen88, Gen87a, Gen87b, War84, Tex86, Xer82b, Pou95]. Licensing [Mey94, Ano94t]. Life [Zapxx, Les15]. ligatures [Kok94]. Like [Gei15]. Likelihood [BSS92]. limelight [Ano93k]. line [Pou95, Wil89, ZL87]. Linear [SV73, Par90]. Linearer [SV73]. lingual [Gho84]. Linguistics [All91]. Linotron [CKT80]. Linux [Ano98]. Liquid [Ano86d]. Lisa [Wag85]. listed [Cen88]. literacy [Big86a]. literate [GW91]. little [Tob89]. LN03 [Sau87]. Load [Ano86i]. loader [War84]. Loading [FN85a, FN85b, Ano93f]. loci [Glu69]. Logical [Flo94, Bro91]. Logo [Che22, Knu86a]. Logotypes [Che22]. London [Ano94a]. look [BL94, Mac91a]. Looking [Cam91, Mac91b]. Looks [Ano94j]. Louisville [Rie89]. lover [Wre87, Wre88a, Wre88b]. lovers [RM87]. Low [BG73, Buy85, Col86, HH83, NN89, Ros91, BS82a, BS82b, Gun85, Joc90, Lin91, LS91, Par90, Bee83]. Low-Cost [BG73, Buy85, Col86, HH83, Joc90, Lin91, Bee83]. low-end [LS91]. low-resolution [Par90]. LSI [KMNO89, NKMO89]. Ltd [Ano98]. Lucida [Lou91, Big86a, BH14, BH18, BB89]. Lyon [Ano85g].

M [GS21, McC16]. MA [USE87]. Mac [Ano91c, Ano91f, Kay96, Tho90]. Machine [AM86, Rei66, Wil89]. Machine-independent [Wil89]. Macintosh [Har92, Flo89, Ado91b, Adv89, Alt85, Alt86a, Ano88b, Ano88c, Ano88d, Ano93w, Ano93x, Cra89, Fen89, Fen91, Fen96, How86, Rei93b, RM87, VCA87, Wil94]. macro [MS89]. macroistruzioni [Bec91b]. macros [Mil81, MS89]. Made [Wag85, Bar95]. magnifications [Chi88, Sch88]. Maintaining [BS81]. mais [And89a]. make [Ano91h, McC16]. MakeFont [Gag93]. Making [Ano90f, Hos24, Nel85b, TX92, Ste92]. Man [F.10, Phe93, Phe94]. Management [Ano98, Dea93b, dGH89, CFG92, Kay96, NNMS94, Wil94, Ber96b]. Manager [Liv95, RSSW89]. Managers [Ano91h, Ano92c, Egl91]. Managing [Kay96, Pea92]. Manipulation [Fur90, Adi90, Ao90, Ano91a, Big86a, Van86]. manner [Pan16]. manual [Bau88, McC87, Sta87, Sta89]. Manuale [Pan16]. manufacture [Ste14, Ste18]. manufacturing [Sou05]. Manutius [Ano85f]. many [Bur18, Wuj88]. mapping [MCP91]. Maps [GM83b]. mark [Ano85g, Ano85f, Ano85h, Ano85i, Ano85j, Ano94u]. Markov [CWB93]. Maryland [Fur90]. Master [Ano93y, Ber96a, Goo95, Har93, Pos92b, Ano93l, Ano94c, Dea92c]. Masters [Ano94d]. Mastery [Web15a]. matching [BL94, HB91, BL94]. materials [Ben93a]. Math [von92a, Jef93b, von92b, von93, Hor93]. Mathcad [Ano94s]. Mathematical [Ost98, Bee85a, Bee85b]. mathematics [FT96, KC75, McW87]. Matrix [BM87, DiV84, MBG77, Nel85b, DeC83, Ken86, Niz90, UB95, Wil80]. matrix/character [DeC83]. matter [MAYB02]. Matters [Dur22, WD18]. Matthew [F.10]. May [Ano87d, Ano94r, Luc85, Mor93]. McNealy [Ano97b]. meaning [Ste92]. mechanical [Ste14, Ste18]. mechanism [WW93]. Media [Ano86e, KG77]. medium [BS82a, BS82b]. meeting [Ano94r, Mor93, MA91]. MEL [Sch90]. Melbert [Car10]. Memory [Buy85, Nas94]. menu [Ano91a, Dub90]. MenuFonts [Ano92g, Dub90]. Mergenthaler [CKT80]. Mesh [Bro98]. message [Knu96]. Meta [Hoe94, Knu81, Knu88b, HG83, Knu82, Knu88a, HG84]. Meta-Font [Knu81, Knu88b, HG83, Knu82, Knu88a, HG84]. Metafont [Ano84a, Knu86e, Sou86, Ano84b, CCL88, Des99, Knu85, LL94, Sou85, Sou98]. METAFONTbook [Knu93, Knu86c]. METAfonts [Wuj88]. metal [HH86, Mac78]. METAPLOT [Wil89]. Method [Ano85e, Ano85d, Ano86d, FN85a, FN85b, HHGH98, HZ93, IKM82, NN89, PS84, Ros91, Sta98, TKMM95, Wad84, Wen68, Ada91, NSA90, PMZS97, TC92]. Methods [GM85, Sch00, Kan84, Kar94c, SO94, Sou05, Cla90]. Metric [Fuc81, Pos89b, Pos90, Pos92a, Kop93]. Metrics [And92, And89b, AB91]. Métrique [AB91]. Métriques [And89b]. MF [Ano86f, Ano86g, Ano87b, Ano87c]. s [Mac91b]. mi [Jon95]. Miami [CVP86]. micro [Sch90]. Microcomputers [DH83]. Microeconomic [Kre90]. Microsoft [Ano92h, Ano93m, Ano93n]. Microsystems [Ano97b]. Middle [Koh89, Rei93b]. Migration [YWL89]. Miscellaneous [Gen94, Phi94]. misconvergence [Seb91]. Missing [McC16]. Mistakes [Ano97b]. Mixed [LB86, Glu69, Tay91]. Mixed-Font [LB86, Glu69, Tay91]. Mode [Ano86i, NW83, Orr94]. Model [CC83, HB91, LC98, Sch98b, MCP91, McC87, NW83, Int82]. Model-based [HB91]. modeling [cYL88]. Models [BS84, KB86, CWB93]. Modern [Ber96a, Car85, Sau86, Sau87, YWL89, Bie90, Hug88, Knu86d, MW92]. Modification [CM94, GWW84]. Modular [SR80]. module [KNT^{+}01, Pet86]. modules [Van93a, Van93b]. Molecular [AKWC82]. Moment [Kan84]. money [Ano92f]. Monotype [Ano97a]. Monterey [USE89]. morphological [KD94]. Most [F.10, Orr94]. Motor [JSG78]. Mould [Nel85b]. MousePaint [Lor86]. Move [Kna96]. movements [DeC83]. MS [Bec91b]. MS-DOS [Bec91b]. Multi [AM86, Ano85d, Ano86e, Ano96b, CKM86, CK86, CCH88, EEN90, KM87, LZ86, Ris88, Sta89, SSST83, TKE83, Ume82, UM85, Pos92b, Ano96c, Gho84, Lin95, LCO94, Sta87, TC92]. Multi-Font [Ano85d, Ano86e, CKM86, CK86, CCH88, KM87, LZ86, SSST83, TKE83, Ume82, UM85, Ano96b, Ano96c, EEN90, Lin95, LCO94, TC92]. multi-lingual [Gho84]. Multi-phase [EEN90]. Multi-system [Ris88]. Multifont [RH75, Krt94]. Multilingual [RAK83, Ao90, KNT^{+}01]. Multiple [Ano85k, Ano93y, Ano94d, Ber96a, Dea92c, Goo95, Har93, HP98, dGH89, Ano93l, Ano94c, Ben93b, Van93a, Van93b]. Multiple-Alphabet [HP98]. Multiplying [Dea92c]. multipurpose [KKL89]. Muse [Adl94]. Museum [Ano97a]. music [Dea85]. muss [Joc90]. My [Adl94, Zap01, Ano91c, Ano92j, Gar10].

Name [GM83b, Lan40]. names [Lan40]. National [Ano97a]. native [Con89]. NATUG [Ara95]. NATUG-7 [Ara95]. Natural [Bri16b, Bri16a]. needs [Cus11]. Netherlands [IEE92]. Network [Ano98, FN85a, FN85b, Yi92, Ano96c, CDV95, cYL88, LO90]. Networked [Won85b, Won85a]. networks [Ano96c, AIDG95]. Neural [CDV95, cYL88, Yi92, Ano96c, AIDG95, LO90, NC89]. neuron [Ano96c]. Never [TL17, Tho19]. Neves [Gre92]. NEWFONT [McC87]. newly [Ano96c]. News [GS21, Wuj89]. Newspaper [IKM82]. newspapers [Bur18]. next [Ano92h, PBC91]. NFSS [Raf93]. Nicole [Dur22]. Nightmare [Sch92b]. Ninth [IEE88]. No [Joc90, Ano93b]. no-font-hassle [Ano93b]. no-fuss [Joc90]. No-muss [Joc90]. Noblest [KB16]. Noise [BB89]. Non [DMK69, HDK68, Pie91, WT92, Wil04, WT06a, Wil06b, WT06b]. Non-Alphabetic [DMK69]. non-designer [WT06a, Wil06b, WT06b]. non-designers [Wil04]. Non-impact [Pie91]. Non-Rotated [WT92]. Non-Standard [HDK68]. Nonfiction [PRB87]. Nonimpact [Tre94]. Normes [Dar89]. notation [MVM09b]. Note [Pos89d, Pos89c, Pos91a, Pos91c, Pos91b, Ado92, Bec91b]. Notebook [Gla04]. Noted [Ano94d, Ano95a]. Notes [BH91, TS85]. Nottingham [Van86]. Nov [Ano93a, Rom16]. November [DC93, IEE88]. novice [Wil04]. Number [USK^{+}17, HYH89, HYH91]. numeral [MVM09b]. Numerals [MVM09b, MVM09a]. Nutritional [Gib90]. Nwanua [Ano98]. NX [Sta87, Sta89]. NY [Ano94r].

obituary [Bur18]. Object

[Sch98b, Sta93b, Van93a, Van93b]. Object-orientation [Sta93b]. Object-Oriented [Sch98b]. objects [Bro91]. Observer [VK93]. Obtention [Mac91a]. occidental [WH76]. OCLC [Tob84]. OCR [ACG94, But94, CJ83, CDV95, Van93a, PC94, RH75, SV73, Van93b]. OCR-A [SV73]. Oct [MA91]. October [AH89, Anoxx, Ara95, IEE86, IEE93, Pie91, Knu89]. offer [Emm95]. offers [Ano92f, Orr94]. Office [Cat83]. OH [JN93]. old [AJ92]. Omni [KB86, Wol86, Wol87]. Omni-Font [KB86, Wol86, Wol87]. on-line [Pou95]. one [Ano91e, Har96]. Only [EM84]. onscreen [Sha07]. open [Woo87]. OpenType [Hos24]. operations [Ano91a]. Operator [AMS83, Pos91b]. Oppositions [Ano94u]. Optical [Tri69, ZI93, ZI95, AV94, AIDG95, ELS94]. Optimal [FK82, FK85, Bow68]. optimization [Kok94]. optimizations [Pos91a]. Optimizing [PZL94]. option [RNR94]. Options [Ano90b, Ano91d, Pat94]. organizing [MRT89, WW93]. Oriental [ICC88, TL92a, LK95]. orientation [MRT89, Sta93b]. Oriented [Sch98b]. Original [Eng87]. Originals [Rig14, Rig14].

Ornamental [Ost98]. ornate [Ano94i]. Orthographies [DMK69]. other [Lan40]. Our [CKT80, Wil93, Cus11]. Out-of-Focused [Yi92]. Outline [And90, CM94, CP91a, Dea92a, HZ93, KKM^{+}93, KKM^{+}94, RSSW89, Sch00, SR96, Ano96a, Hug88, KMNO89, MP94, MPZ95, NKMO89, PM93, PB03, SR97, Sta93a, Hen89]. Outline-Based [Sch00]. outlined [SZ95]. outlines [Hob93, HP93, SR02]. Output [Ano98, Bro85, DS88, RAK83]. Outrageous [EA87]. overly [Mid95]. Overview [HM78, Kom91]. own [WT06b].

P [MAYB02, Wal05]. P-13 [MAYB02]. PA [IEE93]. Pacific [SK93, Egl90]. Pack [Ano93c, Ano93m, Ano93n, Ate92, Bit91, Hav93, Mic92b, Mic92a, Tim87, Tim88]. Package [RG93, Ano92e, Jef93a]. packages [Ano92a, Ano92h, Emm95]. packages— [MS89]. Packard [Ano93n]. Packed [Rok85]. Packs [Ano92l]. page [Van93a, Phe93, Phe94, Van93b]. pages [EA87]. pair [Ano92h]. pak [Ano94p, Ano95c]. Palatino [Bri16a, Bri16b]. Palladino [Fox16]. Palm [Egl90]. pamphlets [Gen94, Phi94]. Panel [AMS83]. PANOSE [Bau88, BL94]. paperback [Tho19]. Paperless [Ano93b]. Papers [MA91]. Paragraphs [von92a, von92b, von93]. Parameter [HHGH98]. Parameterizable [HH01]. Parameters [Moh98]. Parametric [DL91, GM83a, MB93]. Parametrization [Har93]. Paris [IEE86]. Parlor [Dur22]. Part [Adl91, MFW23, MFB^{+}23]. partial [Les95, MG95]. party [Ano95a]. Pascal [Kay85]. passion [Web15a]. path [Van93a, Van93b]. path-based [Van93a, Van93b]. Pattern [FN85a, FN85b, IEE88, CVP86, IEE86, IEE92, KB86]. Patterns [Ano86c, Ost98, PS84]. PC [Ado91a, Ano88a, Ano88e, Pap89a, Pap89b]. PCs [Ano94v, Kay96]. PDF [PB03]. peaceful [Ano91e]. Pen [Pan16]. People [RBY16, IEE93, Kay85]. People/Pascal [Kay85]. peoples [Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00, Fau04a, Fau04b]. Perceived [SB93]. Perception [KB86, Sha07]. Perceptual [RRW93, San92]. Perceptually [HBBG95, MHC98]. Perceptually-Tuned [MHC98]. Performance [AMS83, LB86, LCO94, Mey94, VK93, Ano95b, Ike95]. performs [Mid95]. Personal [Ano93p, Big85, KG77, Big86b, DW90]. PersonalFont [Ano92i]. personality [Sha07]. Persuasion [Jan90]. PHAROS [CP91a]. Phase [DM89, EEN90]. Philadelphia [IEE93]. Phoenix [Sta85]. photoscript [EEN90]. Phototypesetter [CKT80]. Picard [Sha94]. Picture [Buy85]. PIM [Swz93]. PIN.#xx [Ano92j]. pinyin [Won93]. Pitch [PZL94]. pixel [UBR90]. Pixels [Mac91b, Mac91a]. PK [Rok85]. Placement [GM83b]. Plasma [AMS83]. platform [Bro94]. Plotters [DH83]. Plotting [BU78]. plurality [ZZ96]. plus [ALS88, Alt86b, Cen88, Mic87a, Mic87b, Mic87c]. Poetica [Hoe95]. Point [Dre85, Hal07, Mid95]. Points [Ano86a]. police [And89a]. Polices [CP91b, Nic91]. polishes [Nic91]. polit [Nic91]. Polytechnique [AH89]. popular [Jac82]. Portable [WB95]. portfolio [Car10]. Portland [Pie91]. Positional [MVM09b]. Possibilities [Dea92c]. possible [Kla93]. post [Emm95]. post-production [Emm95]. posting [WT06b]. Postprocessing [RH75]. PostScript [AO89, AB91, Ado89, Pos89d, Pos91c, AO89, AB91, Ano90a, Ano90c, Cra90, Egl90, Fel91, GMP90, GMP92b, Har93, Hug88, Joc90, Per88, Pow90]. pour [Lou91, Mac91a, TL91]. Power [Bro90]. powerful [Orr94]. Practice [von92a, von92b, von93]. pratique [Des99]. precise [Ano92f]. Precision [Ano95a, LNN95, LNN00, Web15a]. predicting [Ada91]. preference [Seb91]. Preliminary [Due93b, Dur93b, GW91, Sym81]. Prepaging [FK85]. Preparation [AKWC82, Her70]. Presentation [II88]. Press [Ano85h, Ano85i, Tho19, Dur22]. Pretty [Gai85]. previewing [Ano91f]. prices [Mac35]. primer [Dea92a, Ste14, Ste18, Wil94]. priming [Ten90]. Princeton [Tho19]. Principles [Big85, Big86b, Gib90, BS82a, BS82b, Wil04]. Print [CFW82, Wen67, Ano92b, NNMS94]. Printed [AM86, CCH88, KPB87a, KPB87b, LZ86, Ume82, UM85, WT92, Ano96b, Ben93a, Lin95]. Printer [BS84, BG73, BM87, Cat83, DiV84, Mac91b, Par88, Sou05, Ano85g, Ano85f, Ano85h, Ano85i, Ano85j, Ano95d, Bus92a, Bus92b, Cra90, Hew88, Joc90, May86, McC87, Nas94, Pet86, Sta87]. Printers [BHW83, FN85a, FN85b, HH83, Tre94, Ano90a, BS82a, BS82b, DW90, Egl90, Hew85, Hug88, Mac91a, ZL87]. Printing [Adl91, Ano86b, Ano86a, Ano95d, BS84, BK88, FLS78, Fow73, MN77, Pie91, Upd22, Upd37, Upd62, Upd80, War09, Xer79, Xer80, Bil87, Dea85, Hug88, LS91, NW83, Ste14, Ste18, Xer82a, Xer82b, Les95]. printing-types [Dea85]. Prints [Per88]. Printwheel [May85]. priori [ZI95]. pro [Way92b, Way92c, Ano93i, Van92b]. problem [Bus92a, Bus92b, Mid95, ZZ96]. Problems [AS89, Ste87, Lan40, MRT89, Ueh90, Mal95]. Proceedings [AH89, AGS93b, USE87, Van86, AOM95, DC93, IEE93, SK93, AGS93a, Dés86, Fur90, CVP86, IEE86, Luc85, BDV95]. processes [IEE93]. Processing [Cam92, Hud00, ICC88, IBC87, KKM^{+}94, AJ92, Big86a, BDV95, DC93, GV92, HHG97, KKM^{+}93, MCP91, San92, TL92a, Van86, Woo89]. Processor [CJ83, KKM^{+}94, Ike95, KKM^{+}93]. Processor-Based [CJ83]. Processors [Adl94]. ProCollection [Hew89]. produces [Rul87]. Producing [Gon98, SD86]. Product [Gre96, Pap89b]. Production [GM85, Ben93a, DP93, Emm95, GSKM92]. productions [Zap00]. productivity [Ano93d, Ano94b]. Products [Ano98, IEE93]. Professional [Won85b, VCA88, VCG^{+}89, VCGA90, Wil06b, Won85a]. professional-level [Wil06b]. profiles [ZI93]. Program [Ano86i, Knu86b, Ano92b, Ano93u, Ano94v, Ber96b, McC87, Ols95, Ros89]. Programmable [Cow87]. Programmed [Ano86b]. Programming [Her69, Kay85, GW91, MP92]. Programs [DH83, Pos91b, Ano94s, KNT^{+}01]. Progress [FT96, Wen67]. Project [FG85, Sie85, Phe93, Sou98]. projection [ZI93]. Prolog [Due93b, Due94, Dur93b]. promise [Ros89]. proofs [Ano84a]. propagataion [HYH89]. propagation [HYH91]. property [McC07]. Prosthesis [GL84, GL83]. Protection [Got85]. Providing [Ano85d]. Provisional [GM83b]. PS [Rah93, Wre87, Wre88a, Wre88b]. pseudo [Coh93]. pseudo-dynamic [Coh93]. PSI [TS85]. Psychology [Sha07]. Public [Flo94, PRB87]. Publish [Tim92]. Publishing [Ano93b, Ano94f, Fur90, KM89, Mar87, RM87, Yuz89]. Punch [USK^{+}17, Kar91]. Punchcutting [Dro85]. Punk [AO89, Knu88b, Knu88a]. Pure [Ano98]. purpose [KNT^{+}01]. Put [Tre94].

QMASTER [Ano98]. Quadratic [SV73]. quadratischer [SV73]. qualité [Mac91a]. Quality [AYO91, AS89, BS82a, BS82b, Buy85, DiV84, Mac91b, ST85a, ST85b, SR97, UKKS84, Ada91, Ano96a, Lin91, MP94, Mac91a]. quantizer [Ano96c]. Quattro [Van92b]. que [And89a]. questions [Ano91c]. Quickdraw [Ano94d]. Quietwriter [Ano86h]. quill [Tho90].

R6 [Mey94]. RAMTEK [Chi84]. Random [DM95]. Raster [AGS93a, AGS93b, MA91, Cro78b, Cro78a, AH89]. Rasterization [And90, Her89, Dea92a]. Rasterizing [HP93, KMNO89, NKMO89]. ratio [Som90]. Re [Ash90]. Read [EM84, F.10]. Read-Only [EM84]. Readability [RBY16, Wen67, Seb91, SB93, UBR90]. Reader [UM85]. readers [Kir05]. reading [BSS92, DeC83, MAYB02]. Reads [F.10]. ready [Bee85a, Bee85b]. real [Dub90, Orr94]. real-time [Orr94]. reality [Ano91h]. ReaySetGo [Ano94t]. Recognition [AM86, Ano86e, CKM86, CK86, CW85, CCH88, CBE74, HD88a, IEE88, CVP86, IEE86, IEE92, KPB87a, KB86, KPB87b, KH96, LZ86, Rei66, Ste87, Tri69, Ume82, Wol86, Wol87, All94, Ano96b, Ano96c, AIDG95, Bow68, EEN90, ELS94, HD88b, Woo87, Ken86, KD94, Krt94, cYL88, LO90, Lin95, LCO94, VP94b, WW93, ZI93, VP94a]. Recognizers [LB86]. recognizing [FT96]. Rectangular [NF87]. reducing [Dea85]. redux [Ano95a, Hoe94]. Reference [The03, LNN95, LNN00, Mac90]. regard [Mac91a]. regarding [Knu96, Sha07]. regularisation [Sta93a]. rejecting [WW93]. Relation [CCH88]. Relational [Lin95]. relationships [Ano95b]. Relativity [Gei15]. release [Ano91a, Kna95, Kna95]. Released [Gei15, Ano94v]. releases [Ano92e, Van93a, Van93b]. Remembered [Web15a]. Remembering [Hol15, Rom16]. Remo [BDV95]. Rendering [Cor00, CP91a, KKM^{+}94, Mey94, KKM^{+}93]. Repeat [Ano97b]. repertory [WH76]. repetition [Ten90]. Reply [Gre92]. report [GW91]. Reports [Her70]. Representation [BBK^{+}19, Gos86, Kar92, LB90, Tra89, WT92, Hus95, Kla93, Lin95]. Representations [SR96]. Reprint [MS90a]. Requirements [Ste87]. Research [Ano98, GM85, Kar98, Wen67, Wen68, Ara95, Phe93]. Resistive [Ano86a]. Resolution [BHW83, Koh89, NN89, BS82a, BS82b, Par90]. resources [GC88]. Rest [Hos24]. Restoration [BBK13a, BBK13b, Yi92]. Retaining [But94]. Retrospective [Kar98, Cra89]. Rev. [Fox16]. Reverse [Kad11]. Review [Dur22, Flo94, McC16, Tho19, Ano89e]. Reviewed [Sha94]. Reviews [Ano93s, Ano93q, Ano93r, Ano94w, Pap89b, Par93c]. revised [Bee85b]. Revisiting [BBK13a, BBK13b]. Revival [Car85, Ano95a]. Revivals [Kel19b]. Rhetoricity [Dur22, WD18]. Ribbon [Ano86a]. Rich [TL92b]. RIDT [MA91]. Right [Ano93e, Fen94]. RISC [Ano90a, KKM^{+}93, KKM^{+}94]. RISC-based [Ano90a]. RIT [Pan16]. road [Bur18]. Robert [Fox16, Car85]. Rochester [Ano94r, Car10, Lov24]. Rockies [Sou98]. rods [Ano90a]. Rogers [KB16]. ROM [Ano93e, BMU94, BMU95, Hav93]. Roman [Ano86g, KB16, MVM09b, Sol83, Tob84]. Romanization [MVM09a]. Rome [IEE88]. rood [TS94]. Room [Cat83, Ano91e]. Rotated [WT92]. Rough [Ano91h, Egl91]. rounded [Ano94v]. Row [Ano85e]. RSVP [MAYB02]. Rules [RRW93, Mac78]. run [Dés84]. Russian [ZZ96]. Russians [Ano94v].

Safe [Ast72]. same [Kan84]. sample [War80]. sampler [Hyp90]. San [BDV95, Uni00, VP94b]. Sans [Che23, SA70, Dam16, Mid95]. Santa [Ano94d]. Save [Mid95]. Saveas [Mid95]. Scalable [Con98, Ike95, Lee94, IWM94]. Scale [AYO91, Cro78a]. scaler [Sta93b]. Scaling [Ano90c, CFW82, HWY92, Kar89, Par90, Wad84, Won85b, AV94, NSA90, Won85a]. Scan [Ano88e]. scanned [CWB93]. scanner [Ano94t]. Scans [Ano88e]. Scenes [SFO88]. Scheme [Pon89, Dea85, MS89]. schemes [Ste14, Ste18]. Schrift [Bar80, Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00]. Schriftart [SV73]. Schriften [Kar92, KW93]. Schriftherstellungen [Zap00]. Schriftzeichen [Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00, Zap00, Fau04a, Fau04b]. Science [BH18, Rie89, Dea85]. Scientific [Dés86, GV92, Her72, Luc85, Zap85]. scientifica [Bec91b]. Scott [Dur22, Ano97b]. Scottsdale [Ano93a]. SCREADER [Ano98]. Screen [RBY16, Sch92a, Pos89a, Ano93b, Cyr88]. Scribe [Fox16]. Script [Har98, Kim90, Kok94]. Scripts [Fau85, Fau86a, Fau86b, Fau00, KW93, LC98, Moh98, RAK83, Fau78, Fau80]. Seattle [Mor93]. Second [Dés86, MA91]. Secrets [BG99, Jon94]. section [SR02]. see [Ano92b]. Seeing [Lov24]. Segment [Ast72]. Segmentation [KD94, WT92, Ano96c, WW93]. Segmentation-free [KD94]. Select [Swz93]. Selection [Ano86i, Dun92, MS89, NW83]. selection— [MS90b, MS90a]. Selections [RRW93]. Self [BN94, MRT89, WW93]. Self-correcting [BN94]. self-organizing [MRT89, WW93]. Semantic [Wen68]. Sense [Ano90f]. sensitive [Mid95]. Seoul [SK93]. Sep [SK93]. Separating [HZ93]. September [BDV95, Fur90, ICC88, IEE92, Uni00]. Sequences [RN83]. Serif [Ano24, SA70, Mid95]. Serifs [MAYB02]. Server [CM94, DM92, NNMS94]. Servers [Pea92, TL91]. Serveurs [TL91]. ses [Nic91]. Set [Ano86b, Ano93n, Che94, Jap78, KM87, WW82, Wre87, Wre88a, Wre88b, ZZ96]. sets [NNMS94]. setting [Ano95d]. Seven [Ast72]. Seven-Segment [Ast72]. Seventeenth [Rie89, Uni00]. Seventh [Pie91]. Seybold [Ano93k]. Seymour [Sey94]. Shape [HH01, Par93a]. Shaped [Fox16]. shapes [HH94, PMZS97]. shaping [MNSJ99]. share [Ano93k]. Sheep [SG02, SG03]. ship [Ano92h]. Shoot [Ano91g]. Shoot-out [Ano91g]. Shopping [Ano91b]. Should [War09]. Shoulder [Wil93]. show [Ano93k, Ano94d]. showings [Cen88]. shows [Ros89]. SIGGRAPH [Ano89e]. Signal [DC93]. signs [Bur18, Lov24]. Silenced [Ano92j]. Silicon [CP91a, McC16]. Silver [Rig14]. Simple [Koh89, Ano91a, Dea85]. Simplification [Har98]. simplifies [Ber96b]. Simulating [LC98]. Simulation [MS84, DM95]. Singapore [Ano94e, Ano94q]. single [Bow68]. Singular [Ano97a]. sistema [Bec91a]. site [WT06b]. Sixth [AOM95]. Size [KPB87a, KPB87b, LB86, Wal05, DeC83, HD88b, Wil80]. sized [AV94]. Sizeinvariant [HD88a]. sizes [BSS92, Mac35, MAYB02, SB93, Ste14, Ste18]. skeletal [HHG97]. Skeleton [BS96, Gon98]. Skeleton-based [BS96]. slant [Ken86]. Slavic [Ano87c]. slide [Ano94t]. slow [MAYB02]. slugs [Mac78]. small [MAYB02]. Smart [Ano91b, Hei95]. SmartSuite [Ano95e, Lot93]. smoothness [Kok94]. Society [CVP86]. Soft [WW82]. Soft-Edged [WW82]. SoftUnion [Ano94v]. Software [Ano88e, Ano95f, Ano98, CW85, Dea93a, GL84, Lat95, NNMS94, SD86, Ano90g, Ano93d, Bus92a, Bus92b, GL83, McC87, Win90]. Solid [Ano98]. Solutions [Ano98, Jon94]. solver [Bus92a, Bus92b]. Some [Tam89, BS96, Van88]. something [Emm95]. son [Lou91]. Sonata [Pos89c]. Sophia [Ano87d]. Sophia-Antipolis [Ano87d]. sorts [Mac78]. Source [Wil91]. SP [TA82]. SPEC [Whi87]. Special [AGS93a, AGS93b, Dea93a, Mac90, Xer80]. specific [Phe93, Phe94, San92, SO94]. Specification [WB95, Pos89a, Pos89b, Pos89c, Pos90, Pos92a]. specimen [Car10]. Specimens [MC93]. Spectra [DM89]. Spectral [VK93]. speech [Ano96b]. Speed [KKM^{+}94, Ano93f, KMNO89, KKM^{+}93, ZL87]. speedy [Ano92e]. SpellCaster [Ano98]. spends [Sey94]. Spent [CKT80]. spite [UBR90]. spline [LH91]. Splines [KM83, LS84, Ram87, Kla93]. spotting [CWB93]. Spreadsheet [Ano85e]. Spreadsheets [Ano85d]. SprintScan [Ano94t]. Square [ZST84]. Square-Styled [ZST84]. squares [Kop93]. stability [Kna96]. Stack [Ano88f]. Standard [Flo94, Fow73, HDK68, Hos24, Jae91, SP87, Tri69, Ano91e, MS90b, MS90a, SBSG89, Van88, Xer82b]. standardization [Kom91, Kom93]. Standards [Hos24, Bie90, Dar89, Ano92k]. Stanford [McC19, McC20, Sie85]. Start [TBLJ90, Zapxx]. StarTech [Ano98]. State [Got85, Dea85]. Stealing [SG02, SG03]. Stempelschnitt [Zap00]. stemsnap [Pos91b]. Step [Har96]. stile [Bec91b]. Stofte [Ano93t]. Stonecuttings [Lop85]. Stop [SG02, SG03]. Storage [Won85b, Won85a]. Store [EM84]. Stories [Hal07, Zap07a, Zap07b]. Story [Rig14, Zapxx]. Storyboard [Kur93]. Strasbourg [Dés86]. Strategies [Bro94]. strategy [Ano91e]. Street [Sch92b]. Stroke [AYO91, BBK^{+}19, CCH88, HZ93, Sch00, Ano96a, Lam80, LK95, MPZ95, Sch98a, TC92]. Stroke-Based [Sch00, BBK^{+}19, Sch98a]. Stroked [Wil91]. Strokes [TKE83]. structural [CDV95]. Structure [NN89, PM93, cYL88]. Structured [Ano96c, Big85, Cow87, Due93b, Due94, Dur93b, LK95]. Structures [AKWC82, Tra89]. students [Pan16]. Studies [Moh98]. Study [PRB87, Upd22, Upd37, Upd62, Upd80]. Style [Sla09, WH86, Lin91, RNR94, Tut92]. style-option [RNR94]. Styled [ZST84]. Styles [AMS83, SB93]. Subclasses [RN83]. Subdivision [Kla94]. Subiaco [Ano95a]. subjective [Seb91]. Substituting [PB03]. Substitution [McC84]. Subsystem [BS84, FLS78, NW83]. suggested [Chi88, Sch88]. summary [Jef93b]. Summer [BBK13a, BBK13b, CKT80]. Sun [Ano97b]. Super [Ano93u]. SuperATM [Ano93v]. SuperGlue [Sol87b, Sol88]. superhighway [Ano95a]. supplement [Bar80]. Suppliers [Ano95f]. Support [CM94, PZL94, Jef94, MCL92]. Supporting [Pos89d, Pos91c, CP91a, Kim90]. supports [Cyr88]. survey [Wuj88]. Survivals [Upd22, Upd37, Upd62, Upd80]. Swiss [Van92a]. switching [Pos91a]. Switzerland [AH89]. Symbiosis [AOM95]. Symbol [Ano86b]. symbology [UB95]. Symbols [Gen87a, Gen87b, WT92, Bee85a, Bee85b, WH76, Kna95]. symposium [JN93]. Syntactic [Wol86, Wol87]. Syntax [SA70]. Syntax-Antiqua [SA70]. Synthesis [TKE83]. System [BL94, Car84, CJ83, CY92, CM94, CC83, Chu84, Dea92c, DMK69, Her70, Her72, IBC87, KM89, May85, MB93, MN77, MS84, Nel85a, RH75, SZC94, UKKS84, Wol86, Wol87, dGH89, Adi90, Ao90, Ano95b, Ano96a, Ano96b, Bau88, Bow68, GW91, HFW^{+}96, HH94, Int82, Int83, Woo87, KC75, KKL89, PMZS97, Par93b, Ris88, Sch98a, Xer82a, Xer82b, Ano91c, Pou89, WCW82, Xer79, Xer80, ZI95]. Systems [Ano87d, Ano98, Col86, Cor00, GM83a, Mar87, MCL92, Rug83, SFO88, Ste87, Ano91f, HM78, CS87].

table [Bee82, Har92]. tables [WH76]. Tale [Kur93]. talks [Ano97b]. teacher [Ano93d, Ano94b, Pan16]. Technical [Pos89d, Pos89c, Pos91a, Pos91c, Pos91b, Ado92, SR02, Zap07a, Zap07b, IEE93, Zap00]. technique [Par93a]. Techniques [CKM86, CK86, CY92, LCO94, Wil06b, WH76]. technischen [Zap00]. technischer [Zap07c]. Technological [GM85]. Technologies [Pie91, dGH89]. Technology [Ano94w, Ano98, Ben93a, BS81, BM87, Kar94b, Lee94, Mey94, Ano94c, Dea92a, Kar94c, KMU94, Ueh90, Wil94, Car10]. Telecommunications [Ano98]. Teleprinter [TA82]. Tendencies [Zap85]. Terminal [GL84, SR80, Chi84, GL83, HM78]. tesi [Bec91b]. test [Knu86e]. Testing [LB86, McC83]. TEX [Dés84]. Texas [McC87, Pet86]. TEXnicolor [Ano90e]. TEXnology [Pap89a, Pap89b]. Text [Ano88e, Ano91h, BDN83, BBG85, CW85, Hos24, Kur93, LB90, MCP91, PRB87, Van86, Ben93b, Big86a, CGP95, Egl91, EEN90, GV92, Kna95, Niz90, Seb91]. Texts [AM86, MNSJ99, Ste92]. Thai [Str93a]. thanks [Mac91a]. Their [Che23, MS90c, Upd22, Upd37, Upd62, Upd80, Har98, Pan16, Ste14, Ste18]. Theory [Gei15, Kre90, Ung18]. there [Ano91e]. these [Ano91c]. TheWorldTransliterator [Rei93b]. Thinking [McC19]. Three [Hal07, Har98, Pon89]. TI [Pet86]. tighter [Van93a, Van93b]. TIGS [HM78]. Time [Ano93e, Orr94, Sey94]. Times [YWL89, DeC83, Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00, Fau04a, Fau04b, Har96]. timesaving [Ano91f]. Tips [Ano91b, Gra83]. titling [Emm95]. TMS34010 [Tex86]. Today [Mum89, Emm95, Mur90]. together [Van93a, Van93b]. Tokyo [AOM95]. Tomorrow [Mum89]. too [McC07]. Tool [RRW93, Str93b, TX92, Rul87, Sey94, UB95]. Tools [BK88, Kar94b, Ost98, SD86, Ano90g, Kar94c]. Top [Ano90g, Yuz89]. Topography [MA91]. Toronto [Ano95a, ICC88]. torture [Knu86e]. Touch [HFW^{+}96]. tough [Ano93b]. Tour [Tei85]. town [Ano91e]. Tpsnew [Pow90]. Tracy [Ano95a]. tradeoffs [Som90]. tradition [BS81]. traditional [And89b]. traditionnelle [And89b]. Transliteration [Rei93b]. Transputer [Ara95, Ara95]. Transylvanian [Sta85]. Treacyfaces [Ano94d]. treasure [Hav93]. tree [Ano91f].

Trees [Ano91f]. Treiber [LW88]. Trends [Rie89]. Tri [Pre57]. Tri-Arts [Pre57]. tribute [Pan16]. tricky [Min91]. Troff [IBC87]. true [Ano91h, Ano91e, Her95]. TrueType [Ano92h, Ano92l, Ano93m, Ano93n, BG99, Bit91, Fel91, Gla90, KMU94, LL94, Mic92b, Mic92a, Rei93a, Sla09, Sta98]. Tube [Dam16]. TUG [Doh81]. Tuned [HBBG95, MHC98, Hob93]. tuning [Bil87]. Turbocharged [Ano90a, Cra90]. Turku [MVM09b]. Turning [Dre85]. turns [Ano90a]. Tutorial [Tur03]. TV [Cha83]. twentieth [Sou05]. Two [Her79, Kar98, Bar95]. Twombly [Sto16]. Type [Ano92m, Ano95a, BH02, Ber02, BH25, CGP95, Cen88, Che22, Con98, CBE74, Dre85, DM89, Emm95, Goo95, Gra83, Gre95, GMP90, GMP92a, GM85, Kay96, Kel19b, Les95, LNN95, LNN00, Mal95, MG95, McC16, Moh98, Pou12, RBY16, SG02, Ste14, Ste18, Sto16, TCLM82, Pre57, Tri69, Ung18, Wae92, Wag85, Whi87, WD18, Zap85, Ado87, Ado90a, Ado90b, Pos91b, Pos92b, Ado92, Ado93, BS81, BS82a, BS82b, Big86b, Bus92a, Bus92b, Car10, Cus11, Fon92, Gar10, Gen87a, Gen87b, Gho84, GMP92b, K^{+}02, KGR03, KGR07, Lan94, Par93b, Seb91, Sha07, Sou05, Whi27, Whi47, Whi91, Wil06b, WH76, Ano91e, Dur22, Gla90, Moy95]. Type-1 [MG95]. Type1 [Mal94]. TypeAlign [Ado90c]. Typecase [Ano93w, Ano93x]. Typecasting [Ano94d]. Typeface [Ano88f, Bri16b, Bur18, Car84, Got85, Sch98b, Ano93l, Bri16a, CSS93, Dam16, Fon92, Sch98a, Sha07, Sou86, WH95, BL94]. Typeface-Matching [BL94]. Typefaces [Ano93y, Ano24, Ber96a, Che23, Ear02, Gab93, Knu86d, MC93, Sou85, Wal13, Wen68, BS96, Kar94a, Sha03]. Typemaking [Sou98]. Types [Car85, KB16, Sta85, Upd22, Upd37, Upd62, Upd80, Big86a, Dea85, Kel11, Ste14, Ste18, Way92c]. typeset [FT96, Pow90]. Typesetter [BBK13a, BBK13b]. Typesetting [AKWC82, Col86, GSMN93, Har94, How86, Knu79, Pow90, Str93a, Con89, Gho84, KC75, Knu89, Tut92, WH76, Won93, HH83]. Typewriter [May85, Har92]. typewriting [Kan84]. typewritten [Kan84]. Typing [Ano93y]. Typografie [Poh15, SG03]. Typographic [Dea92c, Her70, HHGH98, Kar98, McV10, MC93, SR96, GC88, HH94, Wil04]. typographics [Man93]. Typographie [Rud88, And89b]. Typography [AH89, AGS93a, AGS93b, BH02, Fur90, Gre95, Her69, Her72, Her81, Her95, LS84, McC20, McC91, PBC91, And89b, Bau88, DD15, Emm95, Dés84]. typophiles [Lan94]. Tzu [Jon95].

ultimative [Poh15]. underlying [Van93a, Van93b]. Underpinnings [McV10]. Unger [Bur18]. Unicode [Che94, Uni00, Ber02, BH93, Hos24, MVM09a, MVM09b]. unified [Kay96]. units [HYH89, HYH91]. universal [HH94]. University [Van86]. UNIX [BBG85, IBC87, Ano91f, Ano97b]. update [Ano92f, Bee81, Whi81]. Updated [Sau87]. Updates [Ado92]. upgrade [Van93a, Van93b]. Upon [AMS83, Ken86]. Uppercase [Bee81]. URW [Kar98]. US$24.95 [Tho19]. US$50 [Gab93]. USA [USE87, USE89, Pie91]. Use [Ano24, Che23, DB93, HP98, IBC87, Upd22, Upd37, Upd62, Upd80, Bus92a, Bus92b, Cro78b, Cro78a, Par93a, RNR94, TL17, Tho19, Jac82]. useful [Van88]. User [Ado91a, Ado91b, MS90b, MS90a, Mic92b, Tex86, Xer79]. Users [Ara95, PRB87, Ano92b, Sta87, Sta89, Xer82a, Xer82b]. Uses [Cla90, II88]. Using [AYO91, ACG94, Bro91, CW85, CY92, Con98, DM92, Due93a, Eng87, Fow73, KB86, KM83, SV73, SR98, SD86, Wae92, WW93, Wri83, Ano96b, Ano96c, AIDG95, CWB93, Con89, CDV95, Dür93a, Gun85, How86, KH94, MP94, Par90, Pow90, Rah92, Rah93, SO94, Tut92, War80, Won93, WH95, Goo95, Lev88]. Utilities [MO92, Ano91f, Ano92a, Ano93k, Hol87, MG95]. Utility [DH83, Ano94k, Wre87, Wre88a, Wre88b].

V2.10 [Ano98]. Vacation

[BBK13a, BBK13b, CKT80]. Valley

[McC16]. Variable

[Ano85k, Kla93, LB86, Rei66, AV94, Cyr88]. Variable-Size [LB86]. variable-sized [AV94]. Variable-Width [Ano85k]. Variation [BU78]. variations [Van88]. varied [SB93]. varies [HYH89, HYH91]. Variety [Ate92, HFW^{+}96]. various [Knu86a]. Vector [BU78, WT92]. vectors [Cro78b, Cro78a]. Velocity [LC98]. Venice [Ano85f]. Vent [Ham90]. Vento [TP90a]. Ventura [Ano93f]. Veränderungen [Zap00]. Vergleich [SV73]. Version [Pos89a, Pos89b, Pos90, Pos92a, Flo94, LNN95, LNN00, Tur00b, Ado91a, Ado91b, MS91, Pad92, Sic85, SA86, Sym81]. versus [Fel91]. very [MAYB02]. VFComb [Ber96b]. VGA [Cha88, Gag93]. VI [Adl91]. Video [Ano98, RS81, Ano89e, Tho90]. viewed [Wil80]. Viewer [Gag93]. Views [Par93c]. Vinyas [Par93b]. Virtual [Ash90, GL84, Knu90a, Ber96b, GL83, Rah92, Won93, Zla91, DP93, Knu90b]. virtueller [Glö92]. Vision [CVP86]. Visual [McC19, Tra89, BL94, KH94, Wil04, Sta98]. Visually [GL84, GL83]. vol [SK93]. Völker [Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00, Fau04a, Fau04b]. Volumes [AKWC82]. Vom [Zap00]. vs [Ano94u, Gab93]. VT220 [Sla09, Wil06a].

W3Control [Ano98]. WA [Mor93]. Walter [Ano95a]. warning [SB93]. warnings [BSS92]. Wars [Ano90d, Gla90]. was [Web15a]. Watch [Gre96]. wave [PBC91]. Way [Koh89, RN83, Swz93]. Web [Uni00, WT06b]. Wechel [Ano85j]. week [Ano92h, Ano93k]. weeks [Bar95]. well [Ano94v]. well-rounded [Ano94v]. West [Kin05]. WhatTheFont [Anoxx]. Wheel [BG73]. Where [Hor93, Kel19b]. which [Ano91e, Ber96b, Cyr88, Emm95, HYH89, HYH91]. Whirligig [Ano94d]. white [Bil87]. Who [Fox16]. Whose [Web15b, Bur18]. Width [Ano85e, Ano85k, Kla93]. widths [Cyr88, McC83]. wild [Ado93]. will [Ano92h, Ano93k, Ano94c, Ano97b]. Windatlas [TP90c]. Window [CM94, HHGH98, Liv95, Pou89, Woo89]. Window-Based [HHGH98]. Windows [Ano94d, Gai85, Tim92, Adl91, AH92, Ano91h, Ano92c, Ano94c, Ano94k, Ano94s, Ano94v, Ano95d, Dea93b, Dun92, Egl91, Hud00, Kay96, Lot93, Luh92, Mic92a, Min91, MP92, Pad92, Swz93]. Wire [BM87]. Wisdom [Ano92d]. Within [Ano85d, Dea85]. Wizards [Ash90, Knu90a, Knu90b]. Won’t [Ano97b]. wood [Car10]. Word [Adl94, Cam92, CWB93, Van93a, IBC87, KH96, Van93b, Web15a, Woo89, Ano91c]. Word-Processing [IBC87]. WordPerfect [Ano92e]. WordStar [Mic87a, Mic87b, Mic87c]. work [Ano95d, Bur18, Zap01]. Working [Hei95]. Works [Ano90c, Ano93g, Ano94k, Bix92, SG02, Swz93]. Workshop [DC93, KC87, Lop85, USE87, Hol87, Jef93b, Kna86, USE89, Ano87d]. Workstation [Big85, SD86, Wri83, Big86b, Gho84]. Workstations [KM87, Won85b, Won85a]. World [Che22, F.10, Hos24, Wal13, Kel19a, Les15, Str93a]. Worldwide [Che94]. Write [Gei15, Bil87]. Write-white [Bil87]. writers [Kir05]. Writing [Cor00]. Written [ZST84, Jef93a, Zap00]. Wrong [HN64]. Wyatt [Dur22]. WYSIWYG [Ano91h, GW91].

[CM94, Flo94, MO92, Pad92, Pea92, Pou89]. X11R6 [Mey94]. Xbase [Ano98, Neb93].

Z [Che23]. Zapf [Hol80, Rom16, Cus11, Hal07, Hol15, Kar92, Kel11, Kel19a, Pan16, SK11, Web15b, Web15a, Zapxx]. Zebrackets [Coh93]. Zeiten [Fau78, Fau80, Fau85, Fau86a, Fau86b, Fau00, Fau04a, Fau04b]. zur [Poh15, Zap00].

References

Andre:1989:DF

[AB89] J. André and B. Borghi. Dynamic fonts. In André and Hersch [AH89], pages 198–204. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Andre:1991:MFP

[AB91] J. André and J. Bur. Métrique de fontes PostScript (English: Post-Script font metrics). Cahiers GUTenberg, 8:29–50, mars 1991. ISSN 1140-9304.

Agazzi:1994:UOE

[ACG94] O. E. Agazzi, K. W. Church, and W. A. Gale. Using OCR and equalization to downsample documents. Proceedings — International Conference on Pattern Recognition, 2: 305–309, 1994. CODEN PICREG. ISSN 1051-4651. IEEE catalog number 94CH3440-5.

Adams:1986:DFL

[Ada86] Debra Anne Adams. A dialogue of forms: letter and digital font design. Thesis (m.s.v.s.), Massachusetts Institute of Technology, Department of Architecture, MIT Microreproduction Lab, 77 Massachusetts Avenue, room 014-0551, Cambridge, MA 02139, USA, 1986. 157 pp. Supervised by Muriel Cooper.

Adams:1989:AAB

[Ada89] D. Adams. abcdefg: a better constraint driven environment for font generation. In André and Hersch [AH89], pages 54–70. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Adams:1991:EFQ

[Ada91] Debra A. Adams. Evaluating font quality: a method of predicting character image defects. In Pietrowski [Pie91], pages 353–364.

Adil:1990:FMA

[Adi90] Zahir Adil. Font manipulation: an Arabic/English system. ????, ????, 1990. 97 pp.

Adler:1991:LWP

[Adl91] Marc Adler. Learning Windows part VI: Bitmaps, fonts, and printing. Microsoft Systems Journal, 6(3):115–??, May 1991. CODEN MSJOED. ISSN 0889-9932.

Adler:1994:WPF

[Adl94] Jerry Adler. Word processors: a font a day keeps my muse away. Newsweek, CXXIV(17):49–??, October 1994. ISSN 0028-9604.

Adobe:1987:FFA

[Ado87] Adobe Systems, Inc. Font & function: Adobe type library collection. Adobe Systems, Mountain View, CA, 1987. 30 pp. LCCN f Z250 .A512 Oversize items shelved as: pffZ250.A512 and pfffZ250.A512. Title supplied by cataloger. User’s manuals: Macintosh Type specimen books and sheets: Adobe acrobat; Adobe Caslon; Adobe Garamond; Adobe type library [i.e., miscellaneous type faces]; Adobe wood, Catfisch script; Charlemagne; Lithos; Mezz; Minion; Minion Cyrillic; Myriad, Nueva; Poetica; Sanvito; Tekton, Trajan; Utopia; Viva Catalogs: Font and function, Spring 1989, Summer 1989, Fall 1989.

Adobe:1989:PLC

[Ado89] Adobe Systems Inc., Mountain View, CA, USA. PostScript language composite font extensions, April 1989.

Adobe:1990:ATFa

[Ado90a] Adobe Systems, Inc. Adobe type 1 font format. Adobe Systems Incorporated, Mountain View, CA, USA, 1990. iii + 101 pp.

Adobe:1990:ATFb

[Ado90b] Adobe Systems, Inc. Adobe type 1 font format. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-57044-0, 0-471-92841-0. iii + 103 pp. LCCN QA76.73.P67 A36 1990. US$14.95, CDN$18.95.

URL http://partners.adobe.
com/asn/developer/pdfs/tn/T1_
SPEC.PDF.

Adobe:1990:AT

[Ado90c] Adobe Systems, Inc. Adobe TypeAlign, 1990. 1 computer disk.

Adobe:1991:UGI

[Ado91a] Adobe Systems, Inc. User guide: IBM PC version. Adobe type library. Adobe Systems, Mountain View, CA, USA, 1991. iv + 50 pp.

Adobe:1991:UGM

[Ado91b] Adobe Systems, Inc. User guide:

Macintosh version. Adobe type library. Adobe Systems, Mountain View, CA, USA, 1991. iii + 24 pp.

Adobe:1992:UAT

[Ado92] Adobe Developer Support, Adobe Systems Inc., Mountain View, CA, USA. Updates to the Adobe type 1 font format — Technical Note 047, February 1992.

Adobe:1993:AWT

[Ado93] Adobe Systems, Inc. Adobe wild type, 1993. 3 computer disks.

Adobe:19xx:FF

[Adoxx] Font and function, page various, 19xx. Adobe Systems Inc., Mountain View, CA, USA.

Advantage:1989:MB

[Adv89] Advantage Computing. Macintosh best 100, 1989. 1 computer disk.

Andre:1993:ESI

[AGS93a] Jacques André, Jakob Gonczarowski, and Richard Southall. Editorial: Special issue: Proceedings of the Raster Imaging and Digital Typography Conference. Electronic Publishing—Origination, Dissemination, and Design, 6(3):115– 116, September 1993. CODEN EPODEU. ISSN 0894-3982.

Andre:1993:ICR

[AGS93b] Jacques André, Jakob Gonczarowski, and Richard Southall, editors.

Editorial: Special issue: Proceedings of the Raster Imaging and Digital Typography Conference, volume 6(3) of Electronic Publishing —Origination, Dissemination, and Design. John Wiley and Sons, Inc., Chichester, UK, September 1993. CODEN EPODEU. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN ????

Andre:1989:RID

[AH89] Jacques André and Roger D. Hersch, editors. Raster Imaging and Digital Typography: Proceedings of the International Conferences, Ecole Polytechnique Fédérale, Lausanne, Switzerland, October 1989. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Angell:1992:WFB

[AH92] David Angell and Brent D. Heslop. The Windows 3.1 font book. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1992. ISBN 1-56609-005-9. x + 174 pp. LCCN ????

Avi-Itzhak:1995:HAO

[AIDG95] Hadar I. Avi-Itzhak, Thanh A. Diep, and Harry Garland. High accuracy optical character recognition using neural networks with centroid dithering. IEEE Transactions on Pattern Analysis and Machine Intelligence, 17(2):218–224, February 1995. CODEN ITPIDJ. ISSN 0162-8828.

Allport:1992:OND

[AJ92] G. Allport and P. Jarratt. The old and the new in document processing. Electronic Library, 10(1):41– 45, February 1992. CODEN EL-LIDZ. ISSN 0264-0473.

Allen:1982:CCD

[AKWC82] Frank H. Allen, Olga Kennard, David G. Watson, and Kathleen M. Crennell. Cambridge Crystallographic Data Center — 6. preparation and computer typesetting of “Molecular structures and dimensions” bibliographic volumes. J. Chem. Inf. Comput. Sci, 22(3):129– 139 (or 129–138??), August 1982. CODEN JCISD8. ISSN 0095-2338.

Aldus:1993:AH

[Ald93] Aldus Corporation. Aldus Home-Publisher, 1993. 6 computer disks + 1 ClickArt collection guide ([3] p.: ill.; 22 cm.) + 1 12 TrueType fonts for Aldus HomePublisher from Bitstream installation guide and typeface sample sheet ([2] p.; 28 cm.).

Allen:1991:CL

[All91] James S. Allen. Computational linguistics, 1991. ISSN 0362-613X. To be published quarterly by MIT Press; fonts not yet finalized.

Allam:1994:ACR

[All94] May Allam. Arabic character recognition. Proceedings of SPIE — The International Society for Optical Engineering, 2181:351–359, 1994. CODEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic).

ALSoft:1988:FJP

[ALS88] ALSoft, Inc. Font/DA juggler plus, 1988. 1 computer disk.

Altsys:1985:FMF

[Alt85] Altsys Corporation. FONTastic Macintosh font editor, 1985. 1 program file (Macintosh) on 1 computer disk.

Altsys:1986:FBF

[Alt86a] Altsys Corporation. FONTastic bitmap font editor for the Macintosh, 1986. 1 program file (Macintosh) on 1 computer disk user’s guide.

Altsys:1986:FPB

[Alt86b] Altsys Corporation. FONTastic plus: bitmap font editor, 1986. Program file (Macintosh) on 1 computer disk user’s guide.

Amin:1986:MRM

[AM86] Adnan Amin and Gerald Masini. Machine recognition of multi font printed Arabic texts. In IEEE [IEE86], pages 392–395. CODEN PICREG. ISBN 0-8186-0742-4. LCCN ???? IEEE Service Cent. Piscataway, NJ, USA.

Abramson:1983:EDE

[AMS83] Sandra R. Abramson, L. Hardy Mason, and Harry L. Snyder. Effects of display errors and font styles upon operator performance with a plasma panel. Proceedings of the Human Factors Society, 1:28–32, 1983. CODEN PHFSDQ.

Andre:1989:MQF

[And89a] J. André. Éditorial : mais que fait la police ? (English: what is the font doing?). Cahiers GUTenberg, 4:1, décembre 1989. ISSN 1140-9304.

Andre:1989:MFT

[And89b] J. André. Métriques de fontes en typographie traditionnelle (English: Font metrics in traditional typography). Cahiers GUTenberg, 4:9–22, décembre 1989. ISSN 1140-9304.

Andler:1990:AGG

[And90] Sten F. Andler. Automatic generation of gridfitting hints for rasterization of outline fonts. In Furuta [Fur90], pages 221–234. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Andre:1992:FM

[And92] Jacques André. Font metrics. Technical Report 676, IRISA, Campus de Beaulieu, F-35042 Rennes cedex, France, September 1992.

Anonymous:1984:GFM

[Ano84a] Anonymous. Gray fonts for Metafont proofs. TUGboat, 5(1):34, May 1984. ISSN 0896-3207.

Anonymous:1984:MGF

[Ano84b] Anonymous. Metafont Generic Font file format. TUGboat, 5(1):31, May 1984. ISSN 0896-3207.

Anonymous:1985:Aa

[Ano85a] Anonymous. The authors. Visible Language, XIX(1):164–166, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_
E.pdf.

Anonymous:1985:C

[Ano85b] Anonymous. Colophon. Visible Language, XIX(1):167–168, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_
E.pdf.

Anonymous:1985:GFF

[Ano85c] Anonymous. Generic font file format. TUGboat, 6(1):8–12, March 1985. ISSN 0896-3207.

Anonymous:1985:MPM

[Ano85d] Anonymous. Method for providing multi-font characters within spreadsheets. IBM Technical Disclosure Bulletin, 28(2):823–??, July 1985. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1985:MAA

[Ano85e] Anonymous. Method to automatically adjust row height and column width of a spreadsheet. IBM Technical Disclosure Bulletin, 28 (2):819–??, July 1985. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1985:PMAb

[Ano85f] Anonymous. The printer’s mark of Aldus Manutius. Venice, 1450– 1515. Visible Language, XIX (1):156, Winter 1985. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (elec-

tronic). URL https://s3-us-
west-2.amazonaws.com/visiblelanguage/
pdf/V19N1_1985_E.pdf.

Anonymous:1985:PMAa

[Ano85g] Anonymous. The printer’s mark of Antoine De Harsy. Lyon, 1550. Visible Language, XIX(1):122, Winter 1985. CODEN VSLGAO.

ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_
E.pdf.

Anonymous:1985:PMG

[Ano85h] Anonymous. The printer’s mark of The Guinta Press. Antwerp. Visible Language, XIX(1):148, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_
E.pdf.

Anonymous:1985:PMH

[Ano85i] Anonymous. The printer’s mark of The Hackiana Press. Leyden, 1671. Visible Language, XIX(1):121, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_
E.pdf.

Anonymous:1985:PMW

[Ano85j] Anonymous. The printer’s mark of Wechel. 1495–1554. Visible Language, XIX(1):54, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_
E.pdf.

Anonymous:1985:VGC

[Ano85k] Anonymous. Variable-width ghost cursor for multiple font characters. IBM Technical Disclosure Bulletin, 27(11):6510–??, April 1985. CO-DEN IBMTAA. ISSN 0018-8689.

Anonymous:1986:APR

[Ano86a] Anonymous. All-points-addressable printing with a resistive ribbon. IBM Technical Disclosure Bulletin, 29(2):609–610, July 1986. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1986:ACP

[Ano86b] Anonymous. Automatic changing of programmed symbol set according to printing font. IBM Technical Disclosure Bulletin, 29(4):1682– ??, September 1986. CODEN IBM-TAA. ISSN 0018-8689.

Anonymous:1986:CFP

[Ano86c] Anonymous. Compression/decompression of font patterns. IBM Technical Disclosure Bulletin, 28(8):3563– 3564, January 1986. CODEN IBM-TAA. ISSN 0018-8689.

Anonymous:1986:LCF

[Ano86d] Anonymous. Liquid crystal font addressing method. IBM Technical Disclosure Bulletin, 29(2):807–808, July 1986. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1986:MMC

[Ano86e] Anonymous. Media and multi-font character recognition. IBM Technical Disclosure Bulletin, 29(4):1562– ??, September 1986. CODEN IBM-TAA. ISSN 0018-8689.

Anonymous:1986:MCF

[Ano86f] Anonymous. The MF Chel Family of Fonts. The Metafoundry, OCLC Online Computer Library Center, Inc., Dublin, OH, USA, 1986.

Anonymous:1986:MRF

[Ano86g] Anonymous. The MF Roman Family of Fonts. The Metafoundry, OCLC Online Computer Library Center, Inc., Dublin, OH, USA, 1986.

Anonymous:1986:QFC

[Ano86h] Anonymous. Quietwriter font cartridge handling. IBM Technical Disclosure Bulletin, 28(11):5062– ??, April 1986. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1986:SDM

[Ano86i] Anonymous. Selection of the display mode, display font, and display buffer by initial program load command. IBM Technical Disclosure Bulletin, 29(4):1524–1525, September 1986. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1987:H

[Ano87a] Anonymous. HiJaak, 1987. 1 computer disk.

Anonymous:1987:MDF

[Ano87b] Anonymous. The MF Decorative Family of Fonts. The Metafoundry, OCLC Online Computer Library Center, Inc., Dublin, OH, USA, 1987.

Anonymous:1987:MSF

[Ano87c] Anonymous. The MF Slavic Families of Fonts. The Metafoundry, OCLC Online Computer Library Center, Inc., Dublin, OH, USA, 1987.

Anonymous:1987:WFD

[Ano87d] Anonymous. Workshop on Font Design Systems (Sophia-Antipolis, France, 18–19 May 1987). TUGboat, 8(1):77, April 1987. ISSN 0896-3207.

Anonymous:1988:CCP

[Ano88a] Anonymous. C code for the PC. Computer Languages, 5(2):52, February 1988. CODEN COLADA. ISSN 0096-0551. advertisement.

Anonymous:1988:MDDa

[Ano88b] Anonymous. Macintosh download disk 14, 1988. 1 computer disk.

Anonymous:1988:MDDb

[Ano88c] Anonymous. Macintosh download disk 16, 1988. 1 computer disk. Title supplied by cataloger. For fuller description of contents see list at Computing and Reserve Library Desk. 3dplt — Compact — DA Appl Font 1.03 — DA Artisto 1.41 — DA Blank Screen — DA Camera — DA ControlPanelPlus 0.85 — DA Coordinates — DA DeskZap 1.3 — DA Dvorak3.0+ — DA f(n)key — DA Fade to Black 2.0 — DA Font Grabber — DA Glass — DA Idle — DA MakePICTfile — DA Mouseometer — DA mouseprint — DA Multi-Scrap — DA New Idle1 — DA New Scrapbook2.0 — DA NumCaps 2.0 — DA ParmBlaster — DA ProMouse1.1 — DA Screendump2 — DA Screen-Save — DA Show Clip — DA Show PICTure — DA Stars1.3 — DA Studdclip — DA ZoomIdle 1.1 — EDIT 2.0d1 — Icon Exchanger 2.0 — Install Twelve-C Demo — Inventory — Keeper — MAC ID — MassCopier — microEMACS — Orion1.4 — PageSetupCustomizer — RamDisk+ — VideoWorks. Collection of freeware and shareware. System requirements: Macintosh. Macintosh (Computer).

Anonymous:1988:MDDc

[Ano88d] Anonymous. Macintosh download disk 20, 1988. 1 computer disk.

Anonymous:1988:PSG

[Ano88e] Anonymous. PC scan to get software that scans German text. InfoWorld, 10(38):29, September 19, 1988. CODEN INWODU. ISSN 0199-6649.

Anonymous:1988:TBS

[Ano88f] Anonymous. Typeface book and stack. Macworld, 5(2):334, February 1988. CODEN MACWEA. ISSN 0741-8647. advertisement.

Anonymous:1989:FDa

[Ano89a] Anonymous. Font disk #1, 1989. 1 computer disk.

Anonymous:1989:FDb

[Ano89b] Anonymous. Font disk #2, 1989. 1 computer disk.

Anonymous:1989:FDc

[Ano89c] Anonymous. Font disk #6, 1989. 1 computer disk.

Anonymous:1989:F

[Ano89d] Anonymous. FontWINDOW, 1989.

5 computer disks.

Anonymous:1989:SVR

[Ano89e] Anonymous. SIGGRAPH video review, 1989. 1 videocassette.

Anonymous:1990:FFP

[Ano90a] Anonymous. Fast fonts: Post-Script gets turbocharged: HanZon’s RISC-based controller turns Laser-Jet printers into PostScript hot rods. Byte Magazine, 15(6):184– ??, June 1990. CODEN BYTEDJ. ISSN 0360-5280.

Anonymous:1990:FOA

[Ano90b] Anonymous. Font options at a glance. PC World, 8(6):162–??, June 1990. CODEN PCWDDV. ISSN 0737-8939.

Anonymous:1990:HPS

[Ano90c] Anonymous. How PostScript scaling works. PC World, 8(6):160– ??, June 1990. CODEN PCWDDV. ISSN 0737-8939.

Anonymous:1990:IDI

[Ano90d] Anonymous. IBM declares independence in font wars. Computer Systems News, ??(458):35, March 1990. CODEN CSYND6. ISSN 0164-9981.

Anonymous:1990:JB

[Ano90e] Anonymous. The Japanese TEXnicolor Book. ASCII Corporation, Tokyo, Japan, 1990. ISBN 4-7561-0405-3. Japanese; contents include chapters on the history of TEX and of the Japanese TEX extension, file formats (fonts and .dvi), and device drivers.

Anonymous:1990:MSF

[Ano90f] Anonymous. Making sense of font jargon. PC World, 8(6):161–??, June 1990. CODEN PCWDDV. ISSN 0737-8939.

Anonymous:1990:TST

[Ano90g] Anonymous. Top 20 software tools, 1990. 3 computer disks.

Anonymous:1990:UH

[Ano90h] Anonymous. Under the hood.

Byte Magazine, 15(8):289–??, August 1990. CODEN BYTEDJ. ISSN 0360-5280.

Anonymous:1991:RSM

[Ano91a] Anonymous. 1-2-3 release 3.1, from simple menu operations to advanced font manipulation. PC World, 9(3):257–258, March 1991. CODEN PCWDDV. ISSN 0737-8939.

Anonymous:1991:BSF

[Ano91b] Anonymous. Buying smart: Font shopping tips. PC World, 9(1): 126–127, January 1991. CODEN PCWDDV. ISSN 0737-8939.

Anonymous:1991:CBS

[Ano91c] Anonymous. Can I have both System 6 and System 7 on my Mac? what’s with this font default in Word 4.0? answers to these questions. MacUser, 7(11): 235–244, November 1991. CODEN MCUSEY. ISSN 0884-0997.

Anonymous:1991:EYF

[Ano91d] Anonymous. Exercising your font options. PC World, 9(1):128–137, January 1991. CODEN PCWDDV. ISSN 0737-8939.

Anonymous:1991:ITR

[Ano91e] Anonymous. Is there room in town for more than one font standard? can True Type and Type 1 fonts enjoy peaceful coexistence? find out which font strategy is the best choice for you. MacUser, 7(10): 165–180, October 1991. CODEN MCUSEY. ISSN 0884-0997.

Anonymous:1991:KTF

[Ano91f] Anonymous. Keys, trees, and fonts: Free timesaving, file-tree, and previewing utilities for DOS, Mac, and Unix systems. Byte Magazine, 16 (8):277–??, August 1991. CODEN BYTEDJ. ISSN 0360-5280.

Anonymous:1991:SAF

[Ano91g] Anonymous. Shoot-out at the font corral. MacUser, 7(10):165, October 1991. CODEN MCUSEY. ISSN 0884-0997.

Anonymous:1991:WDM

[Ano91h] Anonymous. Windows display managers file rough edges from text: How’s your face? Windows font managers finally make true WYSI-WYG a reality. Byte Magazine, 16 (6):311–316, June 1991. CODEN BYTEDJ. ISSN 0360-5280.

Anonymous:1992:ABI

[Ano92a] Anonymous. Adobe, Bitstream, and others introduce new font packages and utilities. InfoWorld, 14 (34):16–??, August 1992. CODEN INWODU. ISSN 0199-6649.

Anonymous:1992:AFP

[Ano92b] Anonymous. Adobe font program lets users see and print fonts they don’t have. InfoWorld, 14(45):6– ??, November 1992. CODEN IN-WODU. ISSN 0199-6649.

Anonymous:1992:FMW

[Ano92c] Anonymous. Font managers for Windows 3.1. PC/Computing, 5 (8):362–417, August 1992. CODEN PCMPEI. ISSN 0899-1847.

Anonymous:1992:FW

[Ano92d] Anonymous. A font of wisdom.

PC Week, 9(36):33–??, September 1992. ISSN 0740-1604.

Anonymous:1992:LRS

[Ano92e] Anonymous. LaserTools releases speedy font package for WordPerfect. InfoWorld, 14(34):18–??, August 1992. CODEN INWODU. ISSN 0199-6649.

Anonymous:1992:LUO

[Ano92f] Anonymous. LetraStudio update offers more precise font control and custom envelope libraries—for less money. InfoWorld, 14(??):26, June 1992. CODEN INWODU. ISSN 0199-6649.

Anonymous:1992:M

[Ano92g] Anonymous. MenuFonts 4.01. Macworld, 9(5):180, May 1992. CODEN MACWEA. ISSN 0741-8647.

Anonymous:1992:MWS

[Ano92h] Anonymous. Microsoft will ship a pair of TrueType font packages next week. InfoWorld, 14(46):20– ??, November 1992. CODEN IN-WODU. ISSN 0199-6649.

Anonymous:1992:P

[Ano92i] Anonymous. PersonalFont. Macworld, 9(12):272–??, December 1992. CODEN MACWEA. ISSN 0741-8647.

Anonymous:1992:SMP

[Ano92j] Anonymous. Silenced by my 24 PIN.#xx font file. New statesman society, 5(216):27–??, August 1992. CODEN NESSEF. ISSN 0954-2361.

Anonymous:1992:S

[Ano92k] Anonymous. Standards. UNIX/ world, 9(8):89–??, August 1992. ISSN 0739-5922.

Anonymous:1992:TFP

[Ano92l] Anonymous. TrueType font packs. PC Week, 9(20):28–32, May 1992. ISSN 0740-1604.

Anonymous:1992:T

[Ano92m] Anonymous. Type. MacUser, 8 (10):205–??, October 1992. CO-DEN MCUSEY. ISSN 0884-0997.

Anonymous:1993:CIC

[Ano93a] Anonymous, editor. 1st Color imaging conference — 1993 Nov: Scottsdale; AZ. ????, ????, 1993. ISBN 0-89208-174-0. LCCN ????

Anonymous:1993:ABP

[Ano93b] Anonymous. Acrobat bounds into the paperless publishing arena: Adobe’s Acrobat has a tough balancing act in delivering electronic, no-font-hassle documents to the screen. Byte Magazine, 18(9):30–??, August 1993. CODEN BYTEDJ. ISSN 0360-5280.

Anonymous:1993:AFP

[Ano93c] Anonymous. Apple font pack. Macworld, 10(3):152–??, March 1993. CODEN MACWEA. ISSN 0741-

Anonymous:1993:ATP

[Ano93d] Anonymous. Apple teacher productivity CD: a collection of software for educators, 1993. 1 computer laser optical disk.

Anonymous:1993:CRD

[Ano93e] Anonymous. CD ROM drives: The time is right. PC World, 11(5):215– ??, May 1993. CODEN PCWDDV. ISSN 0737-8939.

Anonymous:1993:CIF

[Ano93f] Anonymous. Corel improves font choice and loading speed with Ventura 4.2. InfoWorld, 15(43):19–??, October 1993. CODEN INWODU. ISSN 0199-6649.

Anonymous:1993:EFW

[Ano93g] Anonymous. ElseWare’s font works emulates more than 100 fonts. InfoWorld, 15(42):19–??, October 1993. CODEN INWODU. ISSN 0199-6649.

Fontaccess:1993:FA

[Ano93h] Anonymous. Font axcess, 1993. 1 computer laser optical disc.

Anonymous:1993:FIP

[Ano93i] Anonymous. FontMonger and Incubator Pro. MacUser, 9(8):71–??, August 1993. CODEN MCUSEY. ISSN 0884-0997.

Anonymous:1993:G

[Ano93j] Anonymous.  Graphics.  PC

Magazine, 12(19):391–??, November 1993. CODEN PCMGEP. ISSN 0888-8507.

Anonymous:1993:IEF

[Ano93k] Anonymous. Image editing and font utilities will share the limelight at this week’s Seybold show. PC Week, 10(14):37–??, April 1993. ISSN 0740-1604.

Anonymous:1993:IAG

[Ano93l] Anonymous. ITC Avant Garde Gothic is now a multiple master typeface. U & lc: The International Journal of Typographics, 20(1):52– ??, Spring 1993. CODEN ULC-CDC. ISSN 0362-6245.

Anonymous:1993:MTFa

[Ano93m] Anonymous. Microsoft’s TrueType font pack 2. PC/Computing, 6 (1):60–??, January 1993. CODEN PCMPEI. ISSN 0899-1847.

Anonymous:1993:MTFb

[Ano93n] Anonymous. Microsoft’s TrueType font pack 2, Microsoft Hewlett– Packard font set. PC/Computing, 6(5):80–??, May 1993. CODEN PCMPEI. ISSN 0899-1847.

Anonymous:1993:NC

[Ano93o] Anonymous. Now in the channel. Computer Shopper, 13(2):893– ??, February 1993. ISSN 0886-0556.

Anonymous:1993:PF

[Ano93p] Anonymous. Personal font. Macworld, 10(4):172–??, April 1993. CODEN MACWEA. ISSN 0741-8647.

Anonymous:1993:Ra

[Ano93q] Anonymous. Reviews. Publish!: the how-to magazine of desktop publishing, 8(1):58–??, January 1993. CO-DEN PBLIEJ. ISSN 0897-6007.

Anonymous:1993:Rb

[Ano93r] Anonymous. Reviews. Publish!: the how-to magazine of desktop publishing, 8(2):70–??, February 1993. CO-DEN PBLIEJ. ISSN 0897-6007.

Anonymous:1993:RAI

[Ano93s] Anonymous. Reviews: Apple II. In-Cider A+, 11(7):20–??, July 1993. ISSN 1054-6456.

Anonymous:1993:SFC

[Ano93t] Anonymous. Stofte’s font collection. The Lawyer’s PC, 11(5):4–??, December 1993. ISSN 0740-0942.

Anonymous:1993:SAF

[Ano93u] Anonymous. Super ATM font program. InfoWorld, 15(11):108–??, March 1993. CODEN INWODU. ISSN 0199-6649.

Anonymous:1993:S

[Ano93v] Anonymous. SuperATM. MacUser, 9(7):62–??, July 1993. CODEN MCUSEY. ISSN 0884-0997.

Anonymous:1993:TFFa

[Ano93w] Anonymous. Typecase font foundry for Macintosh, 1993. 16 computer disks + 1 user’s guide (24 p.: ill.; 13 x 13 cm.).

Anonymous:1993:TFFb

[Ano93x] Anonymous. Typecase font foundry for the Macintosh, 1993. 12 computer disks + 1 user’s guide (24 p. : ill. ; 13 x 13 cm.).

Anonymous:1993:TDM

[Ano93y] Anonymous. Typing direction:

Multiple master typefaces. Art direction, XLV(2):64–??, May 1993. ISSN 0004-3109.

Anonymous:1994:ICA

[Ano94a] Anonymous, editor. 2nd International conference — 1994 Apr: London, Practical Application of Prolog 2nd. Alinmead Software, ????, 1994.

Anonymous:1994:ATP

[Ano94b] Anonymous. Apple teacher productivity CD, 1994. 1 computer laser optical disk.

Anonymous:1994:AWB

[Ano94c] Anonymous. ATM 3.0 will bring multiple master font technology to Windows. InfoWorld, 16(18):19– ??, May 1994. CODEN INWODU. ISSN 0199-6649.

Anonymous:1994:BNF

[Ano94d] Anonymous. Briefly noted: Font CDs – Quickdraw GX – Multiple Masters for Windows – Ed Cleary – Treacyfaces and Headliners International – American Typecasting Fellowship – Emigre’s Whirligig Catalog – Santa Cruz Book Show. Serif: The Magazine of Type and Typography, 1:10–12, Fall 1994. ISSN 1080-2070.

Anonymous:1994:CCI

[Ano94e] Anonymous, editor. Chinese computing: International conference — June 1994, Singapore, volume 4(1) of COMMUNICATIONS- COLIPS. 1994, Chinese and Oriental Languages Information Processing Society, 1994. ISSN 0218-7019.

Anonymous:1994:DP

[Ano94f] Anonymous. Desktop publishing. Instant and small commercial printer, 13(7):62–??, July 1994. ISSN 1044-3746.

Anonymous:1994:EFCa

[Ano94g] Anonymous. Expresiv fonts classic font collection, 1994. 4 computer disks.

Anonymous:1994:EFCb

[Ano94h] Anonymous. Expresiv fonts classic font collection, 1994. 6 computer disks.

Anonymous:1994:EFO

[Ano94i] Anonymous. Expresiv fonts ornate font collection, 1994. 1 compact laser optical disk.

Anonymous:1994:FL

[Ano94j] Anonymous. First looks. PC Magazine, 13(10):37–??, May 1994. CO-DEN PCMGEP. ISSN 0888-8507.

Anonymous:1994:FWW

[Ano94k] Anonymous. Font Works Windows font utility. InfoWorld, 16(23):117– ??, June 1994. CODEN INWODU. ISSN 0199-6649.

Anonymous:1994:Fb

[Ano94l] Anonymous. FontChameleon 1.0.1.

Macworld, 11(9):67–??, September 1994. CODEN MACWEA. ISSN 0741-8647.

Anonymous:1994:Fd

[Ano94m] Anonymous. FontMixer 1.0. Macworld, 11(10):78–??, October 1994. CODEN MACWEA. ISSN 0741-8647.

Anonymous:1994:Fa

[Ano94n] Anonymous. Fontographer 4.0. MacUser, 10(2):62–??, February 1994.

CODEN MCUSEY. ISSN 0884-0997.

Anonymous:1994:Fc

[Ano94o] Anonymous. Fontographer 4.0.4. Macworld, 11(10):57–??, October 1994. CODEN MACWEA. ISSN 0741-8647.

Anonymous:1994:IIP

[Ano94p] Anonymous. Incredible image pak, 1994. 2 computer disks secret.

Anonymous:1994:ICJ

[Ano94q] Anonymous, editor. International conference — 1994 Jun: Singapore, volume 4(1) of COMMUNICATIONS-COLIPS. Chinese and Oriental Languages Information Processing Society, ????, 1994. ISSN 0218-7019.

Anonymous:1994:JMM

[Ano94r] Anonymous, editor. Joint meeting — 1994 May: Rochester; NY, IS AND T ANNUAL CONFERENCE CONF 47//V2. Society for Imaging Science and Technology; 1994, ????, 1994. ISBN 0-89208-177-5. LCCN ????

Anonymous:1994:LWA

[Ano94s] Anonymous. The latest in Windows accounting programs, Font Fiddler, Mathcad 5.0. PC Magazine, 13 (6):37–??, March 1994. CODEN PCMGEP. ISSN 0888-8507.

Anonymous:1994:NAF

[Ano94t] Anonymous. New Adobe font licensing; SprintScan 35 slide and film scanner; ReaySetGo goes GX, and more. Macworld, 11(8):118–??, August 1994. CODEN MACWEA. ISSN 0741-8647.

Anonymous:1994:OMB

[Ano94u] Anonymous. Oppositions: Mark Batty vs. Chuck Byrne. Print, XLVIII(??):44–??, March 1994. ISSN 0032-8510.

Anonymous:1994:RBF

[Ano94v] Anonymous. Russians bring font drawing to PCs. SoftUnion has released FontLab, a well-rounded font creation program for Windows. The Seybold Report on Desktop Publishing, 8(9):4–??, ???? 1994. CODEN SRDFED. ISSN 0889-9762.

Anonymous:1994:TR

[Ano94w] Anonymous. Technology reviews.

Communication arts, 36(6):230– ??, November 1994. CODEN COATBA. ISSN 0010-3519.

Anonymous:1995:BNW

[Ano95a] Anonymous. Briefly noted: Walter Tracy – Toronto font party – Subiaco revival – Catalogs from Font Bureau and Precision Type – Bitstream lawsuit – Type on the information superhighway redux – Carl Herzog award – ATypI ’95 – FUSE ’95. Serif: The Magazine of Type and Typography, 3, Summer 1995. ISSN 1080-2070.

Anonymous:1995:GFF

[Ano95b] Anonymous. Graphics — fonts, font files, and system performance — what are the relationships? PC Magazine, 14(19):550–??, ???? 1995. CODEN PCMGEP. ISSN 0888-8507.

Anonymous:1995:IIP

[Ano95c] Anonymous. Incredible image pak, 1995. 4 computer disks.

Anonymous:1995:PFC

[Ano95d] Anonymous. Printing and fonts — choosing a new font or setting up a printer? let Windows 95 do the dirty work for you. PC/Computing, 8(10):166–??, ???? 1995. CODEN PCMPEI. ISSN 0899-1847.

Anonymous:1995:S

[Ano95e] Anonymous. SmartSuite, 1995. 23 computer disks start cards.

Anonymous:1995:SS

[Ano95f] Anonymous. Software suppliers. Publishing and production executive, 9(12):77–??, December 1995. CODEN PPEXEP. ISSN 1048-3055.

Anonymous:1996:CSH

[Ano96a] Anonymous. Conversion system of high quality Chinese characters from stroke font to outline font. Ruan Jian Xue Bao/Journal of Software, 7(5):257–263, May 1996. CODEN RUXUEW. ISSN 1000-9825.

Anonymous:1996:MRP

[Ano96b] Anonymous. Multi-font recognition of printed Arabic using the BBN Byblos speech recognition system. ICASSP, IEEE International Conference on Acoustics, Speech and Signal Processing — Proceedings, 4:2136–2139, 1996. CODEN IPRODJ. ISSN 0736-7791. IEEE catalog number 96CB35903.

Anonymous:1996:SNN

[Ano96c] Anonymous. Structured neural networks for multi-font Chinese character recognition using a newly developed digital neural network chip with adaptive segmentation of quantizer neuron architecture (ASQA). Neural Networks for Signal Processing — Proceedings of the IEEE Workshop, pages 330– 336, 1996. IEEE catalog number 96TH8205.

Anonymous:1997:HCS

[Ano97a] Anonymous. Highlights continued:

“Singular Impressions: The Monotype in America,” National Museum of American Art. Smithsonian, 28(2):30–??, May 1997. CO-DEN SMSNA5. ISSN 0037-7333.

Anonymous:1997:IWJ

[Ano97b] Anonymous. Interview: Why Java won’t repeat the mistakes of Unix. Scott McNealy, chairman and CEO of Sun Microsystems, talks about how Java will change computing and the computer industry. Byte Magazine, 22(1):40, January 1997. CODEN BYTEDJ. ISSN 0360-5280. Some McNealy quotes: “And I’m urging every CEO to make sure it [Windows] runs really slowly, so people stop using that stuff. The right way is to ban PowerPoint from your company.” “I banned Power-Point from my company and we have had the best two quarters we’ve ever had… ” “I want to give everybody plastic Mylar sheets and all the pens they need to scribble on them. And I said use what I call the [Sun cofounder] Bill Joy font. You can see where he licked his thumb and erases; it’s so much faster.”.

Anonymous:1998:NPE

[Ano98] Anonymous. New products: ENVI 3.0, Research Systems; Xbase for Linux, StarTech; Empress Hypermedia V2.10, Empress Software; SCREADER, Jos Lemmens; ascp0.8.2, Elumaze Nwanua; FontScope, CurveSoft; Pure Java JDBC Driver, Solid Information Technology Ltd.; QMAS-TER Batch & Output Management, QMASTER Software Solutions Inc.; Babylon, SpellCaster Telecommunications Inc.; Corel Video Network Computer, Corel Computer Corp.; W3Control Filter, W3Control Inc. Linux Journal, 46:??, February 1998. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:19xx:BFC

[Anoxx] Anonymous. Berkeley font catalog, October 1980. ????, ????, 19xx. various pp.

Anonymous:2024:SUC

[Ano24] Anonymous. Serif in Use: a Collection of Serif Typefaces. Victionary, Hong Kong, 2024. ISBN 988-768-443-0 (hardcover). 531 + 45 pp. LCCN Z250.5.S47 S47 2024.

Anonymous:20xx:W

[Anoxx] Anonymous. WhatTheFont. Web site, 20xx. URL http://www.

                                        myfonts.com/WhatTheFont/.

Andre:1989:PPE

[AO89] J. André and V. Ostromoukhov. Punk : de à PostScript (English: Punk: from to PostScript). Cahiers GUTenberg, 4:23–28, décembre 1989. ISSN 1140-9304.

Al-otaibi:1990:BSM

[Ao90] Talak Safer Al-otaibi. A base system for multilingual font manipulation. Includes abstract. thesis (m.s.), ????, ????, 1990. x + 103 pp.

Anzai:1995:SHA

[AOM95] Yuichiro Anzai, Katsuhiko Ogawa, and Hirohiko Mori, editors. Symbiosis of human and artifact: proceedings of the Sixth International Conference on Human-Computer Interaction (HCI International ’95), Tokyo, Japan, 9–14 July 1995, volume 20 of Advances in Human Factors/Ergonomics. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1995. ISBN 0-444-81795-6. ISSN 0921-2647. LCCN QA76.9.A73 I564 1995.

Apple:1992:AFP

[App92] Apple Computer, Inc, 1992. 4 computer disks guide.

Arabnia:1995:TRA

[Ara95] Hamid Arabnia, editor. Transputer research and applications 7: American Transputer Users Group, October 23–25, 1994, Atlanta, GA (NATUG-7), volume 42 of Transputer and occam engineering series. IOS Press, Postal Drawer 10558, Burke, VA 2209-0558, USA, 1995. ISBN 90-5199-187-8 (IOS Press) 4-274-90017-7 (Ohmsha). ISSN 0925-4986. LCCN ????

Adams:1989:PFQ

[AS89] D. Adams and R. Southall. Problems of font quality assessment. In André and Hersch [AH89], pages 213–222. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Asher:1990:RVF

[Ash90] Graham Asher. Re: Virtual fonts: More fun for grand wizards. TEXhax, 90(17), January 1990.

Astle:1972:FFS

[Ast72] Brian Astle. Fail-safe font of sevensegment digits. Electronics, 45 (18):93–94, August 1972. CODEN ELECAD. ISSN 0883-4989.

AtechSoftware:1992:VFP

[Ate92] Atech Software. Variety font pack, 1992. 4 computer disks.

Andre:1994:DOS

[AV94] J. André and I. Vatton. Dynamic optical scaling and variable-sized characters. Electronic Publishing —Origination, Dissemination, and Design, 7(4):231–250, December 1994. CODEN EPODEU. ISSN 0894-3982.

Abe:1991:HQG

[AYO91] Hiroshi Abe, Yoshimichi Yamamoto, and Yoshio Ohno. High quality gray-scale Kanji font generation using automatic stroke displacement. In Morris and André [MA91], pages 147–155. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Barber:1980:SCC

[Bar80] Charles Clyde Barber. A supplement to and corrections of C. Faulmanm’s “Buch der Schrift”. ????, Aberdeen, Scotland, 1980. ???? pp.

Barkovic:1986:EFD

[Bar86] Lily Barkovic. An environment for font design. “major subject: Computer science.” thesis (m.s.), Texas A and M University, College Station, TX, USA, 1986. v + 32 pp.

Barrett:1995:CCG

[Bar95] Bonnie Barrett. Callifonts: Can a good font be made in two weeks? Serif: The Magazine of Type and Typography, 3, Summer 1995. ISSN 1080-2070. Barrett takes a hard look at the quality of the types in this broad and inexpensive collection.

Bauermeister:1988:MCT

[Bau88] Benjamin Bauermeister. A manual of comparative typography: the PANOSE system. Van Nostrand Reinhold, New York, NY, USA, 1988. ISBN 0-442-21187-2. xxvii + 257 pp. LCCN Z250 .B34 1988.

Bowden:1989:NIL

[BB89] P. R. Bowden and D. F. Brailsford. On the noise immunity and legibility of Lucida fonts. In André and Hersch [AH89], pages 205–212. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Buchman:1985:DFA

[BBG85] C. Buchman, D. M. Berry, and J. Gonczarowski. DITROFF/ FFORTID, an adaptation of the UNIX DITROFF for formatting bidirectional text. ACM Transactions on Office Information Systems, 3(4):380–397, 1985. CODEN ATOSDO. ISSN 0734-2047.

Bagley:2013:RSVa

[BBK13a] Steven R. Bagley, David F. Brailsford, and Brian W. Kernighan. Revisiting a summer vacation: Digital restoration and typesetter forensics. Report, School of Computer Science, University of Nottingham, and Department of Computer Science, Princeton University, Nottingham NG8 1BB, UK and Princeton, NJ 08540, USA, July

20, 2013. URL http://www.cs.
princeton.edu/~bwk/202/. See

also the original and reconstructed papers [CKT80].

Bagley:2013:RSVb

[BBK13b] Steven R. Bagley, David F. Brailsford, and Brian W. Kernighan. Revisiting a summer vacation: Digital restoration and typesetter forensics. In Proceedings of the 2013 ACM Symposium on Document Engineering, pages 3–12. ACM Press, New York, NY, USA, 2013.

URL https://dl.acm.org/doi/
10.1145/2494266.2494275.

Balashova:2019:LSR

[BBK^{+}19] Elena Balashova, Amit H. Bermano, Vladimir G. Kim, Stephen DiVerdi, Aaron Hertzmann, and Thomas Funkhouser. Learning a strokebased representation for fonts. Computer Graphics Forum, 38(1): 429–442, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Berkebile:1983:DFC

[BDN83] H. E. Berkebile, J. M. Dunn, and L. R. Nielsen. Draft-to-text font conversion algorithm. IBM Technical Disclosure Bulletin, 25(10): 5173–5174, March 1983. CODEN IBMTAA. ISSN 0018-8689.

Braccini:1995:IAP

[BDV95] Carlo Braccini, Leila DeFloriani, and Gianni Vernazza, editors. Image analysis and processing: 8th International Conference, ICIAP ’95, San Remo, Italy, September 13–15, 1995: proceedings, number 974 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-60298-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1632 .I552 1995.

Beccari:1991:SEE

[Bec91a] Claudio Beccari. LaTEX—Guida a un sistema di editoria elettronica. Editore Ulrico Hoepli, Ulrico Hoepli Editore S.p.A., via Hoepli 5, 20121 Milano, Italy, 1991. ISBN 88-203-1931-4. ix + 399 pp. [From the author]: The book is the first and only one in Italian at this moment [18-Nov-1991]. Although it contains an introductory chapter for the beginner, its main purpose is to give the reader advanced information on LaTEX (and partially on TEX) in order to prepare sophisticated macros, style files, and to make a clever use of fonts. Here is a reduced table of contents (translated to English): Chapter 1 - Introduction

- Introductory LaTEX

- General commands for

text mode

- Mathematical com-

mands - Environments

- Figures and tables - Macros

- Typical documents and

style files

- Fonts Appendix A - TEX Users Group

- Keyboards

- Font tables

- Hyphenation pat-

terns for Italian Bibliography Index The publisher received cameraready pages that were prepared by the author with a 300-dpi laser printer. The page style dimensions were suitably enlarged while the photo plates were reduced so as to minimize the granulosity of the signs. In practice the book was prepared with 12pt option and reduced by the photografic ratio of 0.8 so that the final real font size is 9.6pt. The whole book was prepared with the standard Computer Roman fonts on purpose in order to show the potential user what he can get with a standard TEX/LaTEX kit. The simple pictures were prepared with the “picture” environment. The cover and the title page were designed by a graphic expert of the publishing company.

Beccari:1991:TLS

[Bec91b] Claudio Beccari. La tesi di laurea scientifica—Con note per la composizione in LaTEX—disco MS-DOS da 5.25in contenente macroistruzioni di stile. Editore Ulrico Hoepli, Ulrico Hoepli Editore S.p.A., via Hoepli 5, 20121 Milano, Italy, 1991. ISBN 88-203-1932-2. vii + 116 pp. The book was typeset by the author and printed with a laser printer having a resolution of 300 dpi. The pages were then photographically reduced by a factor of 0.8 to the final size. The fonts used are those of the standard TEX/LaTEX set, that is Computer Modern, with a default normal size of 12pt that the photographic reduction made equivalent to approximately 10pt. The book is a set of instructions for the preparation of the thesis (Master and PhD levels) according to the Italian State Universities regulations; it deals mostly with literary style but it also includes extensive tables of scientific symbols and nomenclature and many references to Standards regulations. The final chapter deal with the typesetting of the thesis with a LaTEX style, TOPTESI, whose source is recorded in a floppy disk that comes with the book.

Beeton:1981:UUF

[Bee81] Barbara Beeton. Uppercase update; Fickle fonts. TUGboat, 2(2):42–43, July 1981. ISSN 0896-3207.

Beeman:1982:DFT

[Bee82] Roger L. Beeman. Display of a font in table form. TUGboat, 3(1):35, March 1982. ISSN 0896-3207.

Beebe:1983:LCD

[Bee83] Nelson H. F. Beebe. Low-cost downloadable font devices. TUGboat, 4(1):11, April 1983. ISSN 0896-3207.

Beeton:1985:MSCa

[Bee85a] Barbara Beeton. Mathematical symbols and cyrillic fonts ready for distribution. TUGboat, 6(2):59–63, July 1985. ISSN 0896-3207.

Beeton:1985:MSCb

[Bee85b] Barbara Beeton. Mathematical symbols and cyrillic fonts ready for distribution (revised). TUGboat, 6 (3):124–128, November 1985. ISSN 0896-3207.

Benjamin:1993:DPP

[Ben93a] Mark Benjamin. The design and production of the printed materials for the Goudy International Center for Font Technology and Aesthetics. Pamphlets, etc. are bound together with the text. thesis (m.s.), Rochester Institute of Technology, Rochester, NY, USA, 1993. v + 23 pp.

Bennet-Tec:1993:AMF

[Ben93b] Bennet-Tec Information Systems. ALLText multiple font text box control, 1993. 1 computer disk.

Berry:1990:FF

[Ber90] Karl Berry. Filenames for fonts. TUGboat, 11(4):517–520, November 1990. ISSN 0896-3207.

Berdnikov:1996:CMT

[Ber96a] A. S. Berdnikov. Computer Modern Typefaces as Multiple Master Fonts. TUGboat, 17(2):115–119, June 1996. ISSN 0896-3207.

Berdnikov:1996:VPW

[Ber96b] A. S. Berdnikov. VFComb 1.3— the program which simplifies virtual font management. TUGboat, 17(2): 120–125, June 1996. ISSN 0896-3207.

Berry:2002:LCT

[Ber02] John D. Berry, editor. Language Culture Type: International Type Design in the Age of Unicode. Graphics Press, Box 430, Cheshire, CT 06410, USA, 2002. ISBN 1-932026-01-0. ix + 373 pp. LCCN ???? US$60.00.

Branson:1973:CWL

[BG73] C. A. Branson and J. B. Gillett.

Character wheel for low-cost changeable font printer. IBM Technical Disclosure Bulletin, 16(5):??, October 1973. CODEN IBMTAA. ISSN 0018-8689.

Bertrand:1999:TFS

[BG99] Michael Bertrand and Dave Grundgeiger.

TrueType font secrets. C/C++ Users Journal, 17(8):32, 34–36, 38– 40, August 1999. CODEN CCU-JEX. ISSN 1075-2838.

Bigelow:1991:NAF

[BH91] Charles A. Bigelow and Kris Holmes. Notes on Apple 4 fonts. Electronic Publishing—Origination, Dissemination, and Design, 4(3):171–181, September 1991. CO-DEN EPODEU. ISSN 0894-3982.

Bigelow:1993:DUF

[BH93] Charles Bigelow and Kris Holmes.

The design of a Unicode font.

Electronic Publishing—Origination, Dissemination, and Design, 6 (3):289–305, September 1993. CO-DEN EPODEU. ISSN 0894-3982.

Baines:2002:TT

[BH02] Phil Baines and Andrew Haslam. Type & Typography. Watson-Gup-till Publications, 1515 Broadway, New York, NY 10036, USA, 2002. ISBN 0-8230-5524-8. 192 pp. LCCN Z250.B15 2002. US$29.95.

Bigelow:2014:HWW

[BH14] Charles Bigelow and Kris Holmes. How and why we designed Lucida. Bigelow & Holmes (blog), 2014.

URL https://bigelowandholmes.
typepad.com/bigelow-holmes/2014/
10/how-and-why-we-designed-lucida.
html.

Bigelow:2018:SHB

[BH18] Charles Bigelow and Kris Holmes. Science and history behind the design of Lucida. TUGboat, 39(3):204–211, 2018. CODEN ???? ISSN 0896-3207. URL

https://tug.org/TUGboat/tb39-
3/tb123bigelow-lucida.pdf.

Bigelow:2025:DTC

[BH25] Charles Bigelow and Kris Holmes. Digital type challenges. Visible Language, 59(1):23–54, ???? 2025. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (elec-

tronic).  URL https://www.
visible-language.org/59-1/digital-
type-challenges.

Bresenham:1983:FCH

[BHW83] J. E. Bresenham, J. S. Heath, and C. N. Wallis. Font compression in high resolution printers. IBM Technical Disclosure Bulletin, 25 (9):4582–4584, February 1983. CO-DEN IBMTAA. ISSN 0018-8689.

Bien:1990:SCM

[Bie90] Janusz S. Bień. On standards for computer modern font extensions. TUGboat, 11(2):175–183, June 1990. ISSN 0896-3207.

Bigelow:1985:PSF

[Big85] Charles Bigelow. Principles of structured font design for the personal workstation. ;login: the USE-NIX Association newsletter, 10(4): 65–83, October/November 1985. CODEN LOGNEM. ISSN 1044-6397.

Bigelow:1986:DLI

[Big86a] C. Bigelow. The design of Lucida: An integrated family of types for electronic literacy text processing and document manipulation. In Van Vliet [Van86], pages 1–17. ISBN 0-511-66313-7 (ebook), 0-521-11031-9 (paperback), 0-521-32592-7 (hardcover). LCCN QA76.9.T48 T48 1986.

Bigelow:1986:PTD

[Big86b] C. Bigelow. Principles of type design for the personal workstation. Gutenberg-Jahrbuch, 61(??): 253–270, 1986.

Billawala:1987:WWP

[Bil87] Neenie Billawala. Write-white printing engines and tuning fonts with . TUGboat, 8(1): 29–32, April 1987. ISSN 0896-3207.

Bitstream:1991:BFP

[Bit91] Bitstream Inc. Bitstream font pack for TrueType, 1991. 6 computer disks playing with type + 4 user’s guides + 4 folded sheets of typeface samples.

Bixby:1992:AW

[Bix92] Robert Bixby. Art works. Compute, 14(11):96–??, December 1992. CO-DEN COMPER. ISSN 0194-357X.

Bentley:1988:TPI

[BK88] J. L. Bentley and B. W. Kernighan. Tools for printing indexes. Electronic Publishing—Origination, Dissemination, and Design, 1(1):3– 18, April 1988. CODEN EPODEU. ISSN 0894-3982.

Boggan:1994:UHP

[BL94] Scott Boggan and Michael De Laurentis. Under the hood: The Panose Typeface-Matching system: a look at a font-matching system based on visual characteristics. Byte Magazine, 19(5):187–??, May 1994. CO-DEN BYTEDJ. ISSN 0360-5280.

Butterfield:1987:DMT

[BM87] Steve Butterfield and Shinichi Mori. 24-wire dot matrix technology changes key printer assumptions. Computer Technology Review, 7(9): 101, 104, 108–109, July 1987. CO-DEN CTERES. ISSN 0278-9647.

BMUG:1994:BFR

[BMU94] BMUG, Inc. BMUG font ROM B1, 1994. 1 computer laser optical disk.

BMUG:1995:BFR

[BMU95] BMUG, Inc. BMUG font ROM B, 1995. 1 computer laser optical disc.

Baird:1994:SC

[BN94] Henry S. Baird and George Nagy.

Self-correcting 100-font classifier. Proceedings of SPIE — The International Society for Optical Engineering, 2181:106–115, 1994. CO-DEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic).

Borde:1992:E

[Bor92] Arvind Borde. TEX by Example.

Academic Press, New York, NY, USA, 1992. ISBN 0-12-117650-9. xiv + 169 pp. LCCN Z253.4.T47 B67 1992. This book was designed by the author, with some assistance from the production department of Academic Press, Boston. The covers were done by the publisher. It was typeset by the author using Plain TEX plus specially written macros. The code for all macros used for the book are reproduced in it so that the book itself contains all the information—page layout specifications, fonts used, and so on— about its own typesetting. Over half the book consists of TEX input and output shown side-by-side, providing an explicit example of how a part of the book was typeset. The Academic Press logo on page iii, the cataloging-in-publication data on page iv, and the diagrams on page 33 were pasted in place by the production department of Academic Press. The diagrams were drawn using LaTEX. Sample printouts of pages were made on various laser printers. When everything seemed satisfactory, the dvi files were copied onto IBM diskettes and were mailed to Academic Press, and from them to the American Mathematical Society. The final copy was produced by the A.M.S. using an APS micro-5 phototypesetter. All fonts used in the book are from the Computer Modern family. Parts of the covers of the book, inside and outside, were done in TEX by the author (the code for this is not included in the book). The outside cover was designed by Camille Pecoul of Academic Press and produced at Allison Associates in Boston. The font used for the title (on the covers) is Meridien. The background pattern was obtained by shooting a handmade marble paper as a black-and-white halftone and then overprinting with the Pantone Matching System (PMS). The placement of material on the cover was done by hand by Ms. Pecoul using a mechanical board. The drop shadow effect was done at Allison Associates.

Bowman:1968:COD

[Bow68] Robert Moffett Bowman. Considerations for the optimal design of a single font character recognition system. Thesis (m.e.e.), University of Virginia, Charlottesville, VA, USA, 1968. ix + 81 pp.

Brailsford:2013:GJC

[Bra13] David Brailsford. The great 202 jailbreak — computerphile. Web video lecture (19m54s)., December 12, 2013. URL

                                        http://www.cs.princeton.edu/
                                        ~bwk/202;   http://www.eprg.
                                        org/papers/202paper.pdf;   https:
                                        //www.youtube.com/watch?v=CVxeuwlvf8w.

Bringhurst:2016:PNHa

[Bri16a] Robert Bringhurst. Palatino: the natural history of a typeface, volume 235 of Book Club of California publication. The Book Club of California, San Francisco, CA, USA, 2016. ISBN 0-9819597-8-4. 292 pp. LCCN Z250.5.P35 B75 2016. URL

http://tug.org/books/reviews/
tb116reviews-bringhurst.html.

Bringhurst:2016:PNHb

[Bri16b] Robert Bringhurst. Palatino: The Natural History of a Typeface. David R. Godine, Publisher, Jaffrey, NH, USA, 2016. ISBN 1-56792-572-3. ???? pp. LCCN

????  URL http://tug.org/
books/reviews/tb116reviews-bringhurst.
html.

Brown:1985:FID

[Bro85] Heather Brown. Font information and device-independent output. NATO ASI Series, Series F: Computer and Systems Sciences, 17:397–407, 1985. CODEN NAS-FEG. ISBN 3-540-13920-6. ISSN 0258-1248.

Brody:1990:NPH

[Bro90] Alan Brody. The new power of the HP LaserJet III. Datamation, 36(19):103–??, October 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Brown:1991:ULO

[Bro91] P. J. Brown. Using logical objects to control hypertext appearance. Electronic Publishing—Origination, Dissemination, and Design, 4(2):109–118, June 1991. CODEN EPODEU. ISSN 0894-3982.

Brown:1994:CDD

[Bro94] J. Randolph Brown. Cross-platform database development. strategies for FoxPro developers. Dr. Dobb’s Journal of Software Tools, 19(6):84, 88, 90–91, 102, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Browne:1998:FDA

[Bro98] Cameron Browne. Font decoration by automatic mesh fitting. Lecture Notes in Computer Science, 1375:23–43, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://diwww.epfl.
                                        ch/w3lsp/ridt98.html;   http:
                                        //link.springer-ny.com/link/
                                        service/series/0558/bibs/1375/
                                        13750023.htm;   http://link.
                                        springer-ny.com/link/service/
                                        series/0558/papers/1375/13750023.
                                        pdf.

Bryan:1984:DDD

[Bry84] Shawn Bryan. Datasouth’s do-it-all dynamo. Microcomputing, 8(9):56– 58, September 1984. CODEN MIR-CDC.

Bigelow:1981:TAT

[BS81] C. Bigelow and J. Seybold. Technology and the aesthetics of type: Maintaining the “tradition” in the age of electronics. The Seybold Report, 10(24):??, ???? 1981.

Bigelow:1982:PDTa

[BS82a] C. Bigelow and J. Seybold. The principles of digital type: Quality type for low, medium and high resolution printers. The Seybold Report, 11(8):??, 1982.

Bigelow:1982:PDTb

[BS82b] C. Bigelow and J. Seybold. The principles of digital type: Quality type for low, medium and high resolution printers. The Seybold Report, 11(12):??, 1982.

Barrera:1984:EPC

[BS84] C. Barrera and A. V. Strietzel. Electrophotographic printer control as embodied in the IBM 3800 printing subsystem models 3 and 8. IBM Journal of Research and Development, 28(3):263–275, May 1984. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Bo:1996:SAA

[BS96] Lu Bo and Cai Shijie. Skeletonbased approach of automatically generating some Chinese typefaces. Journal of computer science and technology, 11(1):30–38, January 1996. CODEN JCTEEM. ISSN 1000-9000.

Braun:1992:LRW

[BSS92] Curt C. Braun, N. Clayton Silver, and Barry R. Stock. Likelihood of reading warnings: the effect of fonts and font sizes. Proceedings of the Human Factors Society, 2:926–930, 1992. CODEN PHFSDQ. ISSN 0163-5182.

Brassel:1978:FVV

[BU78] K. E. Brassel and J. J. Utano. Font variation in vector plotting lettering. Computer Graphics, 11(4):67– 77, March 1978. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Burke:2018:GUO

[Bur18] Christopher Burke. Gerard Unger obituary: Typeface designer whose work adorns many newspapers, books and road signs. The Guardian, ??(??):??, December 5, 2018. ISSN 0261-3077 (print), 1756-3224 (electronic).

URL https://www.theguardian.
com/artanddesign/2018/dec/05/
gerard-unger-obituary.

Busch:1992:FPSa

[Bus92a] David D. Busch. The font problem solver: how to install and use type with your software and printer. Business One Irwin, Homewood, IL, USA, Macintosh edition, 1992. ISBN 1-55623-505-4. various pp. LCCN ????

Busch:1992:FPSb

[Bus92b] David D. Busch. The font problem solver: how to install and use type with your software and printer. Business One Irwin desktop publishing library. Business One Irwin, Homewood, IL, USA, PC edition, 1992. ISBN 1-55623-506-2. 330 pp. LCCN ????

Butler:1994:RDF

[But94] Timothy Butler. Retaining document format in OCR. Proceedings of SPIE — The International Society for Optical Engineering, 2181:78–86, 1994. CO-DEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic).

Buynak:1985:LLM

[Buy85] Buynak, William, Chyron Corp.

Large low-cost memory buffers create improved picture quality. Computer Technology Review, 5(2):169, 172, 174–175, Spring 1985. CODEN CTERES. ISSN 0278-9647.

Baudin:1985:AFA

[BWT85] Fernand Baudin, Inga Wennik, and Margaret Tsai. Abstracts — French; abstracts — German; abstracts — Japanese. Visible Language, XIX(1):157–163, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_
E.pdf.

Campbell:1991:EFG

[Cam91] George Campbell. Easy fonts for great-looking documents. PC World, 9(1):122–125, January 1991. CODEN PCWDDV. ISSN 0737-8939.

Campbell:1992:WPC

[Cam92] George Campbell. Word processing clinic. PC World, 10(2):241– 246, February 1992. CODEN PCWDDV. ISSN 0737-8939.

Carter:1984:IAS

[Car84] K. A. Carter. IMP — a system for computer-aided typeface design. In PROTEXT I. Proc. the First Int. Conf. on Text Processing, pages 114–119. Boole Press, Dublin, Ireland, 1984.

Carter:1985:GMR

[Car85] Matthew Carter. Galliard: a modern revival of types of Robert Granjon. Visible Language, XIX (1):77–98, Winter 1985. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (elec-

tronic). URL https://s3-us-
west-2.amazonaws.com/visiblelanguage/
pdf/V19N1_1985_E.pdf.

Cary:2010:SPW

[Car10] Melbert B. Cary, Jr. A specimen portfolio of wood type in the Cary collection: Melbert B. Cary, Jr. Graphic Arts Collection Rochester Institute of Technology. RIT Carey Graphic Arts Press, Rochester, NY, USA, 2010. ISBN 1-933360-44-5. 305 pp. LCCN Z250.5.W65 M44 2010.

Casti:1993:B

[Cas93] John L. Casti. Bitmap-free TEX. Notices of the American Mathematical Society, 40(5):464–468, May/June 1993. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). A review of Y&Y’s DVIWindo and DVIP-SONE, and comments on outline fonts for TEX.

Catt:1983:CPD

[Cat83] B. R. L. Catt. Computer printer — out of the DP room and into the office. IERE Conference Proceedings, 57:133–139, 1983. CO-DEN IERCB6. ISBN 0-903748-52-

5. ISSN 0538-0006.

Cox:1974:ATF

[CBE74] C. Cox, B. Blesser, and M. Eden.

The application of type font analysis to automatic character recognition. In Proceedings of the Second International Joint Conference on Pattern Recognition, pages 226– 232. ????, ????, 1974.

Cheng:1983:SMC

[CC83] Kuo-Young Cheng and Keh-Jiann Chen. System model for Chinese character fonts generation. Chung-kuo Kung Ch’eng Hsueh K’an/Journal of the Chinese Institute of Engineers, 6(3):175–182, July 1983. CODEN CKCKDZ.

Chen:1988:SRC

[CCH88] P. N. Chen, Y. S. Chen, and W. H.

Hsu. Stroke relation coding — a new approach to the recognition of multi-font printed Chinese characters. International Journal of Pattern Recognition and Artificial Intelligence, 2:149–160, 1988. CO-DEN IPJIEI. ISSN 0218-0014.

Chan:1988:JAC

[CCL88] Kai Ming Chan, Kam Pui Chow, and Wai Man Lo. JACM—just another Chinese Metafont. In ICCP-COL ’88 [ICC88], pages 311–315. LCCN QA76.9.T48 I48 1988.

Cordella:1995:NNC

[CDV95] L. P. Cordella, C. De Stefano, and M. Vento. Neural network classifier for OCR using structural descriptions. Machine Vision and Applications, 8(5):336–342, 1995. CODEN MVAPEO. ISSN 0932-8092.

CentralLinotypers:1988:TFC

[Cen88] Central Linotypers, Inc. Type faces: a complete type library listed in condensed form plus full font showings. Central Linotypers, inc., New York, NY, USA, 1988. 117 pp.

Crookston:1992:FM

[CFG92] Clair Crookston, Florida Marketing International, and Greg Rike Productions. Font management, 1992. 1 videocassette (ca. 71 min.).

Casey:1982:ASD

[CFW82] R. G. Casey, T. D. Friedeman, and K. Y. Wong. Automatic scaling of digital print fonts. IBM Journal of Research and Development, 26:657– 666, November 1982. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Carnase:1995:TBD

[CGP95] Tom Carnase, Baruch Gorkin, and B. Martin Pedersen. Type: the best in digital classic text fonts. Graphis digital font annual; 1. Graphis Press Corp., Zurich, Switzerland, 1995. ISBN 3-85709-457-5. xviii + 217 pp. LCCN ????

Chang:1983:EDN

[Cha83] Chee-Shan S. Chang. The experimental design of a new Chinese TV font. Thesis (m.f.a.), University of Kansas, Design, ????, 1983. 69 pp.

Chalk:1988:IFC

[Cha88] Andrew J. Chalk. An Italic font in C for the EGA and VGA. Dr. Dobb’s Journal of Software Tools, 13(3):36–??, March 1988. CODEN DDJOEB. ISSN 1044-789X.

Cherlin:1994:WIU

[Che94] Edward Cherlin. The Worldwide Impact of the Unicode Character Set. Character Type, 6611 Linville Drive, Weed, CA 96094, 1994. vi + 181 pp. US$995.00.

Cheung:2022:TLB

[Che22] Victor Cheung, editor. From Type to Logo: the Best Logotypes from Around the World. Victionary, Hong Kong, 2022. ISBN 988-756-659-4 (hardcover). 269 + 3 pp. LCCN NC1002.L63 F76 2022.

Cheung:2023:SUC

[Che23] Victor Cheung. Sans in Use: Creative Typefaces and Their Applications. From A to Z. Book A. Victionary, North Point, Hong Kong, 2023. ISBN 988-768-442-2. 576 pp. LCCN Z250.5.S24 S24 2024.

Chi:1984:FER

[Chi84] Chin-Huang Chi. A Font Editor for the RAMTEK 6211 colorgraphics terminal. ????, ????, 1984. vii + 74 pp.

Childs:1988:FSM

[Chi88] Bart Childs. TEX fonts and suggested magnifications. TUGboat, 9 (2):129–131, August 1988. ISSN 0896-3207.

Chung:1984:CFD

[Chu84] Y. Z. Chung. Computerized font design system for Hangul characters (in Korean). Journal of the Korea Information Science Society = Chongbo Kwahakhoe nonmunji, 11 (3):153–163, August 1984. CODEN HJKHDC. ISSN 8264-5521.

Casey:1983:POS

[CJ83] R. G. Casey and C. R. Jih. Processor-based OCR system. IBM Journal of Research and Development, 27(4):386–399, July 1983. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Cannat:1986:LTAb

[CK86] J.-J. Cannat and Y. Kodratoff. Learning techniques applied to multi-font character recognition. Proceedings of SPIE — The International Society for Optical Engineering, 635:469–479, 1986. CO-DEN PSISDG. ISBN 0-89252-670-X. ISSN 0277-786X (print), 1996-756X (electronic).

Cannat:1986:LTAa

[CKM86] J.-J. Cannat, Y. Kodratoff, and S. Moscatelli. Learning techniques applied to multi-font character recognition. In IEEE [IEE86], pages 123–125. CODEN PICREG. ISBN 0-8186-0742-4. LCCN Q 327 I615 1986. IEEE Service Cent. Piscataway, NJ, USA.

Condon:1980:EML

[CKT80] Joe Condon, Brian Kernighan, and Ken Thompson. Experience with the Mergenthaler Linotron 202 phototypesetter, or, how we spent our summer vacation. Technical Memorandum TM-80-1270-1, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 6, 1980.

11 + 2 pp. URL http://www.cs.
princeton.edu/~bwk/202/. Pub-

lication of this internal memo of 1980 was suppressed for 33 years, but both the original, and the reconstructed memo are now available on the Web. See [BBK13b] for the story of its reconstruction and Web publication.

Clark:1990:AUM

[Cla90] Malcolm Clark, editor. TEX Applications, Uses, Methods. Ellis Horwood, New York, NY, USA, 1990. ISBN 0-13-912296-6.

Chen:1994:MXW

[CM94] C. K. Chen and Y. S. Moon. Modification of X window system font server to support Chinese outline fonts. In Anonymous [Ano94q], pages 76–84. ISSN 0218-7019.

Cohen:1993:ZPD

[Coh93] Michael Cohen. Zebrackets: a pseudo-dynamic contextually adaptive font. TUGboat, 14(2):118–122, July 1993. ISSN 0896-3207.

Collins:1986:DFL

[Col86] J. S. Collins. Digital fonts for lowcost typesetting systems. Proceedings — IEEE Computer Society International Conference, pages 82– 87, 1986. CODEN PCICDQ. ISBN 0-8186-0692-4. IEEE Service Cent. Piscataway, NJ, USA.

Conrad:1989:FTU

[Con89] Arvin C. Conrad. Fine typesetting with TEX using native Autologic fonts. TUGboat, 10(4):521– 528, December 1989. ISSN 0896-3207.

Connare:1998:DUS

[Con98] Vincent Connare. Designing and using scalable type for display as graphic elements on digital devices. Lecture Notes in Computer Science, 1375:181–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/1375/13750181.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/1375/
13750181.pdf.

Correll:2000:GER

[Cor00] Sharon Correll. Graphite: An extensible rendering engine for complex writing systems. In Unicode Consortium [Uni00], page 21. ISBN ???? LCCN ???? URL http:

//graphite.sil.org/pdf/IUC17_
paper.pdf.

Cowlishaw:1987:LPS

[Cow87] M. F. Cowlishaw. LEXX — a programmable structured editor. IBM Journal of Research and Development, 31(1):73–80, January 1987. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Corthout:1991:SOF

[CP91a] Marc Corthout and Evert J. Pol. Supporting outline font rendering in dedicated silicon: the PHAROS chip. In Morris and André [MA91], pages 177–189. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Cousquer:1991:PCE

[CP91b] A. Cousquer and É. Picheral. Polices, TEX et cie (English: Fonts, TEX et al). Cahiers GUTenberg, 9: 3–31, juillet 1991. ISSN 1140-9304.

Crabb:1989:MR

[Cra89] D. Crabb. A Macintosh retrospective. Byte Magazine, 14(3):143–144, 146, March 1989. CODEN BYT-EDJ. ISSN 0360-5280.

Crabb:1990:FFP

[Cra90] D. Crabb. Fast fonts: PostScript gets turbocharged (printer interface). Byte Magazine, 15(6):184– 186, June 1990. CODEN BYTEDJ. ISSN 0360-5280.

Crow:1978:UGS

[Cro78a] F. C. Crow. The use of grey scale for improved raster display of vectors and characters. Computer Graphics, 12(3):1–5, August 1978. CO-DEN CGRADI. ISSN 0097-8930.

Crow:1978:UGI

[Cro78b] Franklin C. Crow. The use of grayscale for improved raster display of vectors and characters. Computer Graphics, 12(3):1–5, August 1978. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Wolfram:1987:CS

[CS87] Complex Systems. Complex Systems Publications Inc., 1987. ISSN 0891-2513. Technical journal published six times per year, using Computer Modern fonts.

Cleary:1993:FDT

[CSS93] Ed Cleary, Jurgen Siebert, and Erik Spiekermann. FontBook: digital typeface compendium. FontShop International, Berlin, Germany, 1993. ISBN 3-930023-00-8 (??invalid checksum??). various pp. LCCN ????

Cusick:2011:WOL

[Cus11] Rick Cusick. What our lettering needs: the contribution of Hermann Zapf to calligraphy and type design at Hallmark Cards. RIT Cary Graphic Arts Press, Rochester, NY, USA, 2011. ISBN 1-933360-55-0 (paperback). xv + 117 pp. LCCN Z250.A2 Z374 2011.

IEEE:1986:PCI

[CVP86] Proceedings, CVPR ’86 (IEEE Computer Society Conference on Computer Vision and Pattern Recognition, Miami Beach, FL, June 22–26, 1986), IEEE Publ. 86CH2290-5. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. ISBN 0-8186-0721-1. LCCN TA1632 .I36 1986.

Casey:1985:TRU

[CW85] R. G. Casey and K. Y. Wong. Text recognition using adaptive software. In GLOBECOM ’85: IEEE Global Telecommunications Conference — Conference Record, New Orleans, LA, USA, pages 353– 357. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. IEEE Service Cent. Piscataway, NJ, USA.

Chen:1993:WSS

[CWB93] Francine R. Chen, Lynn D. Wilcox, and Dan S. Bloomberg. Word spotting in scanned images using hidden Markov models. Proceedings — ICASSP, IEEE International Conference on Acoustics, Speech and Signal Processing, 5:V–1–V–4, 1993. CODEN IPRODJ. ISBN 0-7803-0946-4. ISSN 0736-7791. IEEE catalog number 93CH3252-4.

Chang:1992:CAD

[CY92] K.-Y. Chang and G. T. Yang.

CCIGS: a data compression system for Chinese fonts and binary images using classification techniques. Software—Practice and Experience, 22 (12):1027–??, December 1992. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:1988:NNS

[cYL88] Ming chih Yeh Lee. Neural network structure modeling: an application to font recognition. Thesis (m.s.), Texas Tech University, Lubbock, TX, USA, 1988. viii + 125 pp.

Cyrus:1988:DSE

[Cyr88] Wes Tait Cyrus. Design of a screen editor which supports variable font widths. Thesis (m.s.), University of New Mexico, Albuquerque, NM, USA, 1988. ix + 114 pp.

Dako:1990:FAF

[Dak90] Peter Dako. A family affair:

FONTOGRAPHER, THE ART IMPORTER, FAMILY BUILDER. Electronic Composition and Imaging, 4(??):44–47, May 1990. ISSN 0838-9535.

Damon:2016:JST

[Dam16] Dan Damon. Johnston Sans: The Tube typeface that changed everything. BBC News, ??(??):??, March

30, 2016. URL http://www.bbc.
com/news/magazine-35916807.

Dardailler:1989:NFE

[Dar89] D. Dardailler. Normes et fontes (English: standards and fonts). Cahiers GUTenberg, 4:2–8, décembre 1989. ISSN 1140-9304.

DeRee:1993:FUC

[DB93] H. De Ree and G. Blok. Font to use in cockpit documentation. In Jensen and Neumeister [JN93], pages 115–118.

Dubois:1993:SPH

[DC93] E. (Eric) Dubois and L. (Leonardo) Chiariglione, editors. Signal processing of HDTV, IV: proceedings of the International Workshop on HDTV’92, Kawasaki, Japan, November 18–20, 1992, SIG-NAL PROCESSING OF HDTV CONF 4. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1993. ISBN 0-444-81551-1. LCCN TK6679.I57 1992.

Demaine:2015:FFA

[DD15] Erik D. Demaine and Martin L. Demaine. Fun with fonts: Algorithmic typography. Theoretical Computer Science, 586(??):111– 119, June 27, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/S030439751500167X.

Dearborn:1785:SRS

[Dea85] Benjamin Dearborn. A scheme, for reducing the science of music to a more simple state,: and to bring all its characters within the compass of a common fount of printingtypes; especially calculated for the convenience of learners. ????, Portsmouth New-Hampshire, 1785. 16 pp. LCCN Microfiche M 00235 no. 44674 Micro. Bristol B6055. Shipton and Mooney 44674. Microfiche. [New York: Readex Microprint, 1985] 11 x 15 cm. (Early American imprints. First series; no. 44674).

Deach:1992:OFH

[Dea92a] Stephen Deach. Outline font hints and rasterization: a technology primer. The Seybold Report on Desktop Publishing, 6(7):21–31, March 9, 1992. CODEN SRDFED. ISSN 0889-9762. This paper is followed by a dissent from Adobe (Terry O’Donnell), and a response from the author (p. 32). A sidebar (historical note, p. 25) contains a short description of and cites as a forerunner of the hinting mechanism used in TrueType.

Dean:1992:FC

[Dea92b] David Dean. Font comparisons.

Electronic Composition and Imaging, 6(??):20–27, October 1992. ISSN 0838-9535.

Dean:1992:MYT

[Dea92c] David Dean. Multiplying your typographic possibilities Adobe Systems’ Multiple Master font format. Electronic Composition and Imaging, ????(????):8–11, March 1992. ISSN 0838-9535.

Dean:1993:SEF

[Dea93a] David Dean. Special effects font software. Electronic Composition and Imaging, 7(4):16–17, July 1993. ISSN 0838-9535.

Dean:1993:WFM

[Dea93b] David Dean. Windows font management. Electronic Composition and Imaging, 7(3):32–35, May 1993. ISSN 0838-9535.

DeCicco:1983:EDE

[DeC83] Mary Jean DeCicco. An experimental determination of the effects of dot matrix/character size and font on reading times and eye movements. Abstract. thesis (m.s.), Virginia Polytechnic Institute and State University, Blacksburg, VA, USA, 1983. viii + 115 pp.

Desarmenien:1984:HRF

[Dés84] Jacques Désarménien. How to run TEX in a French environment: Hyphenation, fonts, typography. TUGboat, 5(2):91, November 1984. ISSN 0896-3207.

Desarmenien:1986:SDS

[Dés86] Jacques Désarménien, editor. TEX for Scientific Documentation. Second European Conference. Strasbourg, France, June 19–21, 1986. Proceedings, number 236 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Desgraupes:1999:MGP

[Des99] Bernard Desgraupes. Metafont: Guide pratique. Vuibert, Paris, France, 1999. ISBN 2-7117-8642-0. 267 pp. LCCN ????

deBry:1989:MMF

[dGH89] R. deBry, A. W. Griffee, and J. P.

Hofmeister. Management of multiple font technologies in a distributed system environment. In André and Hersch [AH89], pages 223–231. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Dick:1983:UPG

[DH83] C. E. Dick and Joseph Hilsenrath. Utility programs for generating the Hershey character fonts on microcomputers and laboratory plotters. Technical Note TN-1176, U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1983.

diMichele:1995:IF

[di 95] A. di Michele. El Ink (font). The New Orleans review, 21(2):45–??, Summer 1995. ISSN 0028-6400.

DiVittorio:1984:HDI

[DiV84] Mark J. DiVittorio. High-quality, dot-matrix impact printer family. Hewlett–Packard Journal: technical information from the laboratories of Hewlett–Packard Company, 35(11): 30–32, November 1984. CODEN HPJOAX. ISSN 0018-1153.

Dong:1991:PGA

[DL91] Yunmei Dong and Kaide Li. A parametric graphics approach to Chinese font design. In Morris and André [MA91], pages 156–165. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Duan:1989:IPS

[DM89] G. Duan and R. A. Morris. The importance of phase in the spectra of digital type. Electronic Publishing —Origination, Dissemination, and Design, 2(1):47–60, April 1989. CO-DEN EPODEU. ISSN 0894-3982.

Deininger:1992:UNF

[DM92] Axel Deininger and Nathan Meyers. Using the new font capabilities of HP-donated font server enhancements. The X Resource, 3(1): 97–120, June 1992. CODEN XRE-SEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Devroye:1995:RFS

[DM95] Luc Devroye and Michael Mc-Dougall. Random fonts for the simulation of handwriting. Electronic Publishing—Origination, Dissemination, and Design, 8(4):281– 294, December 1995. CODEN EPODEU. ISSN 0894-3982.

Duncan:1969:CGS

[DMK69] S. Duncan, T. Mukaii, and S. Kuno. A computer graphics system for non-alphabetic orthographies. Computer Studies in the Humanities, 2(3):113–132, October 1969.

Doherty:1981:TFC

[Doh81] Barry C. W. Doherty. TUG Font Committee. TUGboat, 2(2):34–35, July 1981. ISSN 0896-3207.

Doob:1993:VFP

[DP93] Michael Doob and Craig Platt. Virtual fonts in a production environment. TUGboat, 14(3):275–281, October 1993. ISSN 0896-3207.

Dreyfus:1985:TPT

[Dre85] John Dreyfus. A turning point in type design. Visible Language, XIX(1):11–22, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (elec-

tronic). URL https://s3-us-
west-2.amazonaws.com/visiblelanguage/
pdf/V19N1_1985_E.pdf.

Drost:1985:PD

[Dro85] Henk Drost. Punchcutting demonstration. Visible Language, XIX (1):99–106, Winter 1985. CO-DEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (elec-

tronic). URL https://s3-us-
west-2.amazonaws.com/visiblelanguage/
pdf/V19N1_1985_E.pdf.

Durbeck:1988:OHD

[DS88] Robert C. Durbeck and Sol Sherr, editors. Output Hardcopy Devices. Academic Press, New York, NY, USA, 1988. ISBN 0-12-225040-0. xii + 526 pp. LCCN TK7887.7 .O981 1988. See article by Mike Parker on printer fonts, and undesirability of white-writing engines. I have not bought this book.

Dubl-ClickSoftware:1990:MRF

[Dub90] Dubl-Click Software, Inc. Menu-Fonts the real font menu, 1990. 1 computer disk.

Duerst:1993:CFD

[Due93a] M. J. Duerst. Coordinateindependent font description using Kanji as an example. In André et al. [AGS93b], pages 133–144. CODEN EPODEU. ISBN 0-471-94823-3. ISSN 0894-3982. LCCN ????

Duerst:1993:SCD

[Due93b] M. J. Duerst. Structured character description for font design: a preliminary approach based on prolog. In Shin and Kunii [SK93], pages 369–382. ISBN 981-02-1529-0. LCCN T385.C5575 1993.

Duerst:1994:PSC

[Due94] M. J. Duerst. A prolog for structured character description for font design. In Anonymous [Ano94a], pages 207–222.

Duncan:1992:WCD

[Dun92] Ray Duncan. The Windows common dialogs for font and color selection. PC Magazine, 11(11):393– 400, June 1992. CODEN PCMGEP. ISSN 0888-8507.

Durst:1993:CFD

[Dür93a] Martin J. Dürst. Coordinateindependent font description using Kanji as an example. Electronic Publishing—Origination, Dissemination, and Design, 6(3):133– 143, September 1993. CODEN EPODEU. ISSN 0894-3982.

Durst:1993:SCD

[Dur93b] Martin J. Durst. Structured character description for font design: a preliminary approach based on prolog. In First Pacific Conference on Computer Graphics and Applications. Korean Information Science Society, Korean Computer Graphics Society, ????, ????, August 1993.

Durazzi:2022:RTM

[Dur22] Allison Durazzi. Review of “Type Matters: The Rhetoricity of Letterforms edited by Christopher Scott Wyatt and Dànielle Nicole De-Voss,” (2018). Parlor Press. Communication Design Quarterly Review, 10(1):54–56, March 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

URL https://dl.acm.org/doi/
10.1145/3507454.3507460.

Diehl:1990:LPG

[DW90] S. Diehl and S. Wszola. Laser printers get personal (buyer’s guide). Byte Magazine, 15(7):138–144, 146, 148, 150–152, 154, 156, July 1990. CODEN BYTEDJ. ISSN 0360-5280.

ElectronicArts:1987:OP

[EA87] Electronic Arts and Advanced Memory Systems, Ltd. Outrageous pages, 1987. 3 computer disks.

Earnest:1976:FF

[Ear76] Les Earnest. Find a font. Operating Note 74, Stanford Artificial Intelligence Laboratory, Stanford, CA, USA, May 1976. iv + 58

pp. URL https://www.saildart.
org/allow/pdf/SAILON/74/find-
a-font.pdf.

Earls:2002:DT

[Ear02] David Earls. Designing Typefaces.

Rotovision, Hove, East Sussex BN3 1DD, England, 2002. ISBN 2-88046-699-7. 160 pp. LCCN ???? US$40.00.

ElGowely:1990:MRM

[EEN90] Khaled El Gowely, Ossama El Dessouki, and Ahmed Nazif. Multiphase recognition of multi-font photoscript Arabic text. Proceedings — International Conference on Pattern Recognition, 1:700–702, 1990.

CODEN PICREG. ISBN 0-8186-2062-5. IEEE catalog number 90CH2898-5.

Eglowstein:1990:PPY

[Egl90] Howard Eglowstein. PostScript in the palm of your hand: Pacific Data’s new cartridge gives HP LaserJet II printers easy PostScript compatibility. Byte Magazine, 15 (1):197–200, January 1990. CO-DEN BYTEDJ. ISSN 0360-5280.

Eglowstein:1991:WDM

[Egl91] H. Eglowstein. Windows display managers file rough edges from text. Byte Magazine, 16(6):311– 312, June 1991. CODEN BYTEDJ. ISSN 0360-5280.

Ellis:1989:PBC

[Ell89] Wm. Edwin Ellis. PostScript, Bézier curves, and Chinese characters. In Riehl [Rie89], pages 162– 165. ISBN 0-89791-299-3. LCCN QA75.5 .A1371 1989. US$36.00.

Esakov:1994:CDO

[ELS94] Jeffrey Esakov, Daniel P. Lopresti, and Jonathan S. Sandberg. Classification and distribution of optical character recognition errors. Proceedings of SPIE — The International Society for Optical Engineering, 2181:204–216, 1994. CO-DEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic).

Elstner:1984:DFR

[EM84] J. Elstner and D. Maignan. Diagnostics for font read-only store. IBM Technical Disclosure Bulletin, 27(5):3049–??, October 1984. CO-DEN IBMTAA. ISSN 0018-8689.

Emmett:1995:TTA

[Emm95] Arielle Emmett. 3D type — with today’s animated 3D titling packages, which offer an array of typography, compositing, and post-production features, conjuring up something more than just the jaded flying font is fairly easy. Computer Graphics World, 18(11):35–??, 1995. CO-DEN CGWODH. ISSN 0271-4159.

Elliott:1990:BB

[EMP90] Alan C. Elliott, Dwight Morgan, and Richard F. Parker. Banner bonanza, 1990. 1 computer disk.

English:1987:UOF

[Eng87] Paul M. English. Using for original font design. August 1987.

F:2010:MRM

[F.10] G. F. The most read man in the world: Matthew Carter designs what everyone reads. Babbage: The Economist, ??(??):??, December 2, 2010. URL https:

//www.economist.com/babbage/
2010/12/02/the-most-read-man-
in-the-world.

Faulmann:1878:BSE

[Fau78] Carl Faulmann. Das Buch der Schrift enthaltend die Schriftzeichen und Alphabete aller Zeiten und aller Völker des Erdkreises. (German) [The Book of Scripts containing the characters and alphabets of all times and peoples of the Earth]. Kaiserlich-königliche Hofund Staatsdruckerei, Wien, Austria, 1878. ???? pp.

Faulmann:1880:BSE

[Fau80] Carl Faulmann. Das Buch der Schrift enthaltend die Schriftzeichen und Alphabete aller Zeiten und aller Völker des Erdkreises. (German) [The Book of Scripts containing the characters and alphabets of all times and peoples of the Earth]. Kaiserlich-königliche Hofund Staatsdruckerei, Wien, Austria, second edition, 1880. xii + 286 pp.

Faulmann:1985:BSE

[Fau85] Carl Faulmann. Das Buch der Schrift enthaltend die Schriftzeichen und Alphabete aller Zeiten und aller Völker des Erdkreises. (German) [The Book of Scripts containing the characters and alphabets of all times and peoples of the Earth]. Franz Greno Verlag, Nördlingen, Germany, 1985. ISBN 3-921568-51-X. 286 pp. LCCN ???? EUR 9.95.

Faulmann:1986:BSEa

[Fau86a] Carl Faulmann. Das Buch der Schrift enthaltend die Schriftzeichen und Alphabete aller Zeiten und aller Völker des Erdkreises. (German) [The Book of Scripts containing the characters and alphabets of all times and peoples of the Earth]. Olms-Weidmann, Hildesheim, Germany, 1986. ISBN 3-487-07827-9. xii + 286 pp. LCCN ????

Faulmann:1986:BSEb

[Fau86b] Carl Faulmann. Das Buch der Schrift enthaltend die Schriftzeichen und Alphabete aller Zeiten und aller Völker des Erdkreises. (German) [The Book of Scripts containing the characters and alphabets of all times and peoples of the Earth]. Eichborn, Frankfurt am Main, Germany, 1986. ISBN 3-8218-1720-8. xii + 286 pp. LCCN ????

Faulmann:2000:BSE

[Fau00] Carl Faulmann. Das Buch der Schrift enthaltend die Schriftzeichen und Alphabete aller Zeiten und aller Völker des Erdkreises. (German) [The Book of Scripts containing the characters and alphabets of all times and peoples of the Earth]. Augustus-Verlag, München, Germany, 2000. ISBN 3-8043-0374-9. xii + 286 pp. LCCN ????

Faulmann:2004:SAAa

[Fau04a] Carl Faulmann. Schriftzeichen und Alphabete aller Zeiten und Völker. (German) [Characters and alphabets of all times and peoples]. Weltbild, Augsburg, Germany, 2004. ISBN 3-8289-0799-7. xii + 286 pp. LCCN ???? EUR 9.95.

Faulmann:2004:SAAb

[Fau04b] Carl Faulmann. Schriftzeichen und Alphabete aller Zeiten und Völker. (German) [Characters and alphabets of all times and peoples]. Matrix Verlag GmbH, Wiesbaden, Germany, 2004. ISBN 3-937715-49-5. xii + 286 pp. LCCN ???? EUR 9.95.

Fournier:1985:GCI

[FB85] Alain Fournier and Brian A. Barsky. Geometric continuity with interpolating Bézier curves. Proceedings — Graphics Interface, pages 337–341, 1985. CODEN PGINEK. ISSN 0713-5424.

Felici:1991:PVT

[Fel91] James Felici. PostScript versus TrueType. Macworld, 8(9):194– 203, September 1991. CODEN MACWEA. ISSN 0741-8647.

Fenton:1989:MFB

[Fen89] Erfert Fenton. The Macintosh font book. Peachpit Press, Berkeley, Calif., 1989. ISBN 0-938151-05-3. xi + 260 pp. LCCN Z286.D47 F46 1989. US$23.95.

Fenton:1991:MFB

[Fen91] Erfert Fenton. The Macintosh font book. Peachpit Press, Berkeley, CA, USA, second edition, 1991. ISBN 0-938151-37-1. xiii + 348 pp. LCCN ????

Fenton:1994:RF

[Fen94] Erfert Fenton. The right font. Macworld, 11(3):114–??, March 1994. CODEN MACWEA. ISSN 0741-8647.

Fenton:1996:MFB

[Fen96] Erfert Fenton. The Macintosh font book. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, third edition, 1996. ISBN 0-201-88364-3 (paperback). xii + 356 pp. LCCN ????

Ferguson:1990:FLE

[Fer90] M. J. Ferguson. Fontes latines européennes et TEX 3.0 (English: European Latin fonts and TEX 3.0). Cahiers GUTenberg, 7:29–32, novembre 1990. ISSN 1140-9304.

Fischer:1985:PI

[FG85] G. Fischer and R. Gunzenhäuser.

Project INFORM. to appear , de-Gruyter Pub, 1985. Also published in/as: summary rcvd Sep.1984.

Fuchs:1982:OFC

[FK82] David R. Fuchs and Donald E.

Knuth. Optimal Font Caching. Technical Report STAN-CS-82-901, Stanford University, 1982.

Fuchs:1985:OPF

[FK85] David R. Fuchs and Donald E.

Knuth. Optimal prepaging and font caching. ACM Transactions on Programming Languages and Systems, 7(1):62–79, January 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

FloridaMarketing:1989:MA

[Flo89] Florida Marketing International.

MacIntosh advanced, 1989. 1 videocassette (ca. 100 min.).

Flowers:1994:XLF

[Flo94] Jim Flowers. X logical font description conventions: Version 1.5: Public review draft (with change bars against version 1.4): X consortium draft standard. The X Resource, Special issue D(1):G–i–G– v, G–1–G–28, February 1994. CO-DEN XRESEA. ISBN 1-56592-071-6. ISSN 1058-5591.

Findley:1978:CIP

[FLS78] G. I. Findley, D. P. Leabo, and A. C. Slutman. Control of the IBM 3800 printing subsystem. IBM Journal of Research and Development, 22(1):2–12, January 1978. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Fujieda:1985:FPLa

[FN85a] Hideo Fujieda and Hiroki Nagai. Font pattern loading method for Kanji-printers in the DIPS-106 Network OS. Denki Tsushin Kenkyusho Kenkyu Jitsuyoka Hokoku, 34(2): 317–325, 1985. CODEN DTKKAA. ISSN 0415-3200.

Fujieda:1985:FPLb

[FN85b] Hideo Fujieda and Hiroki Nagai. Font pattern loading method for Kanji printers in the DIPS-106 Network OS. Review of the Electrical Communication Laboratories (Tokyo), 33(4):725–731, July 1985. CODEN RELTAN. ISSN 0418-6338.

FontBank:1992:FTC

[Fon92] FontBank, Inc. FontBank type companion display typeface collection, 1992. 9 computer disks.

Anonymous:19xx:FF

[Fonxx] Font and flue, page various, 19xx. ????, Pattonsburg, MO, USA.

Anonymous:1991:F

[Fou91] Fount, page various, 1991. Northland Letterpress, Vancouver, BC, Canada.

Fowler:1973:PIU

[Fow73] R. L. Fowler. Printing italics using a standard font. IBM Technical Disclosure Bulletin, 15(9):??, February 1973. CODEN IBMTAA. ISSN 0018-8689.

Fox:2016:RRP

[Fox16] Margalit Fox. Rev. Robert Palladino, scribe who shaped Apple’s fonts, dies at 83. New York Times, ??(??):??, March 4, 2016. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://www.

nytimes.com/2016/03/06/arts/
design/rev-robert-palladino-
83-scribe-who-shaped-apples-
fonts.html.

Fateman:1996:PRT

[FT96] Richard J. Fateman and Taku Tokuyasu. Progress in recognizing typeset mathematics. Proceedings of SPIE — The International Society for Optical Engineering, 2660:37–50, 1996. CO-DEN PSISDG. ISBN 0-8194-2034-4. ISSN 0277-786X (print), 1996-756X (electronic).

Fuchs:1981:FMF

[Fuc81] David Fuchs. TEX Font Metric files. TUGboat, 2(1):12–16, February 1981. ISSN 0896-3207.

Furuta:1990:PIC

[Fur90] R. Furuta, editor. Proceedings of the International Conference on Electronic Publishing, Document Manipulation & Typography, Gaithersburg, Maryland, September 1990.

Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Gable:1993:CVT

[Gab93] Gene Gable. 50 cents vs US$50 typefaces. Publish!: the how-to magazine of desktop publishing, 8 (4):54–??, April 1993. CODEN PBLIEJ. ISSN 0897-6007.

Gagliano:1993:CVF

[Gag93] Joseph V. Gagliano. CUG389 — VGA FontLib, MakeFont and DXF viewer. C Users Journal, 11(12): 111–??, December 1993. ISSN 0898-9788.

Gait:1985:AAH

[Gai85] J. Gait. An aspect of aesthetics in human-computer communications: Pretty windows. IEEE Transactions on Software Engineering, SE-11(8):714–717, August 1985. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Garell:1984:DFD

[Gar84] Lynne Garell. Digital font design: Alice. Thesis (m.f.a.), Rochester Institute of Technology, Rochester, NY, USA, 1984. 25 pp.

Garfield:2010:JMT

[Gar10] Simon Garfield. Just my type: a book about fonts. Gotham Books, New York, NY, USA, 2010. ISBN 1-59240-652-1. 356 pp. LCCN Z250 .G228 2010.

Griffee:1988:ITF

[GC88] Alan W. Griffee and Cherrelyn A. Casey. An introduction to typographic fonts and digital font resources. IBM Systems Journal, 27 (2):206, 1988. CODEN IBMSA7. ISSN 0018-8670.

Geisler:2015:WLG

[Gei15] Harald Geisler. Write like a genius:

New font released on centennial of Einstein’s general theory of relativity. Scientific Computing, ??(??):??, May 7, 2015. URL http://www.

scientificcomputing.com/news/
2015/05/write-genius-new-font-
released-centennial-einsteins-
general-theory-relativity.

GenericSoftware:1987:SLTa

[Gen87a] Generic Software, Inc. Symbols library type fonts, 1987. 1 computer disk.

GenericSoftware:1987:SLTb

[Gen87b] Generic Software, Inc. Symbols library type fonts II, 1987. 1 computer disk.

GEC:1994:MP

[Gen94] General Electric Company.Computer Dept. [Miscellaneous pamphlets]. The Dept., Phoenix, AZ, USA, 1994. 35 volumes in 1.

Ghosh:1984:BDI

[Gho84] P. K. Ghosh. Basic design issues in a multi-lingual type font design and typesetting workstation. J. Inst. Electron. Telecommun. Eng, 30(6): 196–202, November 1984. CODEN JIETAU. ISSN 0377-2063.

Gibson:1990:PNA

[Gib90] Rosalind S. Gibson. Principles of Nutritional Assessment. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-505838-0. xvi + 691 pp. LCCN RC621 .G52 1990. Set by LaTEX, with graphics from Harvard Graphics and Adobe Illustrator, and designed by Ian L. Gibson and Philip Taylor with the assistance of the publisher. The book is typeset in Adobe Times Roman 10.5/12 and 8.5/10.5, with figure annotation in Adobe Helvetica. The reasons for the unusual font size are described in Philip Taylor’s TEX91 conference paper, to appear in TUGboat 12, no. 1, 1991.

Glinert:1983:LFV

[GL83] Ephraim P. Glinert and Richard E. Ladner. A large font virtual terminal interface: a software prosthesis for the visually impaired. ???? 83-01-01, Dept. of Computer Science, University of Washington, Seattle, WA, USA, 1983. 15 pp.

Glinert:1984:LFV

[GL84] Ephraim P. Glinert and Richard E. Ladner. A large font virtual terminal interface: a software prosthesis for the visually impaired. Communications of the Association for Computing Machinery, 27(6):567– 572, June 1984. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:1990:FWT

[Gla90] L. B. Glass. Font wars (Type 1 and TrueType fonts). Byte Magazine, 15(8):289–290, 292–295, August 1990. CODEN BYTEDJ. ISSN 0360-5280.

Glassner:2004:AGNa

[Gla04] Andrew Glassner. Andrew Glassner’s notebook: About face. IEEE Computer Graphics and Applications, 24(1):86–91, January/ February 2004. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://csdl.

computer.org/dl/mags/cg/2004/
01/g1086.pdf.

Glockner:1992:EVF

[Glö92] Jürgen Glöckner. Erzeugung virtueller Fonts. Die TEXnische Komödie, 4(4):9–16, December 1992. Generating virtual fonts.

Anonymous:1969:CMA

[Glu69] H. A. Glucksman. Classification of mixed-font alphabetics by characteristic loci. Technical report, ????, ????, 1969. ???? pp.

Ghosh:1983:PCG

[GM83a] P. K. Ghosh and S. P. Mudur. Parametric curves for graphic design systems. The Computer Journal, 26 (4):312–319, November 1983. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Goldberg:1983:INP

[GM83b] Jeffrey L. Goldberg and Thomas C.

Miller. Interactive name placement for provisional maps. Technical Papers of the American Congress of Surveying and Mapping, pages 314– 321, 1983. CODEN TPAMDF.

Gurtler:1985:FRM

[GM85] André Gürtler and Christian Mengelt. Fundamental research methods and form innovations in type design compared to technological developments in type production. Visible Language, XIX(1):123–147, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_
E.pdf.

Grosvenor:1990:PFH

[GMP90] Jonathan Grosvenor, Kaye Morrison, and Alexandra Pim. The PostScript font handbook 1990: the guide to PostScript Type 1 fonts. Ivanhoe Press, Oxford, England, 1990. ISBN 1-870757-23-8. ix + 257 pp. LCCN ????

Grosvenor:1992:FHA

[GMP92a] Jonathan Grosvenor, Kaye Morrison, and Alexandra Pim. The PostScript Font Handbook: a Directory of Type 1 Fonts. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-57044-0, 0-201-56893-4. x + 425 pp. LCCN Z250.7 .P67 1992. US$24.95.

Grosvenor:1992:PFH

[GMP92b] Jonathan Grosvenor, Kaye Morrison, and Alexandra Pim. The Post-Script font handbook: a directory of type 1 fonts. Addison-Wesley, Wokingham, England, revised edition, 1992. ISBN 0-201-56893-4. x + 425 pp. LCCN ????

Gonczarowski:1998:PSC

[Gon98] Jakob Gonczarowski. Producing the skeleton of a character. Lecture Notes in Computer Science, 1375:66–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/1375/13750066.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/1375/
13750066.pdf.

Goossens:1995:UAT

[Goo95] Michel Goossens. Using Adobe Type 1 Multiple Master fonts with TEX. TUGboat, 16(3):253–258, September 1995. ISSN 0896-3207.

Gossling:1986:BHR

[Gos86] T. H. Gossling. Beyond Hershey:

a representation of fonts for computer graphics. Computer Graphics Forum, 5(4):291–295, December 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gottschall:1985:SAT

[Got85] Edward Gottschall. The state of the art in typeface design protection. Visible Language, XIX(1):149–155, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_
E.pdf.

Gray:1983:TT

[Gra83] Bill Gray. Tips on Type. Van Nostrand Reinhold, New York, NY, USA, 1983. ISBN 0-442-22888-0. 128 pp. LCCN Z250.G78 1983. Little coverage of computer-based typesetting, but an excellent coverage of principles of typography, with lots of examples of fonts.

Gratzer:1993:AII

[Grä93] George Grätzer. Advances in TEX implementations. I. PostScript fonts. Notices of the American Mathematical Society, 40(7):834– 838, September 1993. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Greene:1991:HGA

[Gre91] Andrew M. Greene, Ed. How to Get Around MIT. Technology Community Association, Cambridge, MA, USA, twenty-third edition, 1991. Press run of 3000 copies; sold in local bookstore, and distributed gratis to 2000 students. Typeset with Plain TEX, using Computer Modern fonts, with markup macros written in-house. perl was used to provide a front end, consistent pagination, and the index. Final pages were prepared on 300dpi Apple LaserWriter Plus printer.

Gregory:1992:RKN

[Gre92] Peter Gregory. A reply to Kalb and Neves. Supercomputing Review, 5(5):62–??, May 1992. CO-DEN SURVEG. ISSN 1048-6836.

Green:1995:TDT

[Gre95] Barry Green. Typography: 1995 directory of type foundries. Publish!: the how-to magazine of desktop publishing, 10(10):73–??, October 1995. CODEN PBLIEJ. ISSN 0897-6007.

Green:1996:PW

[Gre96] Barry Green. Product watch. Publish!: the how-to magazine of desktop publishing, 11(3):19–??, March 1996. CODEN PBLIEJ. ISSN 0897-6007.

Garfinkel:2021:NCM

[GS21] Simson Garfinkel and Eugene H.

Spafford. News: Charles M. Geschke (1939–2021). Communications of the Association for Computing Machinery, 64(7):22, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3467481.

Grebert:1992:CGP

[GSKM92] Igor Grebert, David G. Stork, Ron Keesing, and Steve Mims. Connectionist generalization for production: an example from gridfont. Neural Networks, 5(4):699– 710, July–August 1992. CODEN NNETEB. ISSN 0893-6080.

Geyer-Schulz:1993:TAA

[GSMN93] A. Geyer-Schulz, J. Matulka, and G. Neumann. Typesetting APL with LaTEX–the APL front end APL2TEX. APL Quote Quad, 23 (3):22–27, March 1993. CODEN APLQD9. ISSN 0163-6006.

Guntermann:1985:GDL

[Gun85] Klaus Guntermann. GENDRIV— a driver for low cost devices using built-in fonts. In Lucarella [Luc85], pages 197–204. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Goossens:1992:STP

[GV92] M. Goossens and E. Van Herwijnen.

Scientific text processing. International Journal of Modern Physics C [Physics and Computers], 3(3):479– 546, June 1992. CODEN IJMPEO.

ISSN 0129-1831 (print), 1793-6586 (electronic).

Gurari:1991:ACS

[GW91] E. M. Gurari and J. Wu. A WYSIWYG literate programming system (preliminary report). In 1991 ACM Computer Science Conference: March 5–7, 1991, San Antonio Convention Center, San Antonio, Texas: Proceedings: Preparing for the 21st Century, pages 94–

104. ACM Press, New York, NY, USA, March 1991. ISBN 0-89791-382-5. LCCN QA75.5 .A137 1991.

Grieger:1984:ICM

[GWW84] I. Grieger, G. Wagemann, and G. Wu. Interactive creation and modification of Chinese characters. Computers and Graphics, 8(1):81– 92, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Haley:2007:HZT

[Hal07] Allan Haley. Hermann Zapf: Three stories of point & counterpoint. Step Inside Design, ??(??): ??, September/October 2007. URL

https://web.archive.org/web/
20081106134042/https://www.stepinsidedesign.
com/STEPMagazine/Article/28799.

Hammouche:1990:AVL

[Ham90] Rabah Hammouche. Atlas Vent de l’Algérie. Office National de la Météorologie, Alger, Algérie, 1990. 150 pp. Typeset with Bitstream fonts on a Canon Series III 300-dpi laser printer, with extensive graphics and data tables; the data files are available on IBM PC diskettes.

Haralambous:1992:TFM

[Har92] Yannis Haralambous. A typewriter font for the Macintosh 8-bit font table. TUGboat, 13(4):476–477, December 1992. ISSN 0896-3207.

Haralambous:1993:PPF

[Har93] Yannis Haralambous. Parametrization of PostScript fonts through —an alternative to Adobe Multiple Master Fonts. Electronic Publishing— Origination, Dissemination, and Design, 6(3):145–157, September 1993. CODEN EPODEU. ISSN 0894-3982.

Haralambous:1994:TK

[Har94] Yannis Haralambous. Typesetting Khmer. Electronic Publishing —Origination, Dissemination, and Design, 7(4):197–215, December 1994. CODEN EPODEU. ISSN 0894-3982.

Haralambous:1996:FUD

[Har96] Yannis Haralambous. ΩTimes and ΩHelvetica fonts under development: Step One. TUGboat, 17(2): 126–146, June 1996. ISSN 0896-3207.

Haralambous:1998:SAS

[Har98] Yannis Haralambous. Simplification of the Arabic script: Three different approaches and their implementations. Lecture Notes in Computer Science, 1375:138–156, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/1375/13750138.htm;   http:

                                        //link.springer-ny.com/link/
                                        service/series/0558/papers/1375/
                                        13750138.pdf.

Haveman:1993:CTP

[Hav93] Josepha Haveman. CD-ROM treasure pack, 1993. ISBN 1-56737-080-

2. 10 computer laser optical discs

guides.

Hersch:1991:MMH

[HB91] Roger D. Hersch and Claude Bétrisey. Model-based matching and hinting of fonts. Computer Graphics, 25(4):71–80, July 1991. CODEN CGRADI, CPGPBZ. ISSN 0097-8930. URL http://www.

                                        acm.org:80/pubs/citations/proceedings/
                                        graph/122718/p71-hersch/.

Hersch:1995:PTG

[HBBG95] Roger D. Hersch, Claude Bétrisey, Justin Bur, and André Gürtler. Perceptually tuned generation of grayscale fonts. IEEE Computer Graphics and Applications, 15 (6):78–90 (or 78–89??), November 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Holder:1988:FSC

[HD88a] S. Holder and J. Dengler. Fontand sizeinvariant character recognition with greyvalue image features. In IEEE [IEE88], pages 252–254.

Holder:1988:FSI

[HD88b] S. Holder and J. Dengler. Fontand size invariant character recognition with greyvalue image features. Proceedings — International Conference on Pattern Recognition, pages 252–254, 1988. CODEN PI-CREG. ISBN 0-8186-0878-1. Available from IEEE Service Cent (catalog number 88CH2614-6). Piscataway, NJ, USA.

Hayashi:1968:GIN

[HDK68] H. Hayashi, S. Duncan, and S. Kuno. Graphical input of non-standard characters. Communications of the Association for Computing Machinery, 11(9):613– 618, September 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Headley:2005:EF

[Hea05] Gwyn Headley. The encyclopaedia of fonts. Cassell Illustrated, London, UK, 2005. ISBN 1-84403-206-X. 496 pp. LCCN M07.G00913.

Heid:1995:WS

[Hei95] Jim Heid. Working smart. Macworld, 12(5):129–??, May 1995. CO-DEN MACWEA. ISSN 0741-8647.

Henderson:1989:OF

[Hen89] Doug Henderson. Outline fonts with . TUGboat, 10(1): 36–38, April 1989. ISSN 0896-3207.

Hershey:1967:CC

[Her67] Allen V. Hershey. Calligraphy for computers. Technical Report TR-2101, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, August 1967.

Hershey:1969:FIP

[Her69] Allen V. Hershey. FORTRAN IV programming for cartography and typography. Technical Report TR-2339, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1969.

Hershey:1970:PRF

[Her70] Allen V. Hershey. Preparation of reports with the FORTRAN typographic system. Technical Report TN-K/27-70, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1970.

Hershey:1972:CSS

[Her72] Allen V. Hershey. A computer system for scientific typography. Computer Graphics and Image Processing, 1:373–385, 1972. CODEN CGIPBG. ISSN 0734-189X.

Hertzfeld:1979:TFE

[Her79] Andrew Hertzfeld. Two font editors. Master’s thesis, University of California, Berkeley, Berkeley, CA, USA, 1979. various pp. Master of Science, Plan II.

Hershey:1981:ACT

[Her81] Allen V. Hershey. Advanced computer typography. Technical Report NPS012-81-005, U. S. Naval Postgraduate School, Monterey, CA 93940, USA, December 1981.

Hersch:1987:CGU

[Her87] Roger D. Hersch. Character generation under grid constraints. Computer Graphics, 21(4):243–252, July 1987. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

URL http://www.acm.org:80/
pubs/citations/proceedings/graph/
37401/p243-hersch/.

Hersch:1989:IFR

[Her89] R. Hersch. Introduction to font rasterization. In André and Hersch [AH89], pages 1–13. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Hershey:1995:CTT

[Her95] A. V. Hershey. Cartography and typography in True BASIC. Technical Report NPS-09-95-003, Naval Postgraduate School, Monterey CA, USA, September 1995.

HP:1985:CP

[Hew85] Hewlett–Packard Company. Computer printers, 1985. 2 videocassettes (65 min.) (VHS format) workbook.

HP:1988:LPF

[Hew88] Hewlett–Packard Company. Laser-Jet printer family font catalog. Hewlett–Packard, Boise, Idaho?, 1988. various pp.

HP:1989:HLP

[Hew89] Hewlett–Packard Company. HP LaserJet ProCollection font cartridge, 1989. 1 font cartridge + 3 manuals (Using Hewlett–Packard’s AutoFont support; AutoFont support, software application notes; Using the Hewlett–Packard ProCollection font cartridge).

Hasegawa:1996:VTF

[HFW^{+}96] Susumu Hasegawa, Masayuki Fujisawa, Hajime Watanabe, Yoshiko Imaki, Sachiko Iida, and Daisuke Matsunaga. Variety Touch Font: a font generating system. Shapu Giho/Sharp Technical Journal, 64:

27–30, April 1996. CODEN STEJD9. ISSN 0285-0362.

Hobby:1983:CM

[HG83] John Hobby and Gu Guoan. A Chinese Meta-Font. Technical report STAN-CS-83-974, Stanford University, Department of Computer Science, 1983.

Hobby:1984:CMF

[HG84] John D. Hobby and Gu Guoan. A Chinese meta-font. TUGboat, 5(2): 119, November 1984. ISSN 0896-3207.

Handa:1983:LP

[HH83] Michael K. Handa and Jonathan A. Hunt. ‘Typesetting’ with low-cost printers. High Technology (Boston), 3(11):24, 26, 28–??, November 1983. CODEN HTECD3.

Hopkins:1986:CFM

[HH86] Richard L. Hopkins and Hill & Dale Private Press & Type Foundry. Casting a font of metal type, 1986. 1 videocassette (VHS) (109 min.).

Herz:1994:TUA

[HH94] J. Herz and R. D. Hersch. Towards a universal auto-hinting system for typographic shapes. Electronic Publishing—Origination, Dissemination, and Design, 7(4):251– 260, December 1994. CODEN EPODEU. ISSN 0894-3982.

Hu:2001:PFB

[HH01] Changyuan Hu and Roger D. Hersch. Parameterizable fonts based on shape components. IEEE Computer Graphics and Applications, 21 (3):70–85, May/June 2001. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

URL http://computer.org/cga/
cg2001/g3070abs.htm;   http:
//dlib.computer.org/cg/books/
cg2001/pdf/g3070.pdf.

Herz:1997:CPC

[HHG97] Jacky Herz, Roger D. Hersch, and Jakob Gonczarowski. Coherent processing of character skeletal forms. Computers and Graphics, 21(6):727–736, November–December 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

URL http://www.elsevier.com/
cgi-bin/cas/tree/store/cag/cas_
sub/browse/browse.cgi?year=1997&
volume=21&issue=6&aid=9700050.

Hertz:1998:WBM

[HHGH98] Jacky Hertz, Changyuan Hu, Jakob Gonczarowski, and Roger D. Hersch. A window-based method for automatic typographic parameter extraction. Lecture Notes in Computer Science, 1375:44– 54, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/
series/0558/bibs/1375/13750044.
htm;   http://link.springer-
ny.com/link/service/series/0558/
papers/1375/13750044.pdf.

Hohti:1987:GAF

[HK87] Aarno Hohti and Okko Kanerva.

Generating an APL font. TUGboat, 8(3):275–278, November 1987. ISSN 0896-3207.

Heilman:1978:TOT

[HM78] Robert L. Heilman and Jean M. Marchant. TIGS — an overview of the Terminal Independent Graphics Systems. Computer Graphics, 12 (3):293–297, August 1978. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Heesch:1964:WF

[HN64] Henry J. Heesch and Spencer Nesbit. Wrong font. Dept. of Printing Engineering and Management, California Polytechnic State University, San Luis Obispo, CA, USA, 1964. 11 pp.

Hobby:1993:GAT

[Hob93] John D. Hobby. Generating automatically tuned bitmaps from outlines. Journal of the ACM, 40(1): 48–94, January 1993. CODEN JA-COAH. ISSN 0004-5411.

Hoenig:1994:FR

[Hoe94] Alan Hoenig. {Meta}Font Forum redux. TUGboat, 15(2):97, June 1994. ISSN 0896-3207.

Hoenig:1995:PFF

[Hoe95] Alan Hoenig. The Poetica family: fancy fonts with TEX and LaTEX. TUGboat, 16(3):244–252, September 1995. ISSN 0896-3207.

Hoffman:1983:FF

[Hof83] Paul E. Hoffman. Fancy font. Byte Magazine, 8(10):??, October 1983. CODEN BYTEDJ. ISSN 0360-5280.

Holmes:1980:IZC

[Hol80] Kris Holmes. ITC Zapf Chancery. Fine Print, 6(1):29–??, ???? 1980.

Holzner:1987:DWA

[Hol87] Steven Holzner. Desktop workshop 10 assembly language utilities that extend your computing environment, 1987. ISBN 0-13-196214-0, 0-13-188905-2 (reference manual). 1 computer disk.

Holmes:2015:RHZ

[Hol15] Kris Holmes. Remembering Hermann Zapf. Bigelow Holmes (blog), June 2015. URL http://

bigelowandholmes.typepad.com/
bigelow-holmes/2015/06/remembering-
hermann-zapf.html.

Horn:1993:WMF

[Hor93] Berthold K. P. Horn. Where are the math fonts? TUGboat, 14(3):282– 284, October 1993. ISSN 0896-3207.

Hossain:2024:TSR

[Hos24] Anushah Hossain. Text standards for the “rest of world”: The making of the Unicode standard and the OpenType format. IEEE Annals of the History of Computing, 46(1):20– 33, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Howland:1986:TAU

[How86] John E. Howland. Typesetting APL using a Macintosh. ACM SIGAPL APL Quote Quad, 16(4):301–305, 1986. ISBN 0-901865-35-4.

Hussain:1993:ROF

[HP93] Fiaz Hussain and Michael L. V. Pitteway. Rasterizing the outlines of fonts. Electronic Publishing—Origination, Dissemination, and Design, 6(3):171–181, September 1993. CO-DEN EPODEU. ISSN 0894-3982.

Haralambous:1998:DUM

[HP98] Yannis Haralambous and John Plaice. The design and use of a multiple-alphabet font with Ω. Lecture Notes in Computer Science, 1375:126–137, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/1375/13750126.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/1375/
13750126.pdf.

Hudson:2000:WGP

[Hud00] John Hudson. Windows glyph processing. World-Wide Web document, November 7, 2000. URL

http://www.microsoft.com/typography/
developers/opentype/default.
htm.

Huggins:1988:EHM

[Hug88] K. Cleo R. Huggins. Egyptian hieroglyphs for modern printing devices: an outline font of Egyptian hieroglyphs for PostScript printers. Thesis, Department of Computer Science, Stanford University, Stanford, CA, USA, June 1,

1988. 61 pp. URL https://apps.
dtic.mil/dtic/tr/fulltext/u2/
a326695.pdf.

Hussain:1995:CRC

[Hus95] F. Hussain. On computerisation and representation of characters. IEE Colloquium (Digest), 191:11/1–11/9, 1995. CODEN DCILDN. ISSN 0963-3308.

Hustwit:2007:HDF

[Hus07] Gary Hustwit. Helvetica: a documentary film. Feature-length independent film., 2007. URL http://

www.helveticafilm.com/.

Huang:1992:AFS

[HWY92] Y. Huang, X. Wang, and C. Yuan.

An algorithm for font scaling and generating from a kind of curvecontoured Chinese character fonts. In Tou and Liang [TL92a], pages 191–196.

Hirose:1989:BAW

[HYH89] Yoshio Hirose, Koichi Yamashita, and Shimpei Hijiya. Backpropagataion algorithm which varies the number of hidden units. In International Joint Conference on Neural Networks — volume II (Jun 18–22 1989: Washington, DC, USA), pages 625–?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. IEEE catalog number 89CH2765-6.

Hirose:1991:BAW

[HYH91] Yoshio Hirose, Koichi Yamashita, and Shimpei Hijiya. Backpropagation algorithm which varies the number of hidden units. Neural Networks, 4(1):61–66, 1991. CO-DEN NNETEB. ISSN 0893-6080.

HyperActiveSoftware:1990:FSC

[Hyp90] HyperActive Software. The font sampler catalog. HyperActive Software, Minneapolis, MN, USA, 1990. various pp.

Hu:1993:AHC

[HZ93] C. Hu and F. Zhang. Automatic hinting of Chinese outline font based on stroke separating method. In Shin and Kunii [SK93], pages 359–368. ISBN 981-02-1529-0. LCCN T385.C5575 1993.

Ip:1987:CCW

[IBC87] Chok-Ho Ip, Daniel M. Berry, and Kam Pui Chow. Cwpr, a Chinese/Japanese Word-Processing System for Use with the UNIX Device-Independent Troff System. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. ISBN 0-8186-0780-7. 570–577 pp. LCCN QA75.5 .I5761 1987. IEEE Service Cent. Piscataway, NJ, USA.

ICCPCOL:1988:ICC

[ICC88] 1988 International Conference on Computer Processing of Chinese and Oriental Languages, August 29–September 1, 1988, Toronto, Canada. Concordia University, Montréal, Québec, Canada, 1988. LCCN QA76.9.T48 I48 1988.

IEEE:1986:PEI

[IEE86] IEEE, editor. Proceedings, Eighth International Conference on Pattern Recognition (Paris, France, October 27–31, 1986), IEEE Publ. 86CH2342-4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. CODEN PICREG. ISBN 0-8186-0742-4. LCCN ????

ICPR:1988:NIC

[IEE88] IEEE, editor. Ninth International Conference on Pattern Recognition (Rome, Italy, November 14– 17, 1988). Computer Society Press, Washington, DC, 1988.

IEEE:1992:IIC

[IEE92] IEEE, editor. 11th IAPA International Conference on Pattern Recognition: August 30–September 3, 1992, Hague, The Netherlands. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2930-4, 0-8186-2917-7, 0-8186-2915-0. ISSN 1051-4651. LCCN Q 327 I615 1992. Four volumes.

IEEE:1993:IPN

[IEE93] IEEE, editor. IPCC 93 proceedings:

The new face of technical communication: people, processes, products: October 5–8, 1993, Philadelphia, PA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1466-2, 0-7803-1465-4, 0-7803-1467-0. LCCN T 10.5 I54 1993.

Ion:1988:TUI

[II88] Patrick D. F. Ion and Bonita D. Ross Ion. TEX and its uses for information presentation. Bulletin of the American Society for Information Science, 14(4):21–22, April– May 1988. CODEN BASICR. ISSN 0095-4403.

Ikedo:1995:SHG

[Ike95] Tsuneo Ikedo. Scalable highperformance graphics processor: GVIP. Visual Computer, 11(3): 121–133, 1995. CODEN VICOE5. ISSN 0178-2789.

Ishii:1982:HED

[IKM82] Atsushi Ishii, Kenji Kouno, and Yasunori Maezawa. Highly efficient data compression method for newspaper image data. Fujitsu scientific and technical journal, V 18(N 2):199–225, June 1982. CODEN FUSTA4. ISSN 0016-2523.

IBM:1982:IID

[Int82] International Business Machines Corporation. IBM 6670 information distributor: Model II font editing system. Field developed program: Program description/operations manual. IBM Corporation, Southfield, MI, USA, 1982. viii + 48 pp.

IBM:1983:IFE

[Int83] International Business Machines Corporation. IBM 6670 font editing system. IBM system/370, 43XX processors, 30XX processors Program offering. IBM Corporation, Irving, Tx., 1983. 4 pp.

Nesterov:1991:IC

[INT91] Vyacheslav M. Nesterov, editor. Interval Computations, 1991. ISSN 0135-4868. Box 52, St.Petersburg 195256, Russia, E-mail: nest@nit. spb.su (Internet). A refereed international journal, Interval Computations is the only periodical in the world devoted specifically to various aspects of reliable numerical computations based on the interval approach. It is managed by an international editorial board from Bulgaria, Germany, Japan, Russia, and the United States, and printed in Russia. The journal includes various items in the fields of theoretical research, computer tools, applications, interdisciplinary research and other relevant areas. The IC has been prepared with TEX since 1992. Recently a new style has been designed based on standard LaTEXarticle style with changes in fonts, headings layout, running heads etc. It uses 10.8pt Baskerville for body text (the unusual font size is selected for better wedding with CM math) and horizontally-squeezed Palatino for headings. For mathematical formulae, standard Computer Modern is preserved. Proofs are made using emTEX DVIHPLJ on a Hewlett– Packard LaserJet IIIp, and final pages are printed at 600dpi. Using VTEX is now under consideration. Papers are accepted in any TEX variety; LaTEXarticle.sty preferred.

Itoh:1993:CFA

[IO93] Koichi Itoh and Yoshio Ohno. A curve fitting algorithm for character fonts. Electronic Publishing—Origination, Dissemination, and Design, 6(3):195, September 1993. CODEN EPODEU. ISSN 0894-3982.

Ip:1994:FCC

[IWM94] Horace H. S. Ip, Helena T. F. Wong, and Florence Y. Mong. Fractal coding of Chinese scalable calligraphic fonts. Computers and Graphics, 18 (3):343–351, May–June 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jackson:1982:FCP

[Jac82] Calvin Jackson. Font codes in popular use. TUGboat, 3(2):26, October 1982. ISSN 0896-3207.

Jaeschke:1991:DSC

[Jae91] Rex Jaeschke. The Dictionary of Standard C. Professional Press Books, 101 Witmer Road, Horsham, PA 19044, USA, 1991. ISBN 1-878956-07-8. x + 164 pp. LCCN QA76.73.C15 J335 1991. Prepared with LaTEX, with output to Personal TEX’s 2000dpi phototypesetter using Computer Modern fonts.

Jantz:1990:AP

[Jan90] Richard Jantz. Art of persuasion. PC World, 8(10):259–??, October 1990. CODEN PCWDDV. ISSN 0737-8939.

JSA:1978:CJG

[Jap78] Japanese Standards Association, 1-24, Akasaka 4 Chome, Minatoku, Tokyo, 107 Japan. Code of the Japanese Graphic Character Set for Information Interchange, JIS-C-6226 (1978), 1978.

Jeffrey:1993:FIP

[Jef93a] Alan Jeffrey. A PostScript font installation package written in TEX. TUGboat, 14(3):285–292, October 1993. ISSN 0896-3207.

Jeffrey:1993:MFE

[Jef93b] Alan Jeffrey. Math font encodings: a workshop summary. TUGboat, 14 (3):293–295, October 1993. ISSN 0896-3207.

Jeffrey:1994:FS

[Jef94] Alan Jeffrey. PostScript font support in LaTEX2e. TUGboat, 15(3): 263–268, September 1994. ISSN 0896-3207.

Jensen:1993:ISA

[JN93] R. S. Jensen and D. Neumeister, editors. 7th International symposium — 1993 Apr: Columbus; OH, volume 7 of PROCEEDINGS OF THE SYMPOSIUM ON AVI-ATION PSYCHOLOGY TH//V1. ????, ????, 1993.

Joch:1990:NNL

[Joc90] A. Joch. No-muss, no-fuss, low-cost PostScript printer. Byte Magazine, 15(8):144–148, August 1990. CO-DEN BYTEDJ. ISSN 0360-5280.

Jones:1994:FSS

[Jon94] Gerald E. Jones. Font secrets and solutions. SYBEX, San Francisco, 1994. ISBN 0-7821-1246-3. xxviii + 384 pp. LCCN Z250.7 .J66 1994. US$21.99.

Jones:1995:THA

[Jon95] Gerald E. Jones. Tzu hsing ao mi chi chieh ta. Tien tzu kung yeh chu pan she, Pei-ching shih, ti 1 pan. edition, 1995. ISBN 7-5053-2782-8. 4 + iv + 216 pp. LCCN ????

Jones:1978:BDM

[JSG78] Dan Jones, Don Stupeck, and Gerald Gallagher. Brushless DC motor font drive. In Proceedings of the Annual Symposium on Incremental Motion Control System Devices, 7th, Chicago, IL, USA, May 24–26 1978, pages 189–202. Incremental Motion Control Syst Soc, Urbana-Champaign, Ill, 1978. CO-DEN IMCSD3.

Kegler:2002:IFC

[K^{+}02] Richard Kegler et al., editors. Indie fonts: a compendium of digital type from independent foundries. P-Type Publications, Buffalo, NY, USA, 2002. ISBN 0-9631082-2-0. 408 pp. LCCN Z250.7 .I53 2002. See also volume 2 [KGR03] and volume 3 [KGR07].

Kadavy:2011:DHR

[Kad11] David Kadavy. Design for Hackers: Reverse Engineering Beauty. John Wiley and Sons, Inc., Chichester, UK, 2011. ISBN 1-119-99895-6. 352 (est.) pp. LCCN ????

Kane:1984:MMI

[Kan84] Michael Christopher Kane. Moment methods for the identification of typewriting equipment from same font typewritten evidence. Thesis (m.s.), Ohio State University, Columbus, OH, USA, 1984. 208. pp.

Karow:1989:AHI

[Kar89] P. Karow. Automatic hinting for intelligent font scaling. In André and Hersch [AH89], pages 232–241. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Karow:1991:DPC

[Kar91] Peter Karow. Digital punch cutting. Electronic Publishing—Origination, Dissemination, and Design, 4(3):151–170, September 1991. CO-DEN EPODEU. ISSN 0894-3982.

Karow:1992:DSD

[Kar92] Peter Karow. Digitale Schriften, Darstellung und Formate, Geleitwort von Hermann Zapf (Digital Fonts, Representation and Formats, Forward by Hermann Zapf). Springer-Verlag and URW-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc. and Hamburg, Germany, 1992. xiii + 457, with 230 illustrations pp. English translation in [Kar94a].

Karow:1994:DTD

[Kar94a] Peter Karow. Digital typefaces: description and formats. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-56509-4. xiii + 448 pp. LCCN Z250.7 .K37 1994. Foreword by Hermann Zapf. Translated by Bill Horton from German edition [Kar92].

Karow:1994:FTD

[Kar94b] Peter Karow. Font Technology:

Description and Tools. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-57223-6. xvi + 460 pp. US$79.00.

Karow:1994:FTM

[Kar94c] Peter Karow. Font technology:

methods and tools. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-57223-6 (Berlin), 0-387-57223-6 (New York). xiv + 457 pp. LCCN Z250.7 .K3813 1994.

Karow:1998:TDT

[Kar98] Peter Karow. Two decades of typographic research at URW: a retrospective. Lecture Notes in Computer Science, 1375:265–280, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/1375/13750265.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/1375/
13750265.pdf.

Kay:1985:PPP

[Kay85] David G. Kay. Programming for People/Pascal. Mayfield, Palo Alto, CA, USA, 1985. ISBN 0-87484-717-6. xiii + 620 pp. LCCN QA76.73.P3 K3. Prepared with TEX using Baskerville fonts, with output to Stanford Autologic APS-5 phototypesetter.

Kay:1996:AFM

[Kay96] Russell Kay. Adobe’s finally the managing type — get unified font management on the Mac and Windows PCs with Adobe’s new ATM 4.0. Byte Magazine, 21(9):42–??, September 1996. CODEN BYT-EDJ. ISSN 0360-5280.

Kerrick:1986:ERO

[KB86] David D. Kerrick and Alan C. Bovik. Efficient recognition of omni-font characters using models of human pattern perception. In Proceedings of the 1986 IEEE International Conference on Systems, Man, and Cybernetics, Atlanta, GA, USA, pages 1287–1291. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. IEEE Service Cent. Piscataway, NJ, USA.

Kelly:2016:NRH

[KB16] Jerry Kelly and Misha Beletsky.

The Noblest Roman: a History of the Centaur Types of Bruce Rogers. David R. Godine, Publisher, in association with Sherwin Beach Press, Boston, MA, USA, 2016. ISBN 1-56792-582-0. 120 pp. LCCN Z250.5.C44 K45 2016.

Kernighan:1975:STM

[KC75] B. W. Kernighan and L. L. Cherry.

A system for typesetting mathematics. Communications of the Association for Computing Machinery, 18(??):151–157, ?? 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kindersley:1987:DKW

[KC87] David Kindersley and Lida Lopes Cardozo. David Kindersley’s workshop. World-Wide Web document, 1987. URL http://www.

icce.rug.nl/erikjan/bluefuzz/
David-Kindersley/David-Kindersley/
David-Kindersley.html.

Kraus:1994:SMC

[KD94] Eugene J. Kraus and Edward R.

Dougherty. Segmentation-free morphological character recognition. Proceedings of SPIE — The International Society for Optical Engineering, 2181:14–23, 1994. CO-DEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic).

Kelly:2011:AMA

[Kel11] Jerry Kelly. About more alphabets: the types of Hermann Zapf, volume 3 of Typophile chap book. N.S. The Typophiles, New York, NY, USA, 2011. ISBN 0-9842744-0-5. 111 pp. LCCN ???? With a foreword by Robert Bringhurst.

Kelly:2019:HZW

[Kel19a] Jerry Kelly. Hermann Zapf and the world he designed. The Kelly-Winterton Press, New York, NY, USA, 2019. ISBN 0-578-42521-1 (hardcover). 364 pp. LCCN Z250.A2 Z3754 2019.

Kelly:2019:TRW

[Kel19b] Jerry Kelly. Type Revivals: What Are They? Where Did They Come From? Where Are They Going? David R. Godine, Publisher, Boston, MA, USA, 2019. ISBN 1-56792-647-9 (hardcover). ???? pp. LCCN Z250 .K446

                                        2019. URL http://www.godine.
                                        com/book/type-revivals/.

Kenney:1986:EFS

[Ken86] Michael Joseph Kenney. The effect of font, slant, and dot composition upon dot-matrix letter recognition. Thesis (m.a. in psychology), [s.n.], Dayton, OH, USA, 1986. viii + 36 pp.

Kay:1977:PDM

[KG77] A. Kay and A. Goldberg. Personal dynamic media. Computer, 10(3):31–41, March 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kegler:2003:IFC

[KGR03] Richard Kegler, James Grieshaber, and Tamye Riggs, editors. Indie fonts 2: a compendium of digital type from independent foundries. Rockport Publishers, Gloucester, MA, 2003. ISBN 1-59253-124-

5. 400 pp. LCCN Z250.7 .I535 2003. URL http://www.

loc.gov/catdir/enhancements/
fy0740/2002117837-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0740/2002117837-d.html. See

also volume 1 [K^{+}02] and volume 3 [KGR07].

Kegler:2007:IFC

[KGR07] Richard Kegler, James Grieshaber, and Tamye Riggs, editors. Indie fonts 3: a compendium of digital type from independent foundries. Rockport Publishers, Glouster, MA, 2007. ISBN 1-59253-313-

2. 414 pp. LCCN Z250.7 .I536 2007. URL http://www.

loc.gov/catdir/enhancements/
fy0835/2006928773-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0837/2006928773-d.html. See

also volume 1 [K^{+}02] and volume 2 [KGR03].

Khoubyari:1994:FIU

[KH94] Siamak Khoubyari and Jonathan J.

Hull. Font identification using visual global context. In Vincent and Pavlidis [VP94b], pages 116–124. CODEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TK7895.O6 D63 1994.

Khoubyari:1996:FFW

[KH96] Siamak Khoubyari and Jonathan J. Hull. Font and function word identification in document recognition. Computer Vision and Image Understanding: CVIU, 63(1):66–74, January 1996. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Kim:1981:ICC

[Kim81] Scott Kim. Inversions: a catalog of calligraphic cartwheels. BYTE Books, Peterborough, NH, USA, 1981. ISBN 0-07-034546-5. 122 pp. LCCN N7430.5 .K5. Foreword by Douglas R. Hofstadter and backword by Jef Raskin.

Kim:1990:DSH

[Kim90] Kyongsok Kim. Databases supporting Hangul (or Korean script). Proceedings of the IEEE International Conference on Systems, Man and Cybernetics, pages 485– 490, November 1990. CODEN PICYE3. ISBN 0-87942-597-0. ISSN 0884-3627. IEEE catalog number 90CH2930-6.

King:2005:NFC

[Kin05] Emily King. New faces (chapter 2: The west coast). Typotheque, ??(??):??, January 23, 2005.

                                        URL https://www.typotheque.
                                        com/articles/new_faces_chapter_
                                        two_the_west_coast.

Kirschenbaum:2005:GGH

[Kir05] Valerie Kirschenbaum. Goodbye Gutenberg: hello to a new generation of readers and writers. Designer Writers series. Global Renaissance Society, New York, NY, USA, 2005. ISBN 0-9745750-3-8. 416 pp. LCCN Z116.A3

K57 2005. URL http://www.
goodbyegutenberg.com/.

Kitamura:1984:DFG

[Kit84] J. Kitamura. Draft font generation. IBM Technical Disclosure Bulletin, 27(4B):2504–??, September 1984. CODEN IBMTAA. ISSN 0018-8689.

Kim:1989:MFG

[KKL89] Cha Jong Kim, Han Jong Kang, and Soo Youn Lee. A multipurpose font generation system. Journal of the Korea Information Science Society = Chongbo Kwahakhoe nonmunji, 16(5):422–433, September 1989. CODEN HJKHDC. ISSN 0258-9125.

Kawata:1993:OFR

[KKM^{+}93] Tetsuro Kawata, Kenich Kawachi, Nobuaki Miyakawa, Ichiro Kawazome, Hiromi Ysumatsu, Susumu Haga, and Masaya Takenaka. Outline font rendering processor with an embedded RISC CPU for highspeed hint processing. Proceedings of the Custom Integrated Circuits Conference, pages 12.5.1–12.5.4, 1993. CODEN PCICER. ISBN 0-7803-0826-3. ISSN 0886-5930. IEEE catalog number 93CH3214-4.

Kawata:1994:OFR

[KKM^{+}94] Tesuro Kawata, Kenichi Kawauchi, Nobuaki Miyakawa, Ichiro Kawazome, Hiromi Yasumatsu, Susmumu Haga, and Masaya Takenaka. An outline font rendering processor with an embedded RISC CPU for high-speed hint processing. IEEE Journal of Solid-State Circuits, 29(3):280–289, March 1994. CODEN IJSCBC. ISSN 0018-9200.

Klassen:1993:VWS

[Kla93] R. Victor Klassen. Variable width splines: a possible font representation? Electronic Publishing—Origination, Dissemination, and Design, 6(3):183–194, September 1993. CO-DEN EPODEU. ISSN 0894-3982.

Klassen:1994:EIH

[Kla94] R. Victor Klassen. Exact integer hybrid subdivision and forward differencing of cubics. ACM Transactions on Graphics, 13 (3):240–255, July 1994. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0730-
0301/197476.html.

Krishnan:1983:ACG

[KM83] R. K. Krishnan and C. S. Moghe. An approach to character generation using cubic splines. IEEE Transactions on Consumer Electronics, CE-29:438–442, 1983. CO-DEN ITCEDA. ISSN 0098-3063.

Kobayashi:1987:MHL

[KM87] Makoto Kobayashi and Yoshiaki Mima. Multi-Font Handling for Large Character Set on Workstations. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. ISBN 0-8186-0770-X. 2–8 pp. LCCN HF 5548.2 I15 1987. IEEE Service Cent. Piscataway, NJ, USA.

Kawabata:1989:FCD

[KM89] Atsushi Kawabata and Dan C.

Marinescu. Font cache design issues in a distributed electronic publishing system for Japanese language documents. Technical Report SERC-TR-55-P, Software Engineering Research Centre, September 1989.

Kai:1989:HSO

[KMNO89] Naoyuki Kai, Tsutomu Minagawa, Ichiro Nagashima, and Masahide Ohhashi. High speed outline font rasterizing LSI. Proceedings of the Custom Integrated Circuits Conference, pages 24.6/1–4, May 1989. CODEN PCICER. ISSN 0886-5930. Available from IEEE Service Cent. Piscataway.

Kohen:1994:TDF

[KMU94] Eliyezer Kohen, Microsoft Corporation, and University Video Communications. TrueType the digital font technology, 1994. 1 videocassette (48 min.).

Knapp:1986:FW

[Kna86] Walter W. Knapp. The font workshop, 1986. 1 program file (Macintosh) on 1 computer disk manual.

Knappen:1993:FAF

[Kna93] Jörg Knappen. Fonts for Africa: the fc-fonts. TUGboat, 14(2):104–106, July 1993. ISSN 0896-3207.

Knappen:1995:RDF

[Kna95] Jörg Knappen. Release 1.2 of the dc-fonts: Improvements to the European letters and first release of text companion symbols.

TUGboat, 16(4):381–387, December 1995. ISSN 0896-3207.

Knappen:1996:TDF

[Kna96] Jörg Knappen. The dc fonts 1.3: Move towards stability and completeness. TUGboat, 17(2):99–101, June 1996. ISSN 0896-3207.

Kakugawa:2001:GPF

[KNT^{+}01] H. Kakugawa, M. Nishikimi, N. Takahashi, S. Tomura, and K. Handa. A general purpose font module for multilingual application programs. Software—Practice and Experience, 31(15):1487–1508, ???? 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

                                        /www3.interscience.wiley.com/
                                        cgi-bin/abstract/86511587/START;
                                         http://www3.interscience.wiley.
                                        com/cgi-bin/fulltext?ID=86511587&
                                        PLACEBO=IE.pdf.

Knuth:1979:TMN

[Knu79] Donald E. Knuth. TEX and - : New Directions in Typesetting. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1979. ISBN 0-932376-02-9. xi + 201 + 105 pp. LCCN Z253.3 .K58 1979.

Knuth:1981:CM

[Knu81] Donald E. Knuth. The concept of a meta-font. Technical Report STAN-CS-81-886, Stanford University, October 1981.

Knuth:1982:CM

[Knu82] Donald E. Knuth. The Concept of a Meta-Font. Visible Language, 16 (1):3–27, winter 1982. French translation by M. R. Delorme in Communication et Langages no. 55, (1983), 40–53; reprinted in Typographie et Informatique, proceedings of IN-RIA conference held 21–25 January 1985, organized by Jacques André and Patrick Sallio (Rennes, France: INRIA/IRISA – CCETT, 1985), 119–132.

Knuth:1985:LLM

[Knu85] Donald E. Knuth. Lessons learned from Metafont. Visible Language, XIX(1):35–53, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_
E.pdf.

Knuth:1986:LVF

[Knu86a] Donald E. Knuth. The TEX logo in various fonts. TUGboat, 7(2):101– 102, June 1986. ISSN 0896-3207.

Knuth:1986:MP

[Knu86b] Donald E. Knuth. : The Program, volume D of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13438-1. xv + 560 pp. LCCN Z250.8.M46 K578 1986.

Knuth:1986:MB

[Knu86c] Donald E. Knuth. The - book, volume C of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13445-4 (hardcover), 0-201-13444-6 (paperback). xi + 361 pp. LCCN Z250.8.M46 K58 1986.

Knuth:1986:CMT

[Knu86d] Donald E. Knuth. Computer Modern Typefaces, volume E of Computers and Typesetting. Addison-Wes-ley, Reading, MA, USA, 1986. ISBN 0-201-13446-2. xv + 588 pp. LCCN Z250.8.M46 K574 1986.

Knuth:1986:TTM

[Knu86e] Donald Ervin Knuth. A torture test for METAFONT. Dept. of Computer Science, Stanford University, Stanford, CA, USA, 1986. 78 pp.

Knuth:1987:FDH

[Knu87] Donald E. Knuth. Fonts for digital halftones. TUGboat, 8(2):135–160, July 1987. ISSN 0896-3207.

Knuth:1988:PMF

[Knu88a] Donald E. Knuth. A Punk Meta-Font. TUGboat, 9(2):152–168, August 1988. ISSN 0896-3207.

Knuth:1988:PM

[Knu88b] Donald E. Knuth. A punk metafont. TUGboat, 9(2):152–168, August 1988. ISSN 0896-3207.

Knuth:1989:CTE

[Knu89] Donald E. Knuth. Computers and typesetting: errata and changes as of 11 October 1989. Computers and typesetting. TEX Users Group, San Francisco, CA, USA, 1989. various pp.

Knuth:1990:VFMa

[Knu90a] Donald E. Knuth. Virtual fonts: More fun for grand wizards. TEXhax, 90(11 and 12), January 1990.

Knuth:1990:VFMb

[Knu90b] Donald E. Knuth. Virtual Fonts:

More Fun for Grand Wizards. TUGboat, 11(1):13–23, April 1990. ISSN 0896-3207.

Knuth:1991:LF

[Knu91] D. E. Knuth. L’avenir de TEX et de (English: The Future of TEX and ). Cahiers GUTenberg, 8:1–2, mars 1991. ISSN 1140-9304.

Knuth:1992:AH

[Knu92] Donald E. Knuth. Axioms and Hulls. Number 606 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-55611-7 (New York), 3-540-55611-7 (Berlin). ix + 109 pp. LCCN QA640.3 .K68 1992.

Knuth:1993:M

[Knu93] Donald Ervin Knuth. The META-FONTbook, volume C of Computers and typesetting. Addison-Wes-ley, Reading, MA, USA, 1993. ISBN 0-201-13445-4 (hard), 0-201-13444-6 (soft). xi + 361 pp. LCCN Z250.8.M46K58 1986.

Knuth:1996:IMR

[Knu96] Donald E. Knuth. Important message regarding CM fonts. TUGboat, 17(1):29, March 1996. ISSN 0896-3207.

Kohen:1989:SEW

[Koh89] E. Kohen. A simple and efficient way to design middle resolution fonts. In André and Hersch [AH89], pages 22–33. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Kokula:1994:AGS

[Kok94] Michael Kokula. Automatic generation of script font ligatures based on curve smoothness optimization. Electronic Publishing—Origination, Dissemination, and Design, 7(4):217–229, December 1994. CO-DEN EPODEU. ISSN 0894-3982.

Komachi:1991:ISO

[Kom91] Y. Komachi. International standardization overview on document description languages and fonts. Joho Shori, 32(10):1110– 1117, 1991. CODEN JOSHA4. ISSN 0447-8053.

Komachi:1993:ISF

[Kom93] Y. Komachi. International standardization for front information interchange. Joho Shori, 34(7):915– 921, July 1993. CODEN JOSHA4. ISSN 0447-8053.

Kopec:1993:LFM

[Kop93] Gary E. Kopec. Least-squares font metric estimation from images. IEEE Transactions on Image Processing, 2(4):510–519, October 1993. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

Kahan:1987:RPC

[KPB87a] Simon Kahan, Theo Pavlidis, and Henry S. Baird. On the recognition of printed characters of any font and size. IEEE Trans. on Pattern Anal. and Machine Intel., PAMI-9 (2):274–288, March 1987. CODEN ITPIDJ. ISSN 0162-8828.

Khan:1987:RPC

[KPB87b] Simon Khan, Theo Pavlidis, and Henry S. Baird. On the recognition of printed characters of any font and size. IEEE Transactions on Pattern Analysis and Machine Intelligence, PAMI-9(2):274–288, March 1987. CODEN ITPIDJ. ISSN 0162-8828.

Kreps:1990:CMT

[Kre90] David M. Kreps. A Course in Microeconomic Theory. Princeton University Press, Princeton, NJ, USA, 1990. ISBN 0-691-04264-0. xviii + 850 pp. LCCN HB172 .K74 1990. Typeset using PostScript and Computer Modern fonts.

Krtolica:1994:BCA

[Krt94] Radovan V. Krtolica. Box connectivity approach to multifont character recognition. Proceedings of SPIE — The International Society for Optical Engineering, 2181:315– 321, 1994. CODEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic).

Kurz:1993:FFT

[Kur93] Phil Kurz. Fractured font tale: How to give your text a storyboard ending. AV video, 15(9):38–??, September 1993. ISSN 0747-1335.

Kapr:1993:FFG

[KW93] Albert Kapr and Hans Peter Willberg. Fraktur: Form und Geschichte der gebrochenen Schriften. (German) [Fraktur: Form and History of Broken Scripts]. Schmidt, Mainz, Germany, 1993. ISBN 3-87439-260-0. 248 pp. LCCN ????

Lam:1980:EEH

[Lam80] Siu Ho Lam. Efficient encoding of high-density Chinese character fonts by stroke composition. Thesis (m.s.), M.I.T., Department of Electrical Engineering and Computer Science, ????, 1980. 81 pp. Supervised by Francis F. Lee.

Langenfelt:1940:FNE

[Lan40] Gosta Langenfelt. Family names as English font-names and other name problems. Almqvist and Wiksells Boktryckeri, Uppsala, 1940. 107– 136 pp. LCCN CS2505 .L36 1940 off campus. Reprinted from: Studier i modern sprakvetenskap, v. 14.

Lange:1994:DTF

[Lan94] Gerald Lange. Digital type foundry for typophiles: Carter & Cone Type Inc. Serif: The Magazine of Type and Typography, 1:50–56, 58, Fall 1994. ISSN 1080-2070. A font-by-font accounting of the new designs by Matthew Carter as part of his new enterprise.

Larsson:1989:CF

[Lar89] Jan Eric Larsson. A chess font for TEX. TUGboat, 10(3):351, November 1989. ISSN 0896-3207.

Latousek:1995:SDC

[Lat95] Rob Latousek. Software directory for the classics. American Classical League, Teaching Materials and Resource Center, Miami University, Oxford, OH, USA, 1995. ISBN 0-939507-45-5. iv + 120 pp. LCCN ????

Lam:1986:PTM

[LB86] S. W. Lam and H. S. Baird.

Performance testing of mixed-font, variable-size character recognizers. Computing Science Technical Report 126, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 1986.

Lamb:1990:CRT

[LB90] D. A. Lamb and D. T. Barnard.

Compact representation of text files. Technical report, Queen’s University, Kingston, Ontario, March 1990. 1–7 pp.

Lee:1998:BVM

[LC98] Do-Hoon Lee and Hwan-Gue Cho.

The beta-velocity model for simulating handwritten Korean scripts. Lecture Notes in Computer Science, 1375:252–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/1375/13750252.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/1375/
13750252.pdf.

Logar:1994:PCC

[LCO94] Antonette M. Logar, Edward M.

Corwin, and William J. B. Oldham. Performance comparisons of classification techniques for multi-font character recognition. International Journal of Human Computer Studies, 40(3):403–423, March 1994. ISSN 1071-5819.

Lee:1994:CES

[Lee94] T.-L. Lee. CTEX evolving into scalable font technology. In Anonymous [Ano94q], pages 85–92. ISSN 0218-7019.

Lesenko:1995:PDP

[Les95] Sergey Lesenko. Printing TEX documents with partial Type 1 fonts. TUGboat, 16(3):265–268, September 1995. ISSN 0896-3207.

Lester:2015:GBH

[Les15] Valérie Lester. Giambattista Bodoni: his life and his world. David R. Godine, Publisher, Jaffrey, NH, USA, 2015. ISBN 1-56792-528-6. 255 + 16 pp. LCCN Z232.B66 L47 2015.

LeVeque:1981:FDA

[LeV81] William J. LeVeque. Font development at the AMS. TUGboat, 2(2): 39–40, July 1981. ISSN 0896-3207.

Levy:1988:UGF

[Lev88] Silvio Levy. Using Greek fonts with TEX. TUGboat, 9(1):20–24, April 1988. ISSN 0896-3207.

Liao:1991:FGB

[LH91] Chia-Wei Liao and Jun S. Huang. Font generation by beta-spline curve. Computers and Graphics, 15(4):527–534, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Linderholm:1991:AIL

[Lin91] O. Linderholm. Apple introduces low-cost laser quality with style. Byte Magazine, 16(5):48, 50, May 1991. CODEN BYTEDJ. ISSN 0360-5280.

Lin:1995:RAC

[Lin95] Zhen Yong Lin. Relational attribute code representation for printed multi-font Chinese character recognition. Proceedings of the IEEE International Conference on Systems, Man and Cybernetics, 2:1924–1927, 1995. CODEN PICYE3. ISSN 0884-3627. IEEE catalog number 95CB35767.

Livingston:1995:WM

[Liv95] Brian Livingston. Window manager. InfoWorld, 17(22):33–??, May 1995. CODEN INWODU. ISSN 0199-6649.

Lim:1995:OCF

[LK95] Soon-Bum Lim and Myung-Soo Kim. Oriental character font design by a structured composition of stroke elements. Computer Aided Design, 27(3):193–207, March 1995. CODEN CAIDA5. ISSN 0010-4485.

Lin:1994:CMF

[LL94] Y.-J. Lin and T.-L. Lee. Conversion from Metafont files to TrueType. In Anonymous [Ano94e], pages 65–75. ISSN 0218-7019.

Lai:1991:FTH

[LM91] P. Lai and U. Manber. Flying through hypertext. In Proc. of ACM Hypertext’91, Presentation Issues, page 123. ACM Press, New York, NY, USA, 1991.

Level:1995:PTF

[LNN95] Jeff Level, Bruce Newman, and Brenda Newman. Precision Type Font Reference Guide, Version 5.0. Precision Type, Commack, NY, USA, 1995. ISBN 0-9646252-0-2. xxv + 653 pp. LCCN Z250.7 .P74 1995.

Level:2000:PTF

[LNN00] Jeff Level, Bruce Newman, and Brenda Newman. Precision Type Font Reference Guide, Version 5.0. Hartley & Marks Publishers, Point Roberts, WA, USA and Vancouver, BC, Canada, 2000. ISBN 0-88179-182-2. xxv + 653 pp. LCCN Z250.7 .P74 2000. US$39.95, CDN$59.95. Republication of [LNN95].

Lee:1990:FRN

[LO90] Ming-Chih Lee and William J. B. Oldham. Font recognition by a neural network. International Journal of Man-Machine Studies, 33(1):41– 61, July 1990. CODEN IJMMBC. ISSN 0020-7373.

LopesCardozo:1985:SKW

[Lop85] Lida Lopes Cardozo. Stonecuttings from Kindersley’s workshop. Visible Language, XIX(1):55–59, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_
E.pdf.

Lorin:1986:BFM

[Lor86] Lorin Software. Braille font for MousePaint, 1986. 1 computer disk.

Lotus:1993:SW

[Lot93] Lotus Development Corporation. SmartSuite for Windows, 1993. 25 computer disks cards.

Louarn:1991:LFC

[Lou91] Ph. Louarn. Lucida, une fonte complète pour LaTEX et son installation (English: Lucida, a complete font for LaTEX and its installation). Cahiers GUTenberg, 9:32–40, juillet 1991. ISSN 1140-9304.

Lovenheim:2024:SRS

[Lov24] Peter Lovenheim. Seeing Rochester signs through Chuck Bigelow’s eyes. Rochester Beacon, ?? (??):??, July 29, 2024. URL

https://rochesterbeacon.com/
2024/07/29/seeing-rochester-
signs-through-chuck-bigelows-
eyes/.

Leitch:1984:CTC

[LS84] S. Leitch and F. J. Smith.

Computer-aided typography, cubic splines in font design. In PRO-TEXT I. Proc. the First Int. Conf. on Text Processing, pages 173–178. Boole Press, Dublin, Ireland, 1984.

Lubeck:1991:HPL

[LS91] J. H. Lubeck and B. D. Schatzman. High-end printing on a lowend budget. Byte Magazine, 16(9): 315–316, 318–320, September 1991. CODEN BYTEDJ. ISSN 0360-5280.

Lucarella:1985:PFE

[Luc85] Dario Lucarella, editor. Proceedings of the First European Conference on TEX for Scientific Documentation, 16–17 May 1985, Como, Italy. Addison-Wesley, Reading, MA, USA, 1985. ISBN 0-201-13399-7. LCCN Q224.E97

Luhn:1992:W

[Luh92] Robert Luhn. Windows 3.1. PC World, 10(2):174–181, February 1992. CODEN PCWDDV. ISSN 0737-8939.

Lohoff-Werner:1988:TTL

[LW88] Joachim Lohoff-Werner. Treiber für TEX mit Laden von Fonts. Studienarbeit, 1988.

Li:1986:NAR

[LZ86] Bin Li and Shuxiang Zhao. New approach to recognition of both handwritten and multi-font printed Chinese characters. In IEEE:1986:PEI [IEE86], pages 641–643. CODEN PICREG. ISBN 0-8186-0742-4. IEEE Service Cent. Piscataway, NJ, USA.

Morris:1991:RID

[MA91] Robert A. Morris and Jacques André, editors. Raster Imaging and Digital Topography II: Papers from the second RIDT meeting, held in Boston, Oct. 14–16, 1991. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991. MackenzieHarris:1935:CAF

[Mac35] Mackenzie and Harris. Centaur and Arrighi: font sizes and prices. Mackenzie and Harris, San Francisco, 1935. 1 pp. LCCN pf Z250.M12 C4 1935z.

Mackenzie-HarrisCorp:1978:FSL

[Mac78] Mackenzie-Harris Corp. Font, sorts, leads, slugs, metal furniture, rules, borders. Mackenzie-Harris Corp., San Francisco, CA, USA, 1978. 95 pp.

MacKay:1990:ITS

[Mac90] Pierre A. MacKay. The internationalization of TEX with special reference to Arabic. Proceedings of the IEEE International Conference on Systems, Man and Cybernetics, pages 481–484, November 1990. CODEN PICYE3. ISBN 0-87942-597-0. ISSN 0884-3627. IEEE catalog number 90CH2930-6.

MacKay:1991:RPO

[Mac91a] P. A. MacKay. Un regard sur les pixels. obtention de fontes de qualité pour imprimantes à laser à 300 dpi grâce à (English: A look at pixels. Getting fonts of quality for 300-dpi laser printers thanks to - ). Cahiers GUTenberg, 12:21– 36, décembre 1991. ISSN 1140-9304.

MacKay:1991:LAP

[Mac91b] Pierre A. MacKay. Looking at the pixels: Quality control for 300 dpi laser printer fonts, especially s. In Morris and André [MA91], pages 205–215. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Malyshev:1994:ACF

[Mal94] Basil Malyshev. Automatic conversion of fonts to Type1 PostScript. TUGboat, 15(3):200, September 1994. ISSN 0896-3207.

Malyshev:1995:PCF

[Mal95] Basil Malyshev. Problems of the conversion of fonts to PostScript Type 1. TUGboat, 16 (1):60–68, March 1995. ISSN 0896-3207.

Mandel:1993:DAT

[Man93] Ladislas Mandel. Developing an awareness of typographics letterforms. Electronic Publishing—Origination, Dissemination, and Design, 6(1):3–22, March 1993. CO-DEN EPODEU. ISSN 0894-3982.

Marovac:1987:HFE

[Mar87] Nenad Marovac. Handling fonts in electronic publishing systems. Computers and Graphics, 11(3):289– 295, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mayo:1985:SCP

[May85] Randy D. Mayo. System control for a printwheel typewriter. IBM Journal of Research and Development, 29(5):488–493, September 1985. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

May:1986:DFG

[May86] Harry L. May. Design of a font generator for a laser printer. Thesis (m.s.), University of Illinois at Urbana-Champaign, Urbana-Champaign, IL 61801, USA, 1986. vii + 34 pp.

Morris:2002:PSS

[MAYB02] R. A. Morris, K. Aquilante, D. Yager, and C. Bigelow. P-13: Serifs slow RSVP reading at very small sizes, but don’t matter at larger sizes. SID Symposium Digest of Technical Papers, 33(1):244–247, ???? 2002.

McQueen:1993:IPF

[MB93] Clyde D. McQueen III and Raymond G. Beausoleil. Infinifont: a parametric font generation system. Electronic Publishing—Origination, Dissemination, and Design, 6(3):117–132, September 1993. CO-DEN EPODEU. ISSN 0894-3982.

Maddox:1977:FCM

[MBG77] Michael E. Maddox, James T. Burnette, and James C. Gutmann. Font comparisons for 5 × 7 dot matrix characters. Human Factors, 19(1): 89–93, February 1977. CODEN HUFAA6. ISSN 0018-7208.

Meggs:1993:TSG

[MC93] Philip Meggs and Rob Carter. Typographic Specimens: The Great Typefaces. John Wiley and Sons, Inc., Chichester, UK, 1993. ISBN 0-471-28429-7. 416 (est.) pp. LCCN ???? US$44.95.

McClure:1983:TWF

[McC83] Robert M. McClure. Testing the widths of a font. TUGboat, 4(1): 38, April 1983. ISSN 0896-3207.

McCord:1984:FS

[McC84] E. McCord. Font substitution. IBM Technical Disclosure Bulletin, 27 (1B):547–548, June 1984. CODEN IBMTAA. ISSN 0018-8689.

McClure:1987:NFC

[McC87] Rebecca Lamar McClure. NEW-FONT: font creation program for the Texas Instruments model 855 printer: software manual. Thesis (honors), Auburn University, Auburn, AL, USA, 1987. 33 pp.

McCormack:1991:IDT

[McC91] Jscr. McCormack. Iconographic dental typography. A dental character font for computer graphics. British dental journal, 170(11):417– ??, June 1991. CODEN BDJOHJ. ISSN 0007-0610.

McCallum:2007:FIP

[McC07] John McCallum. Fonts are intellectual property too! IT-NOW, 49(4):6–7, July 2007. ISSN 1746-5702. URL http://

                                        itnow.oxfordjournals.org/cgi/
                                        content/abstract/49/4/6?etoc.

McCarthy:2016:MTD

[McC16] Steven McCarthy. Missing a type of design: A review of Make It New: The History of Silicon Valley Design by Barry M. Katz. The Design Journal, 19(2):535–539, ???? 2016. ISSN ????

McCarthy:2019:DSV

[McC19] Steven McCarthy. Design at Stanford: a Visual History of Thinking and Doing. Blurb, San Francisco, CA, USA, 2019. ISBN ???? ???? pp. LCCN ????

McCarthy:2020:DTS

[McC20] Steven McCarthy. Digital typography at Stanford. She Ji: The Journal of Design, Economics, and Innovation, 6(4):546–560, Winter 2020. ISSN 2405-8726.

Martin:1992:SAA

[MCL92] James Martin, Kathleen Kavanagh Chapman, and Joe Leben. Systems application architecture: common communications support: distributed applications. Prentice Hall, Englewood Cliffs, NJ, USA, 1992. ISBN 0-13-785908-2. xxiv + 536 pp. LCCN QA76.9.A73 M373 1992.

Maeng:1991:TPG

[MCP91] Seung-Rul Maeng, Zeung-Seong Cho, and Dong-Soon Park. Text processing with a glyph mapping model. Korea Information Science Society Review, 9(6):74–78, 1991. CODEN CHKWEN. ISSN 1015-9908.

McVarish:2010:CGU

[McV10] Emily McVarish. “The crystal goblet”: The underpinnings of typographic convention. Design and Culture, 2(2):285–307, ???? 2010.

McWorter:1987:NFM

[McW87] William A. McWorter. New fonts for mathematics? TUGboat, 8(1): 76–777, April 1987. ISSN 0896-3207.

Mei:1980:LLC

[Mei80] Tung Yun Mei. LCCD, a language for Chinese character design. Report STAN-CS-80-824, Stanford University, Department of Computer Science, 1980. ?? pp. See [Mei81a, Mei81b].

Mei:1981:LAL

[Mei81a] Tung Yun Mei. LCCD, a language for Chinese character design. Software—Practice and Experience, 11(12):1273–1292, December 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See also [Mei80].

Mei:1981:LLC

[Mei81b] Tung Yun Mei. LCCD, a language for Chinese character design. Software—Practice and Experience, 11(12):1273–1292, December 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). See also [Mei80].

Meyers:1994:NFT

[Mey94] Nathan Meyers. New font technology for X11R6: Changes for R6 increase font rendering capabilities, improve performance, and enable licensing. The X Resource, 9(1):211– 227, January 1994. CODEN XRE-SEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Mittelbach:2023:LCPb

[MFB^{+}23] Frank Mittelbach, Ulrike Fischer, Javier Bezos, Johannes Braams, and Joseph Wright. The La-TeX Companion: Part II. Tools and techniques for computer typesetting. Addison-Wesley, Reading, MA, USA, third edition, 2023. ISBN 0-201-36300-3 (Part II) (hardcover), 0-13-816657-9 (Part II) (PDF), 0-13-816652-8 (Part II) (ePub). xxxviii + 970 pp. LCCN Z253.4.L38 M58 2023.

Mittelbach:2023:LCPa

[MFW23] Frank Mittelbach, Ulrike Fischer, and Joseph Wright. The La-TeX Companion: Part I. Tools and techniques for computer typesetting. Addison-Wesley, Reading, MA, USA, third edition, 2023. ISBN 0-13-465894-9 (Part I) (hardcover), 978-013-465-9-5-9-6 (Part I) (PDF), 978-013-465-9-6-0-2 (Part I)

(e-pub), 0-13-816648-0-X (Part I + II). xxviii + 947 pp. LCCN Z253.4.L38 M58 2023.

Malyshev:1995:PFE

[MG95] Basil Malyshev and Michel Goossens.

Partial font embedding utilities for PostScript Type-1 fonts. TUGboat, 16(1):69–77, March 1995. ISSN 0896-3207.

Morris:1998:LCP

[MHC98] R. A. Morris, R. D. Hersch, and A. Coimbra. Legibility of condensed perceptually-tuned grayscale fonts. Lecture Notes in Computer Science, 1375:281–293, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/1375/13750281.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/1375/
13750281.pdf.

MicroPro:1987:WPa

[Mic87a] MicroPro International Corporation. WordStar 2000 plus, 1987. 12 computer cartridges command booklet.

MicroPro:1987:WPb

[Mic87b] MicroPro International Corporation. WordStar 2000 plus, 1987. 19 computer disks and learning manual + 1 companion programs manual + 1 command booklet + 1 customer service guide + 3 templates.

MicroPro:1987:WPc

[Mic87c] MicroPro International Corporation. WordStar 2000 plus, 1987. 20 computer disks command booklet + 1 instruction leaflet.

Microsoft:1992:MTFb

[Mic92a] Microsoft Corporation. Microsoft TrueType font pack 2 for Windows, 1992. 2 computer disks.

Microsoft:1992:MTFa

[Mic92b] Microsoft Corporation. Microsoft TrueType font pack 2 user’s guide. Microsoft Corp., Redmond, WA, USA, 1992. xii + 106 pp.

Middlebrook:1995:BSC

[Mid95] Mark Middlebrook. bugs: Save command that performs as a Saveas, a jumping point problem, overly sensitive dimensions, documentation fallacies, a Sans Serif font problem, and commands that cause graphics anomalies. CADalyst: the newsletter of the Auto-CAD Users’ Group, ????(????):13– ??, May 1995. ISSN 0820-5450.

Milligan:1981:MAF

[Mil81] Patrick Milligan. TEX macros for automatic font code allocation. TUGboat, 2(2):44–45, July 1981. ISSN 0896-3207.

MECC:1983:AFE

[Min83] Minnesota Educational Computing Consortium. Atari font editor. Minnesota Educational Computing Consortium, St. Paul, MN, USA, 1983. 43 pp.

Minasi:1991:LFW

[Min91] Mark J. Minasi. LaserJets, fonts, and Windows: Getting fancy fonts from Windows is tricky but doable.

Byte Magazine, 16(7):271–??, July 1991. CODEN BYTEDJ. ISSN 0360-5280.

Mittelbach:1992:CF

[Mit92] Frank Mittelbach. Comments on “Filenames for Fonts” (TUGboat 11#4). TUGboat, 13(1):51–53, April 1992. ISSN 0896-3207.

Millar:1977:AFP

[MN77] J. B. Millar and M. C. Newey. Arbitrary font printing system. Papers and Presentations — Proceedings of the Digital Equipment Computer Users Society, 3(5):1641–1647, ???? 1977. CODEN DEPPAD.

Mudur:1999:CGIb

[MNSJ99] S. P. Mudur, Niranjan Nayak, Shrinath Shanbhag, and R. K. Joshi. Computer graphics in India — an architecture for the shaping of Indic texts. Computers and Graphics, 23(1):7–24, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://www.

elsevier.com/cas/tree/store/
cag/sub/1999/23/1/647.pdf.

McNutt:1992:XAF

[MO92] Dinah McNutt and Miles O’Neal.

The X administrator: Font formats and utilities. The X Resource, 2(1): 14–34, April 1992. CODEN XRE-SEA. ISBN 0-937175-97-8. ISSN 1058-5591.

Modegi:1993:HFG

[Mod93] T. Modegi. Halftone font generator for HDTV display. In Dubois and Chiariglione [DC93], pages 465– 471. ISBN 0-444-81551-1. LCCN TK6679.I57 1992.

Mohanty:1998:FPT

[Moh98] S. K. Mohanty. The formulation of parameters for type design of Indian scripts based on calligraphic studies. Lecture Notes in Computer Science, 1375:157–166, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/1375/13750157.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/1375/
13750157.pdf.

MolecularDesign:1989:Ca

[Mol89a] Molecular Design Ltd. ChemText, 1989. 13 computer disks + 1 quick guide (1 v. ; 30 cm.) + 1 quickreference card + 1 installation guide (25 p. ; 28 cm.) + 1 addendum (21 p. ; 28 cm.) + 1 hardware requirements data sheet.

MolecularDesign:1989:Cb

[Mol89b] Molecular Design Ltd. ChemText, 1989. 14 computer disks guide + 1 quick-reference card + 1 Chem-Text 1.2/1.3 Addendum + 1 Chem-Text 1.3 installation + 1 hardware requirements sheet.

Morreale:1993:AMM

[Mor93] J. Morreale, editor. Annual meeting — 1993 May: Seattle; WA, volume 24 of SID INTERNA-TIONAL SYMPOSIUM DIGEST OF TECHNICAL PAPERS. SID, ????, 1993. ISSN 0097-966X (print), 2154-6738 (electronic), 2154-6746.

Moye:1995:FTD

[Moy95] Stephen Moye. Fontographer: Type by Design. MIS:Press, New York, NY, USA, 1995. ISBN 1-55828-447-8. xi + 275 pp. LCCN Z250.8.F65 M69 1995. US$30.00 (est.).

Murray:1992:WP

[MP92] William H. Murray and Chris H.

Pappas. Windows 3.1 programming. Osborne McGraw-Hill, Berkeley, 1992. ISBN 0-07-881855-9. xviii + 805 pp. LCCN ????

Ma:1994:AGA

[MP94] Xiaohu Ma and Zhigeng Pan. Automatic generation algorithm of highquality outline font using Bézier curve. Zidonghua Xuebao/Acta Automatica Sinica, 20(1):121–125, January 1994. CODEN THHPAY. ISSN 0254-4156.

Ma:1995:AGC

[MPZ95] Xiaohu Ma, Zhigeng Pan, and Fuyan Zhang. Automatic generation of Chinese outline font based on stroke extraction. Journal of computer science and technology, 10 (1):42–52, January 1995. CODEN JCTEEM. ISSN 1000-9000.

McThompson:1983:ECF

[MRS83] O. W. McThompson, R. F. Romon, and L. A. Sendelbach. Ergonomic character font for CRT displays. IBM Technical Disclosure Bulletin, 26(4):2120–??, September 1983. CODEN IBMTAA. ISSN 0018-

Morris:1989:CFS

[MRT89] R. J. T. Morris, L. D. Rubin, and H. Tirri. A comparison of feedforward and self-organizing approaches to the font orientation problems. In International Joint Conference on Neural Networks — volume II (Jun 18–22 1989: Washington, DC, USA), pages 291– 297. IEEE Computer Society Press, IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. IEEE catalog number 89CH2765-6.

Mursi:1984:CFS

[MS84] Mona M. Mursi and Mohamed Saba. Character font simulation system. Modeling and Simulation on Microcomputers, pages 131–135, 1984. CODEN MSMIE8.

Mittelbach:1989:NFS

[MS89] Frank Mittelbach and Rainer Schöpf. A new font selection scheme for TEX macro packages—the basic macros. TUGboat, 10(2):222–238, July 1989. ISSN 0896-3207.

Mittelbach:1990:RNF

[MS90a] Frank Mittelbach and Rainer Schöpf. Reprint: The new font family selection—User interface to standard LaTEX. TUGboat, 11(2): 297–305, June 1990. ISSN 0896-3207.

Mittelbach:1990:NFF

[MS90b] Frank Mittelbach and Rainer Schöpf. The new font family selection—User interface to standard LaTEX. TUGboat, 11(1):91–97, April 1990. ISSN 0896-3207.

Moon:1990:CFT

[MS90c] Y. S. Moon and T. Y. Shin. Chinese fonts and their digitization. In Furuta [Fur90], pages 235–248. ISBN 0-521-40246-8. LCCN Z 286 E43 I58 1990.

Mullen:1991:EFD

[MS91] Donald C. Mullen and Simon L.

Smith. EXP, the font development kit: version 2.0. Brooks/Cole Pub. Co., Pacific Grove, CA, USA, 1991. ISBN 0-534-13326-6 (singleuser version), 0-534-13329-0 (sitelicense). vii + 61 pp. LCCN ????

Mumford:1989:CTT

[Mum89] A. M. Mumford. The CGM today and tomorrow. Computer Graphics Forum, 8(2):125–128, June 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Murayama:1990:DTF

[Mur90] N. Murayama. DTP today and future. Information Processing Society of Japan, 31(11):1483–1494, 1990. CODEN JOSHA4. ISSN 0447-8053.

Miller:2009:AAA

[MVM09a] Frederic P. Miller, Agnes F. Vandome, and John McBrewster, editors. Arabic Alphabet: Arabic alphabet, Abjad numerals, ‘i’rab, Arabic diacritics, Arabic numerals, History of the Arabic alphabet, Arabic Unicode, Islamic calligraphy, … Romanization of Arabic, Arabic chat alphabet. Alphascript Publishing, 17 Rue Meldrum, Beau Bassin, 1713-01 Mauritius, 2009. ISBN 613-0-04650-2. 100 (est.) pp. LCCN ???? US$50.

Miller:2009:RNR

[MVM09b] Frederic P. Miller, Agnes F. Vandome, and John McBrewster, editors. Roman Numerals: Roman numeral analysis, Roman arithmetic, Roman abacus, Kharosthi, Unicode numerals, Etruscan numerals, Positional notation, Arabic numerals, Districts of Turku. Alphascript Publishing, 17 Rue Meldrum, Beau Bassin, 1713-01 Mauritius, 2009. ISBN 613-0-06480-2. 88 (est.) pp. LCCN ???? US$50.

Mylonas:1992:MGA

[MW92] C. Mylonas and R. Whitney. Modern Greek with adjunct fonts. TUGboat, 13(1):39–50, April 1992. ISSN 0896-3207.

Nass:1994:LCE

[Nas94] Richard Nass. Laser-printer controller eases memory demands. Electronic Design, 42(22):139–140, October 1994. CODEN ELODAW. ISSN 0013-4872.

Sejnowski:1989:NC

[NC89] Neural Computation. MIT Press, 1989. ISSN 0899-7667. Technical journal published four times per year, using PostScript and Computer Modern fonts.

Nebel:1993:XL

[Neb93] Stephen Nebel. Xbase and the LaserJet. DBMS, 6(1):55–??, January 1993. CODEN DBMSEO. ISSN 1041-5173.

Nelson:1985:JCG

[Nel85a] Greg Nelson. Juno, a constraintbased graphics system. Computer Graphics, 19(3):235–243, July 1985. CODEN CPGPBZ. ISBN 0-89791-166-0. ISSN 0097-8930.

Nelson:1985:MMM

[Nel85b] Stan Nelson. Mould making, matrix fitting, and hand casting. Visible Language, XIX(1):107–120, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_
E.pdf.

Naiman:1987:RCF

[NF87] Avi Naiman and Alain Fournier.

Rectangular convolution for fast filtering of characters. Computer Graphics, 21(4):233–242, July 1987. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

URL http://www.acm.org:80/
pubs/citations/proceedings/graph/
37401/p233-naiman/.

Nicole:1991:EPS

[Nic91] O. Nicole — trad. The Economist polit ses polices (English: The Economist polishes its fonts). Cahiers GUTenberg, 9:41–48, juillet 1991. ISSN 1140-9304.

Nizamani:1990:DMD

[Niz90] A. S. Nizamani. On dot matrix displays for Arabic text. IEEE Transactions on Consumer Electronics, 36(2):76–83, May 1990. CODEN ITCEDA. ISSN 0098-3063.

Nagashima:1989:DOF

[NKMO89] Ichiro Nagashima, Naoyuki Kai, Tsutomu Minagawa, and Masahide Ohhashi. Design of an outline font rasterizing LSI. In Symposium on VLSI Circuits Digest of Technical Papers (May 25–27 1989: Kyoto, Japan), pages 117–118. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. IEEE catalog number 89TH0262-6.

Nanard:1989:AML

[NN89] M. Nanard and J. Nanard. An anti-aliasing method for low resolution fonts based on font structure. In André and Hersch [AH89], pages 111–122. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Nanard:1989:DAF

[NNGP89] M. Nanard, J. Nanard, M. Gandara, and N. Porte. Declarative approach for font design by incremental learning. In André and Hersch [AH89], pages 71–82. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Namiki:1994:SAF

[NNMS94] M. Namiki, K. Nakajima, A. Morma, and S. Sasagawa. Software architecture and font management of a print server for large character sets. In Anonymous [Ano94r], pages 714– 717. ISBN 0-89208-177-5. LCCN ????

Namane:1990:CSC

[NSA90] A. Namane and M. A. Sid-Ahmed. Character scaling by contour method. IEEE Transactions on Pattern Analysis and Machine Intelligence, 12(6):600–606, June 1990. CODEN ITPIDJ. ISSN 0162-8828.

Ngan:1983:MPS

[NW83] David K. Ngan and Scott W. Wood.

3800 model 3 printing subsystem compatibility mode font selection guide. Technical bulletin. IBM Corporation, Gaithersburg, MD, USA, 1983. vi + 44 pp.

Olsak:1995:PAF

[Ols95] Petr Olsak. The program a2ac— Font handling on the PostScript level. TUGboat, 16(1):54–59, March 1995. ISSN 0896-3207.

Ou:1989:FGA

[OO89] C. Ou and Y. Ohno. Font generation algorithms for Kanji characters. In André and Hersch [AH89], pages 123–133. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Orr:1994:HBC

[Orr94] Joel Orr. Hoops 4.0: Beyond 3-D:

The common API to most 2-D and 3-D libraries and hardware now offers a real-time mode and a powerful font engine. Byte Magazine, 19 (2):125–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280.

Ostromoukhov:1998:MTC

[Ost98] Victor Ostromoukhov. Mathematical tools for computergenerated ornamental patterns. Lecture Notes in Computer Science, 1375:193–223, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/

                                        bibs/1375/13750193.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/series/0558/papers/1375/
                                        13750193.pdf.

Padovano:1992:GBF

[Pad92] Michael Padovano. Gain better font control with latest X Windows version. Systems integration, XXV(2): 23–24, February 1992. CODEN SYINE9. ISSN 1044-4262.

Pantagraph:1987:FB

[Pan87] Pantagraph Printing and Stationery Company. Font book. Pantagraph Printing and Stationery Company, Bloomington, IL, USA, 1987. various pp.

Pankow:2016:MCE

[Pan16] David Pankow, editor. Manuale calligraphicum: examples of calligraphy by students of Hermann Zapf at RIT, in the manner of Pen and graver, as a tribute to their teacher. The Cary Graphic Arts Collection, Rochester, NY, USA, 2016. xi + + 58 pp. LCCN ???? Foreword by Steven Galbraith.

Pappas:1989:IP

[Pap89a] T. L. Pappas. TEXnology on the IBM PC. Computer, 22(8):111–120, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pappas:1989:PRT

[Pap89b] T. L. Pappas. Product reviews — TEXnology on the IBM PC. Computer, 22(8):111–120, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parker:1988:OHD

[Par88] Mike Parker. Printer fonts. In Robert C. Durbeck and Sol Sherr, editors, Output Hardcopy Devices, chapter 19, pages 497–516. Academic Press, New York, NY, USA, 1988.

Parker:1989:F

[Par89] Mike Parker. Fonts and Post-Script. TUGboat, 10(3):346–350, November 1989. ISSN 0896-3207.

Parker:1990:SLB

[Par90] J. R. Parker. Scaling low-resolution bitmap fonts using linear feature extraction. Visual Computer, 6(2):74– 78, March 1990. CODEN VICOE5. ISSN 0178-2789.

Parida:1993:CTG

[Par93a] Laxmi Parida. Computational technique for general shape deformations for use in font design. Computers and Graphics, 17(4):349– 356, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Parida:1993:VIC

[Par93b] Laxmi Parida. Vinyas: An interactive calligraphic type design system. IFIP Transactions B: Computer Applications in Technology, B-9: 287–299, 1993. CODEN ITBTEH. ISBN 0-444-81564-3. ISSN 0926-5481.

Parker:1993:VR

[Par93c] Tim Parker. Views and reviews. I.T. magazine, 25(9):34– ??, September 1993. CODEN IT-MGEX. ISSN 0008-3364.

PattonPattonSoftware:1992:FC

[Pat92] Patton and Patton Software Corporation. Flow charting 3, 1992. 6 computer disks 1 network supervisor’s guide.

Patz:1994:NDF

[Pat94] Debby Patz. New digital fontdelivery options. Folio, 23(21):64– ??, December 1994. ISSN 0046-4333.

Probets:2003:SOF

[PB03] S. G. Probets and D. F. Brailsford. Substituting outline fonts for bitmap fonts in archived PDF files. Software—Practice and Experience, 33(9):885–899, July 25, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

                                        tronic). URL http://www.cs.
                                        nott.ac.uk/~dfb/Publications/
                                        Download/2003/Probets03.pdf.

Poynor:1991:TNN

[PBC91] Rick Poynor and Edward Booth-Clibborn, editors. Typography now—-the next wave. Internos Books, London, England, 1991. ISBN 0-904866-90-4. 223 pp. LCCN Z116 .T97 1991.

Pal:1994:OBI

[PC94] U. Pal and B. B. Chaudhuri. OCR in Bangla: An Indo-Bangladeshi language. Proceedings — International Conference on Pattern Recognition, 2:269–273, 1994. CO-DEN PICREG. ISSN 1051-4651. IEEE catalog number 94CH3440-5.

Pearce:1992:XAM

[Pea92] Eric Pearce. The X administrator:

Managing font servers. The X Resource, 3(1):35–46, June 1992. CO-DEN XRESEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Perry:1988:PAC

[Per88] Tekla S. Perry. ‘PostScript’ prints anything: a case history. IEEE Spectrum, 25(5):42–46, May 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Peterson:1986:AAT

[Pet86] John Thomas Peterson. Automatic assembly of the Texas Instruments font module for the TI 855 printer. Thesis (m.s. in engin.), University of Texas at Austin, Austin, TX, USA, 1986. x + 162 pp.

Petzold:1993:E

[Pet93] Charles Petzold. Environments. PC Magazine, 12(3):361–??, February 1993. CODEN PCMGEP. ISSN 0888-8507.

Phelps:1993:DIB

[Phe93] Thomas A. Phelps. Domain-specific information browsers for man page, file, and font: research project. Master of science, plan ii., Dept. of Electrical Engineering and Computer Science, University of California, Berkeley, Berkeley, CA, USA, 1993. 40 pp.

Phelps:1994:DIB

[Phe94] Thomas A. Phelps. Domain-specific information browsers for man page, file, and font. Report UCB/CSD 94/802 4004667978, Computer Science Division (EECS), University of California, Berkeley, Berkeley, CA, USA, 1994. 40 pp. Cover title. “March 1994.” Originally presented as the author’s thesis (Master)– University of California, Berkeley, 1994.

Philco:1994:MP

[Phi94] Philco Corporation. [Miscellaneous pamphlets]. The Corporation, Willow Grove, PA, USA, 1994. 26 volumes in 1.

Pietrowski:1991:ITS

[Pie91] Ken Pietrowski, editor. IS&T’s Seventh International Congress on Advances in Non-impact Printing Technologies, October 6–11, Portland, OR, USA, volume 2. Imaging Science & Technology, ????, 1991.

Park:1993:SEA

[PM93] Seung Woon Park and Seung Ryoul Maeng. Structure extraction and automatic hinting of Chinese outline characters. Electronic Publishing—Origination, Dissemination, and Design, 6(2):67–91, June 1993. CODEN EPODEU. ISSN 0894-3982.

Pan:1997:CFC

[PMZS97] Zhigeng Pan, Xiaohu Ma, Mingmin Zhang, and Jiaoying Shi. Chinese font composition method based on algebraic system of geometric shapes. Computers and Graphics, 21(3):321–328, May– June 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://

www.elsevier.com/cgi-bin/cas/

tree/store/cag/cas_sub/browse/
browse.cgi?year=1997&volume=
21&issue=3&aid=9700009.

Pohlen:2015:LBU

[Poh15] Joep Pohlen. Letterfontäne (über Buchstaben); [das ultimative Handbuch zur Typografie]. Taschen, Köln, Germany, 2015. ISBN 3-8365-5450-X. 640 pp. LCCN ????

Ponting:1989:TCB

[Pon89] Bob Ponting. Three companies break Adobe encryption scheme. InfoWorld, 11(9):8, February 2, 1989. CODEN INWODU. ISSN 0199-6649. URL https:

//books.google.com/books?id=
IToEAAAAMBAJ&pg=PT7.

Adobe:1989:ABS

[Pos89a] PostScript Developer Tools & Strategies Group, Adobe Systems Inc., Mountain View, CA, USA. Adobe binary screen font files (ABF) specification — Version 2.0, January 1989.

Adobe:1989:AFM

[Pos89b] PostScript Developer Tools & Strategies Group, Adobe Systems Inc., Mountain View, CA, USA. Adobe font metric files specification — Version 2.0, January 1989.

Adobe:1989:SFD

[Pos89c] PostScript Developer Tools & Strategies Group, Adobe Systems Inc., Mountain View, CA, USA. Sonata font design specification — Technical Note 045, May 1989. pr: discontinued by Adobe; may be still available to members of Adobe Developers Association.

Adobe:1989:SDP

[Pos89d] PostScript Developer Tools & Strategies Group, Adobe Systems Inc., Mountain View, CA, USA. Supporting downloadable PostScript fonts — Technical Note 040, January 1989. pr: discontinued by Adobe; may be still available to members of Adobe Developers Association.

Adobe:1990:AFM

[Pos90] PostScript Developer Tools & Strategies Group, Adobe Systems Inc., Mountain View, CA, USA. Adobe font metric files specification — Version 3.0, March 1990.

Adobe:1991:FSO

[Pos91a] PostScript Developer Support Group, Adobe Systems Inc., Mountain View, CA, USA. Font switching optimizations — Technical Note 048, March 1991. pr: discontinued by Adobe; may be still available to members of Adobe Developers Association.

Adobe:1991:SHO

[Pos91b] PostScript Developer Support Group, Adobe Systems Inc., Mountain View, CA, USA. The stemsnap hint operator for type 1 font programs — Technical Note 049, February 1991. pr: discontinued by Adobe; may be still available to members of Adobe Developers Association.

Adobe:1991:SFP

[Pos91c] PostScript Developer Support Group, Adobe Systems Inc., Mountain View, CA, USA. Supporting fonts in the PostScript environment — Technical Note 075, May 1991. pr: discontinued by Adobe; may be still available to members of Adobe Developers Association.

Adobe:1992:AFM

[Pos92a] PostScript Developer Support Group, Adobe Systems Inc., Mountain View, CA, USA. Adobe font metric file format specification — Version 4.0, February 1992.

Adobe:1992:ATF

[Pos92b] PostScript Developer Support Group, Adobe Systems Inc., Mountain View, CA, USA. Adobe type 1 font format multi master extensions, February 1992.

Pountain:1989:XWS

[Pou89] D. Pountain. The X Window System. Byte Magazine, 14(1):353–355, 336–360, January 1989. CODEN BYTEDJ. ISSN 0360-5280.

Pountain:1995:BLC

[Pou95] D. Pountain. The British Library’s catalog is on-line. Byte Magazine, 20(5):62–64, 66, 68, 70, May 1995. CODEN BYTEDJ. ISSN 0360-5280.

Poulin:2012:TD

[Pou12] Richard Poulin. Type designers. In Allan Haley, Richard Poulin, Jason Tselentis, Tony Seddon, Gery Leonidas, Ina Saltz, Kathryn Henderson, and Tyler Alterman, editors, Typography, Referenced: a Comprehensive Visual Guide to the Language, History, and Practice of Typography, pages 68–121. Rockport Publishers, Beverly, MA, USA, 2012. ISBN 1-59253-702-2. LCCN Z250 .T985 2012.

Powell:1990:TDP

[Pow90] Patrick Powell. Typesetting with DITROFF PostScript fonts and characters using Typeset — Tpsnew. Technical report TR 90-3, University of Minnesota, Institute of Technology, Computer Science Dept., Minneapolis, MN, USA, 1990. 9 pp.

Prabha:1987:AFT

[PRB87] Chandra Prabha, Duane Rice, and John Bunge. Access to the full text of nonfiction books: Design considerations from a study of public library users. Proceedings of the ASIS annual meeting, 24:196–200, 1987. CODEN PAISDQ. ISBN 0-938734-19-9. ISSN 0044-7870.

Tri-ArtsPress:1957:TTF

[Pre57] Tri-Arts Press. Type: type faces at Tri-Arts Press, Inc. Tri-Arts Press, New York, NY, USA, 1957. 51 + [78] pp.

Panoutsopoulos:1984:MEB

[PS84] A. Panoutsopoulos and C. Y. Suen. Method of encoding binary patterns and font generation. In Proceedings of the 1984 IEEE International Conference on Systems, Man and Cybernetics, Halifax, NS, Can, pages 131–137. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. IEEE Service Cent. Piscataway, NJ, USA.

Pfeiffer:1990:LFB

[PWH90] Katherine Shelly Pfeiffer and Toni Will-Harris. The LaserJet font book. Peachpit Press, Berkeley, Calif., 1990. ISBN 0-938151-06-1. LCCN TK7887.7 .P44 1990. US$24.95.

Pong:1994:HFF

[PZL94] Man-Chi Pong, Youngguang Zhang, and Pak-Keung Lai. The HBF font format: Optimizing fixed-pitch font support. The X Resource, 10(1):113–123, April 1994. CO-DEN XRESEA. ISBN 1-56592-067-8. ISSN 1058-5591.

Zbikowski:1993:HGN

[Raf93] RafałŻbikowski. Hacker’s Guide to AMSFonts and NFSS in the Context of LaTEX. TUGboat, 14(1):62– 69, April 1993. ISSN 0896-3207.

Rahtz:1992:IUV

[Rah92] Sebastian Rahtz. Invisibility using virtual fonts. TUGboat, 13(1):34– 36, April 1992. ISSN 0896-3207.

Rahtz:1993:IEL

[Rah93] Sebastian Rahtz. Implementing the extended TEX layout using PS fonts. TUGboat, 14(2):107–117, July 1993. ISSN 0896-3207.

Radhakrishnan:1983:MIO

[RAK83] T. Radhakrishnan, J. W. Atwood, and S. G. Krishnamoorthy. A multilingual Input/Output device for Indian scripts. International Journal of Man-Machine Studies, 19: 137–146, August 1983. CODEN IJMMBC. ISSN 0020-7373.

Ramshaw:1987:BAC

[Ram87] Lyle Ramshaw. Blossoming: a connect-the-dots approach to splines. Technical Report 19, Digital Equipment Corporation, Systems Research Centre, June 21, 1987. 172 pages. pp.

Rello:2016:EFT

[RBY16] Luz Rello and Ricardo Baeza-Yates. The effect of font type on screen readability by people with dyslexia. ACM Transactions on Accessible Computing, 8(4):15:1–15:??, May 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Reines:1966:VFC

[Rei66] J. Reines. Variable font character recognition machine. In Pattern Recognition, Proceedings of IEEE Workshop, Las Croabas and Dorado, Oct 24–26, 1968. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1966.

Reis:1992:FBF

[Rei92] Reis. The font business: fading fast as we know it. Advanced Imaging, 7(6):34–42, June 1992. CO-DEN ADIMEZ. ISSN 1042-0711.

Reichenthal:1993:GKT

[Rei93a] Steven Reichenthal. Getting to know TrueType. Dr. Dobb’s Journal of Software Tools, 18(10):56–??, Fall 1993. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Reinhart:1993:TMM

[Rei93b] A. Kevin Reinhart. TheWorld-Transliterator: Macintosh Middle Eastern-language transliteration font. Middle East Studies Association bulletin, 27(2):294–??, December 1993. CODEN MESBEL. ISSN 0026-3184.

Reynolds:1993:BE

[Rey93] Rick Reynolds. Bleeding edge. Publish!: the how-to magazine of desktop publishing, 8(3):19–??, March 1993. CODEN PBLIEJ. ISSN 0897-6007.

Rechsteiner:1993:FHP

[RG93] Hubert Rechsteiner and Fran cois Grize. The font handler package. The X Resource, 6(1):183–216, March 1993. CODEN XRESEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Rosenbaum:1975:MOP

[RH75] W. S. Rosenbaum and J. J. Hilliard.

Multifont OCR postprocessing system. IBM Journal of Research and Development, 19(4):398–421, July 1975. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Riehl:1989:CTS

[Rie89] Arthur M. Riehl, editor. Computing Trends in the 1990’s: Seventeenth annual ACM Computer Science Conference, February 21–23, 1989, Commonwealth Convention Center, Louisville, KY. ACM Press, New York, NY, USA, 1989. ISBN 0-89791-299-3. LCCN QA75.5 .A1371 1989. US$36.00.

Riggs:2014:AOS

[Rig14] Tamye Riggs. The Adobe Originals silver anniversary story: Expanding the originals. Typekit (blog), June 30, 2014. URL https:/

                                        /blog.typekit.com/2014/06/30/
                                        the-adobe-originals-silver-anniversary-
                                        story-expanding-the-originals/
                                        ..

Ristroph:1988:MFG

[Ris88] John H. Ristroph. Multi-system font generation. Computers & industrial engineering, 15:467–474, 1988. CODEN CINDDL. ISSN 0360-8352.

Ryan:1987:MBF

[RM87] Tim Ryan and Douglas G. Miles. The Macintosh book of fonts: a desktop publishing handbook for font lovers. SourceNet, Santa Barbara, CA, USA, 1987. ISBN 0-915051-02-

8. various pp. LCCN ????

Riis-Nielson:1983:CSA

[RN83] Hanne Riis-Nielson. Computation sequences: a way to characterize subclasses of attribute grammars. Acta Informatica, 19:255–268, 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See also: report DAIMI PB-138, Computer Science Department, Aarhus University (1981).

Ramasubramanian:1994:ASN

[RNR94] R. Ramasubramanian, R. W. D. Nickalls, and M. A. Reed. ASCII.sty: A new style-option and encoded font with IBM graphics control characters for use with TEX and LaTEX. TUGboat, 15(2):98–103, June 1994. ISSN 0896-3207.

Rokicki:1985:PPF

[Rok85] Tomas Rokicki. Packed (PK) font file format. TUGboat, 6(3):115–120, November 1985. ISSN 0896-3207.

Romano:2016:RHZ

[Rom16] Frank Romano. Remembering Hermann Zapf (Nov. 8, 1918–June 4, 2015). Museum of Printing Web site, 2016. URL https:

//museumofprinting.org/blog/
remembering-hermann-zapf/.

Rosenberg:1989:CDS

[Ros89] S. Rosenberg. Corel Draw shows great promise (graphics program). Byte Magazine, 14(6):213–216, June 1989. CODEN BYTEDJ. ISSN 0360-5280.

Rosenberg:1991:LCM

[Ros91] Charles Rosenberg. A low complexity method of compressing Kanji font bitmaps. In Morris and André [MA91], pages 137–146. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Rogowitz:1993:CTC

[RRW93] B. E. Rogowitz, D. A. Rabenhorst, and T. J. Watson. CRAFT: a tool for customizing color and font selections guided by perceptual rules. In Anonymous [Ano93a], pages 140– 143. ISBN 0-89208-174-0. LCCN ????

Ramsay:1981:CVC

[RS81] Mike Ramsay and Theodore A.

Shaffer. Custom video control delivers advanced font graphics. Electronics, 54(25):151–154, December 1981. CODEN ELECAD. ISSN 0013-5070.

Raine:1989:AOF

[RSSW89] Neil Raine, David Seal, William Stoye, and Roger Wilson. The Acorn Outline Font Manager. In USENIX Association [USE89], pages 25–36. ISBN ???? LCCN ????

Ruder:1988:TG

[Rud88] Emil Ruder. Typographie; Ein Gestaltungsbuch. Niggli/Hatje, Heiden, Stuttgart, Germany, fifth revised edition, 1988. Contains complete English and French translation.

Ruggles:1983:LDS

[Rug83] Lynn Ruggles. Letterform design systems. Report STAN CS-TR-83-971, Department of Computer Science, Stanford University, Stanford, CA, USA, 1983. iv + 23 pp.

URL http://i.stanford.edu/
pub/cstr/reports/cs/tr/83/971/
CS-TR-83-971.pdf.

Rullman:1987:IFD

[Rul87] Bradley R. Rullman. An interactive font design tool that produces device independent font descriptions. Thesis (m.s.), University of Washington, Seattle, WA, USA, 1987. vi + [124] pp.

Schulz-Anker:1970:SAS

[SA70] E. Schulz-Anker. Syntax-Antiqua, a sans serif on a new basis. Gebrauchsgraphik, 7(??):??, ???? 1970. ISSN 0016-5743.

Sicherman:1986:BFC

[SA86] G. L. (George L.) Sicherman and John Y. Arrasjid. The Buffalo font catalogue, version 4. Technical report. State University of New York at Buffalo. Dept. of Computer Science; 86-13. State University of New York at Buffalo. Dept. of Computer Science. Technical report; 86-13. State University of New York at Buffalo, Dept. of Computer Science, Buffalo, NY, USA, 1986. 15 pp.

Sanocki:1992:EFL

[San92] Thomas Sanocki. Effects of fontand letter-specific experience on the perceptual processing of letters. The American journal of psychology, 105(3):435–458, Fall 1992. CO-DEN AJPCAA. ISSN 0002-9556 (print), 1939-8298 (electronic).

Sauter:1986:BCM

[Sau86] John Sauter. Building Computer Modern Fonts. TUGboat, 7(3):151– 152, October 1986. ISSN 0896-3207.

Sauter:1987:UCM

[Sau87] John Sauter. Updated Computer Modern fonts for the LN03. TUGboat, 8(2):129–130, July 1987. ISSN 0896-3207.

Srouji:1992:AFD

[SB92] Johny Srouji and Daniel Berry.

Arabic formatting with ditroff/ ffortid. Electronic Publishing —Origination, Dissemination, and Design, 5(4):163–208, December 1992. CODEN EPODEU. ISSN 0894-3982.

Silver:1993:PRW

[SB93] N. Clayton Silver and Curt C. Braun. Perceived readability of warning labels with varied font sizes and styles. Safety Science, 16(5-6): 615–625, August 1993. CODEN SS-CIEO. ISSN 0925-7535.

Smura:1989:FII

[SBSG89] Edwin J. Smura, Barbara Beeton, Karla Savage, and Alan Griffee. Font information interchange standard ISO/IEC 9541. Computer Communications, 12(2):93– 96, April 1989. CODEN COCOD7. ISSN 0140-3664.

Schrod:1988:FSM

[Sch88] Joachim Schrod. LaTEX fonts and suggested magnifications. TUGboat, 9(3):241–243, November 1988. ISSN 0896-3207.

Schneider:1990:MME

[Sch90] Walter Schneider. MEL micro experimental laboratory, 1990. 4 computer disks leaf.

Schelin:1992:GSF

[Sch92a] Carl Schelin. Generating screen fonts. C Users Journal, 10(11):84– ??, November 1992. ISSN 0898-9788.

Schettini:1992:NFS

[Sch92b] Stephen Schettini. Nightmare on Font Street. Electronic Composition and Imaging, 6(??):50–53, October 1992. ISSN 0838-9535.

Schneider:1998:DSB

[Sch98a] Uwe Schneider. DaType: a strokebased typeface design system. Computers and Graphics, 22(4):515– 526, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cas/tree/
store/cag/sub/1998/22/4/583.
pdf.

Schneider:1998:OOM

[Sch98b] Uwe Schneider. An objectoriented model for the hierarchical composition of letterforms in computer-aided typeface design. Lecture Notes in Computer Science, 1375:109–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/1375/13750109.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/1375/
13750109.pdf.

Schneider:2000:HAS

[Sch00] Uwe Schneider. A hybrid approach for stroke-based letterform composition including outline-based methods. Computer Graphics Forum, 19 (4):??, December 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sherman:1986:PSU

[SD86] Mark Sherman and Robert L. Drysdale. Producing software using tools in a workstation environment. Technical Report PCS-TR86-134, Dartmouth College, Hanover, NH,

1986. URL mailto::reports@cs.
dartmouth.edu. (telephone +1

603-646-2206) (fax +1 603-646-1672) (hardcopy only).

Searfoss:1993:CFB

[Sea93] Glenn Searfoss. The computer font book. McGraw-Hill, Inc., New York, NY, USA, 1993. ISBN 0-07-881800-1. xxiv + 384 pp. LCCN Z 250.7 S43 1993. US$27.95.

Sebok:1991:ECC

[Seb91] Angelia L. Sebok. Effects of color CRT misconvergence and font type on text readability and subjective preference. Abstract. thesis (m.s.), Virginia Polytechnic Institute and State University, Blacksburg, VA, USA, 1991. vii + 100 pp.

Seymour:1994:JSS

[Sey94] Jim Seymour. Jim Seymour spends his free time fiddling with cool tool Font Chameleon. PC Week, 11(7): 64–??, February 1994. ISSN 0740-1604.

Sanfeliu:1988:ABH

[SFO88] A. Sanfeliu, J. Font, and I. Orteu. An architecture based on hybrid systems for analyzing 3D industrial scenes. In IEEE [IEE88], pages 368–370.

Spiekermann:2002:SSS

[SG02] Erik Spiekermann and E. M. Ginger. Stop Stealing Sheep & Find Out How Type Works. Adobe Press, Mountain View, CA, USA, second edition, 2002. ISBN 0-201-70339-4. 188 pp. LCCN Z250 .S738 1993. US$30.00.

Spiekermann:2003:SSS

[SG03] Erik Spiekermann and E. M. Ginger. Stop Stealing Sheep: Das Geheimnis guter Typografie. Markt & Technik, Haar bei München, Germany, 2003. ISBN 3-8272-5989-4. 220 pp. LCCN ????

Shaw:1994:WHF

[Sha94] Rick Shaw. Are we having font yet?

also: Apple’s Picard reviewed. AV video, 16(1):50–??, January 1994. ISSN 0747-1335.

Shamir:2003:CBA

[Sha03] Ariel Shamir. Constraint-based approach for automatic hinting of digital typefaces. ACM Transactions on Graphics, 22(2):131–151, April 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shaikh:2007:POT

[Sha07] Audrey Dawn Shaikh. Psychology of onscreen type: Investigations regarding typeface personality, appropriateness, and impact on document perception. Ph.D. dissertation, Department of Psychology, Wichita State University, Wichita, KS, USA, May 2007. xxviii + 624 pp. URL https:/

/www.proquest.com/pqdtglobal/
docview/304804882.

Sicherman:1985:BFC

[Sic85] G. L. (George L.) Sicherman. The Buffalo font catalogue, version 2. Technical report / University at Buffalo, State University of New York, Department of Computer Science; 85-04 Technical report (State University of New York at Buffalo. Dept. of Computer Science); 85-04. Dept. of Computer Science, University at Buffalo (SUNY), Buffalo, NY, USA, 1985. 14 pp.

Siegel:1985:EPS

[Sie85] David R. Siegel. The Euler Project at Stanford. Stanford University Library Archives, 1985.

Simonsen:1988:GFE

[Sim88] Mark Simonsen. GS font editor, 1988. 1 computer disk.

Shin:1993:CGA

[SK93] Sung Yong Shin and Tosiyasu L. Kunii, editors. Computer graphics and applications: proceedings of the First Pacific Conference on Computer Graphics and Applications, Pacific Graphics ’93 — vol. 1, Seoul, Korea, Aug. 30–Sep. 2, 1993, volume 1 of Computer Graphics and Applications. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1993. ISBN 981-02-1529-0. LCCN T385.C5575 1993.

Sando:2011:ZEC

[SK11] Minako Sando and Akira Kobayashi. Zapf Exhibition: the Calligraphy of Hermann and Gudrun Zapf. Japan Letter Arts Forum, Tokyo, Japan, 2011. 64 pp. LCCN ????

Slavinsky:2009:HST

[Sla09] Viacheslav Slavinsky. Hacking with style: TrueType VT220 font. Web site., 2009. URL

                                        http://svo.2.staticpublic.s3-
                                        website-us-east-1.amazonaws.
                                        com/glasstty/.

Sennhauser:1994:DIC

[SO94] Rene Sennhauser and Krystyna W. Ohnesorge. Document image compression using document analysis and block-class-specific data compression methods. Proceedings of SPIE — The International Society for Optical Engineering, 2186:146– 155, 1994. CODEN PSISDG. ISBN 0-8194-1481-6. ISSN 0277-786X (print), 1996-756X (electronic).

SoftCraft:1987:FF

[Sof87] SoftCraft, Inc. Fancy font, 1987.

6 computer disks font editing: Efont/Cfont user’s manual + 1 fancy WP user’s manual + 1 font documentation sheet + 1 starter font pack sheet + 1 Lotus conversion sheet + 1 Pfont summary sheet + 1 package content sheet + 1 product and font catalog + 1 license agreement sheet.

Solo:1983:CRA

[Sol83] Dan X. Solo. Classic Roman Alphabets: 100 Complete Fonts. Dover pictorial archive series. Dover, New York, NY, USA, 1983. ISBN 0-486-24517-9. LCCN Z250.5.R64 C57 1983.

Solutions:1987:C

[Sol87a] Solutions International, Inc. The curator, 1987.

Solutions:1987:S

[Sol87b] Solutions International, Inc. Super-Glue, 1987. 1 computer disk.

Solutions:1988:S

[Sol88] Solutions, Inc. SuperGlue, 1988.

ISBN 0-929100-01-8 (manual). 1 computer disk.

Somberg:1990:CAR

[Som90] Benjamin L. Somberg. Character aspect ratio and design tradeoffs. Proceedings of the Human Factors Society, pages 1461–1464, 1990. CODEN PHFSDQ. ISSN 0163-5182.

Southall:1985:DNT

[Sou85] Richard Southall. Designing new typefaces with Metafont. Report STAN-CS-TR-85-1074, Department of Computer Science, Stanford University, Stanford, CA, USA, September 1985. ii + 37 pp.

URL http://i.stanford.edu/
pub/cstr/reports/cs/tr/85/1074/
CS-TR-85-1074.pdf.

Southall:1986:DNT

[Sou86] Richard Southall. Designing a new typeface with METAFONT. In Désarménien [Dés86], pages 161– 179. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Southall:1998:MRC

[Sou98] Richard Southall. Metafont in the Rockies: The Colorado Typemaking Project. Lecture Notes in Computer Science, 1375:167–180, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/1375/13750167.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/1375/
13750167.pdf.

Southall:2005:PTT

[Sou05] Richard Southall. Printer’s type in the twentieth century: manufacturing and design methods. British Library, London, UK, 2005. ISBN 0-7123-4812-3 (British Library), 1-58456-155-6 (Oak Knoll Press). xv + 238 pp. LCCN Z250.A2 S68 2005.

Sowa:1991:BHB

[Sow91] Friedhelm Sowa. Bitmaps and halftones with BM2FONT. TUGboat, 12(34):534–538, November 1991. ISSN 0896-3207.

Smura:1987:TNB

[SP87] Edwin J. Smura and Archie D.

Provan. Toward a new beginning: the development of a standard for font and character encoding to control electronic document interchange. IEEE Transactions on Professional Communication, PC-30(4):259–264, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Sinha:1980:MDT

[SR80] R. M. K. Sinha and A. Raman. A modular data terminal for Indian languages. Computer Graphics, 14 (1–2):39–72, July 1980. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Shamir:1996:ETE

[SR96] A. Shamir and A. Rappoport. Extraction of typographic elements from outline representations of fonts. Computer Graphics Forum, 15(3):259–268, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shamir:1997:QED

[SR97] Ariel Shamir and Ari Rappoport.

Quality enhancements of digital outline fonts. Computers and Graphics, 21(6):713–725, November–December 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

URL http://www.elsevier.com/
cgi-bin/cas/tree/store/cag/cas_
sub/browse/browse.cgi?year=1997&
volume=21&issue=6&aid=9700049.

Shamir:1998:FBD

[SR98] Ariel Shamir and Ari Rappoport. Feature-based design of fonts using constraints. Lecture Notes in Computer Science, 1375:93–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/1375/13750093.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/1375/
13750093.pdf.

Sarfraz:2002:TSA

[SR02] M. Sarfraz and M. F. A. Razzak. Technical section: An algorithm for automatic capturing of the font outlines. Computers and Graphics, 26(5):795–804, October 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://www.

elsevier.com/gej-ng/10/13/20/
68/56/42/abstract.html.

Sugita:1983:MKG

[SSST83] Takuya Sugita, Hisao Sakamoto, Kimiko Shima, and Masahide Tsukamoto. Multi-font Kanji generator. Transactions of the Institute of Electronics and Communication Engineers of Japan, Section E (English), E66(6):377–382, June 1983. CODEN TIEEDU. ISSN 0387-236X.

Sakamoto:1985:CGHa

[ST85a] M. Sakamoto and M. Takagi. Computerized generation of high-quality fonts of the Japanese alphabets (in Japanese). Transactions of the Institute of Electronics, Information and Communication Engineers D-II, J68D(4):702–709, April 1985. CODEN DTGDE7. ISSN 0915-1923.

Sakamoto:1985:CGHb

[ST85b] Munekazu Sakamoto and Mikio Takagi. Computerized Generation of High-Quality Fonts of the Japanese Alphabets. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. ISBN 0-8186-0621-5 (paperback), 0-8186-8461-9 (hardback) (??invalid checksum??), 0-8186-4621-7 (microfiche). 786– 791 pp. LCCN QA 75.5 C5865 1985. IEEE Service Cent. Piscataway, NJ, USA.

Stauffacher:1985:TPK

[Sta85] Jack Stauffacher. The Transylvanian phoenix: The Kis–Janson types in the digital era. Visible Language, XIX(1):60–76, Winter 1985. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_
E.pdf.

StarMicronics:1987:NMF

[Sta87] Star Micronics Company. NX — 1000C multi — font printer users manual. Star Micronics Co., Shizuoka, Japan, 1987. 89 pp.

StarMicronics:1989:MFN

[Sta89] Star Micronics Company. Multi — font NX — 2410 users manual. Star Micronics Co., New York, NY, USA, 1989. 111 pp.

Stamm:1993:DRI

[Sta93a] Beat Stamm. Dynamic regularisation of intelligent outline fonts. Electronic Publishing—Origination, Dissemination, and Design, 6(3):219–230, September 1993. CO-DEN EPODEU. ISSN 0894-3982.

Stamm:1993:OEF

[Sta93b] Beat Stamm. Object-orientation and extensibility in a font-scaler. Electronic Publishing—Origination, Dissemination, and Design, 6 (3):159–170, September 1993. CO-DEN EPODEU. ISSN 0894-3982.

Stamm:1998:VTG

[Sta98] Beat Stamm. Visual TrueType: a graphical method for authoring font intelligence. Lecture Notes in Computer Science, 1375:77–92, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://link.springer-ny.
                                        com/link/service/series/0558/
                                        bibs/1375/13750077.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/series/0558/papers/1375/
                                        13750077.pdf.

Stewart:1914:TPI

[Ste14] Alexander A. Stewart. Type; a primer of information about the mechanical features of printing types: their sizes, font schemes, and c., with a brief description of their manufacture. Typographic technical series for apprentices, pt. 1, no. 1. Committee on apprentices, United typothetae and Franklin clubs of America, Boston, MA, USA, 1914. 42 pp.

Stewart:1918:TPI

[Ste18] Alexander A. Stewart. Type, a primer of information about the mechanical features of printing types: their sizes, font schemes, and c., with a brief description of their manufacture. Typographic technical series for apprentices, pt. I, no.

1. Committee on education, United typothetae of America, Chicago, IL, USA, 1918. 35 pp.

Steven:1987:PRC

[Ste87] Peter M. Steven. Problems of and requirements for character recognition systems. In RESNA ’87: Meeting the Challenge, Proceedings of the 10th Annual Conference on Rehabilitation Technology, San Jose, CA, USA, pages 666–668. RESNA-Assoc for the Advancement of Rehabilitation Technology, Washington, DC, USA, 1987.

Stevens:1992:FEM

[Ste92] Beth Stevens. Font effects: making and enhancing meaning in texts. Research paper, 3 hrs. thesis (m.a.), Ball State University, Muncie, IN, USA, 1992. i + 27 pp.

Stock-Allen:2016:CTH

[Sto16] Nancy Stock-Allen, In. Carol Twombly: Her Brief But Brilliant Career in Type Design. Oak Knoll Press, New Castle, DE, USA, 2016. ISBN 1-58456-346-X (hardcover). 176 pp. LCCN Z250.A2 T898 2016.

Strange:1993:AWD

[Str93a] Catherine L. Strange. Around the world in 80 days. typesetting Arabic, Hebrew, Cyrillic, and Thai. In IEEE [IEE93], pages 43–49 (or 43– 48??). ISBN 0-7803-1466-2. LCCN T 10.5 I54 1993. IEEE catalog number 93CH3367-0.

Strysko:1993:TF

[Str93b] Brian Strysko. Tool font. Thesis (b.s.), California Polytechnic State University, ????, 1993. ii + 25 pp.

Schuermann:1973:VLQ

[SV73] Juergen Schuermann and Leonore Vogt. Vergleich Linearer und quadratischer Klassifikatoren am Beispiel der Schriftart OCR-A. [comparison of linear and quadratic classifiers using the font OCR-A as an example]. NTZ Nachrichtentechnische Zeitschrift, NTZ Communications Journal, 26(6):269–274, June 1973. CODEN NNNCAZ.

Swzola:1993:ISC

[Swz93] S. Swzola. Info Select clears clutter: this PIM for Windows works the way you do. Byte Magazine, 18(5):97–98, Spring 1993. CODEN BYTEDJ. ISSN 0360-5280.

Symbolics:1981:FEP

[Sym81] Symbolics. Font editor, preliminary version. Technical report, Symbolics, Inc., Chatsworth, CA, September 1981.

Shao:1995:NCF

[SZ95] Lejun Shao and Hao Zhou.

A new contour fill algorithm for outlined character image generation. Computers and Graphics, 19(4):551–556, July– August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://

www.elsevier.com/cgi-bin/cas/
tree/store/cag/cas_sub/browse/
browse.cgi?year=1995&volume=
19&issue=4&aid=9500033.

Shao:1994:CFG

[SZC94] L. Shao, H. Zhou, and Kim Wah Chang. A Chinese font generation system. In Anonymous [Ano94q], pages 93–99. ISSN 0218-7019.

Tuescher:1982:LDS

[TA82] Daniel Tuescher and Max Arnold.

Latin/Arabic dual-font SP 300 ATU teleprinter. Hasler Review, 15 (3):79–81, Autumn 1982. CODEN HSLRA5. ISSN 0374-3306.

Tamari:1989:DHF

[Tam89] I. J. Tamari. Digitalisation of Hebrew fonts, or: Some evolutional evaluations. In André and Hersch [AH89], pages 188–197. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Taylor:1985:DDF

[Tay85] Don Taylor. Directory of digital font libraries. LPI Pub. Co., Agoura, CA, USA, second edition, 1985. 104 pp.

Taylor:1991:IMF

[Tay91] Philip Taylor. Improving the æsthetics of mixed-font documents. TUGboat, 12(1):35–44, March 1991. ISSN 0896-3207.

Tomayko:1990:AMG

[TBLJ90] James E. Tomayko, Bob Bemer, Ernie Lassen, and Robert Johnson. Anecdotes: More on General Electric’s start in the computer business; Dr. Robert Johnson interview; comments on interview. Annals of the History of Computing, 12(2):130–137, April/June 1990. CODEN AHCOE5. ISSN 0164-1239. URL http://dlib.

computer.org/an/books/an1990/
pdf/a2130.pdf;   http://www.
computer.org/annals/an1990/a2130abs.
htm.

Tao:2016:GBG

[TBP16] Nigel Tao, Chuck Bigelow, and Rob Pike. The Go blog: Go fonts. Web site, November 16,

2016. URL https://blog.golang.

org/go-fonts. This Web site announces the release of high-quality sansserif and monospaced font family, Go, designed by the Bigelow & Holmes type foundry for the Go Project, and under the same open source license as the rest of the Go project’s software.

Tseng:1992:EKS

[TC92] L. Y. Tseng and C. T. Chuang. Efficient knowledge-based stroke extraction method for multi-font Chinese characters. Pattern Recognition, 25(12):1445–1458, December 1992. CODEN PTNRA8. ISSN 0031-3203 (print), 1873-5142 (electronic).

Tou:1982:ATF

[TCLM82] Julius T. Tou, Jack M. Cheng, D. Z. Lu, and John McCarthy. Automatic type font identification for falsified documents. Proceedings — International Conference on Pattern Recognition, pages 1234–1240, 1982. CODEN PICREG. IEEE Service Cent. Piscataway, NJ, USA.

Teitelman:1985:TTC

[Tei85] W. Teitelman. A tour through Cedar. IEEE Transactions on Software Engineering, SE-11(3):285– 302, March 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Tenpenny:1990:EFF

[Ten90] Patricia Lynn Tenpenny. The effects of frequency, font, and context on episode-based repetition priming. Thesis (ph. d.), University of Illinois at Urbana-Champaign, Urbana-Champaign, IL 61801, USA, 1990. v + 117 pp.

TI:1986:TFL

[Tex86] Texas Instruments Incorporated.

TMS34010 Font library user’s guide. Texas Instruments, Houston, TX, USA, 1986. various pp.

FreeTypeProject:2000:FF

[The00] The FreeType Project. The FreeType 2 FAQ. World-Wide Web document, 2000. URL http:

//www.freetype.org/freetype2/
docs/ft2faq.html.

FreeTypeProject:2003:FAR

[The03] The FreeType Project. FreeType-2.1.7 API reference. World-Wide Web document, 2003. URL http:

                                        //www.freetype.org/freetype2/
                                        docs/reference/ft2-toc.html.

Thompson:1990:DQM

[Tho90] T. Thompson. A digital ‘quill’ for Mac video displays. Byte Magazine, 15(13):229–231, December 1990. CODEN BYTEDJ. ISSN 0360-5280.

Thomas:2019:BRD

[Tho19] Douglas Thomas. Book review: Douglas Thomas, Never use Futura. Princeton Architectural Press, 2017, paperback, 208pp., US$24.95, ISBN 978-1-61689-572-3. TUGboat, 40(1):??, ???? 2019. ISSN 0896-3207. URL

                                        http://tug.org/books/reviews/
                                        tb124reviews-thomas.html.

Timeworks:1987:FP

[Tim87] Timeworks, Inc. Font pack 1, 1987. 1 computer disk user’s manual + 1 instruction card.

Timeworks:1988:FP

[Tim88] Timeworks, Inc. Font pack 2, 1988. 2 computer disks user’s manual + instruction card.

Timeworks:1992:PIW

[Tim92] Timeworks, Inc. Publish it! for windows, 1992. 5 computer disks + reference guide (1 folded card) + ATM user guide (iv, 24 p. : ill. ; 19 cm.) + ATM quick installation (1 card).

Tachikawa:1983:MKS

[TKE83] Michiyoshi Tachikawa, Tatsuo Kasahara, and Koichi Ejiri. Multifont Kanji synthesis by chain coded strokes. In Japan Display ’83, Proceedings of the 3rd International Display Research Conference, Kobe, Japan, pages 172–175. Soc for Information Display, Los Angeles, Calif, USA, 1983.

Takayama:1995:DFC

[TKMM95] K. Takayama, H. Kano, Y. Maeda, and K. Misue. Dynamic font: Its concept and generation method. In Anzai et al. [AOM95], pages 135– 140. ISBN 0-444-81795-6. ISSN 0921-2647. LCCN QA76.9.A73 I564 1995.

Trepos:1991:SFF

[TL91] R. Trépos and Ph. Louarn. Serveurs de fichiers et de fontes pour TEX (English: File and font servers for TEX). Cahiers GUTenberg, 8:21–28, mars 1991. ISSN 1140-9304.

Tou:1992:ICC

[TL92a] J. T. Tou and J. J. Liang, editors. International conference on computer processing of Chinese and Oriental languages — 1992 Dec: Gainesville; FL, INTER-NATIONAL CONFERENCE ON COMPUTER PROCESSING OF CHINESE AND ORIENTAL LAN-GUAGES. Chinese Language Computer Society; 1992, ????, 1992.

TsangKaChe:1992:FFE

[TL92b] E. Tsang Ka Che and C. H. Lee.

A feature-rich font editor for Asian characters. In Tou and Liang [TL92a], pages 319–324.

Thomas:2017:NUFa

[TL17] Douglas Thomas and Ellen Lupton. Never use Futura. Princeton Architectural Press, New York, NY, 2017. ISBN 1-61689-572-1 (paperback), 1-61689-666-3 (ebook). 207 pp. LCCN Z250.5.F88 T48 2017.

Tobin:1984:ORF

[Tob84] Georgia K. M. Tobin. The OCLC Roman family of fonts. TUGboat, 5 (1):36, May 1984. ISSN 0896-3207.

Tobin:1989:HLF

[Tob89] Georgia K. M. Tobin. A handy little font. TUGboat, 10(1):28–30, April 1989. ISSN 0896-3207.

Tofsted:1990:ICF

[Tof90] David Tofsted. An improved chess font. TUGboat, 11(4):542–544, November 1990. ISSN 0896-3207.

Troen:1990:AED

[TP90a] Ib Troen and Erik Lundtang Petersen. Atlante Europeo del Vento. Risø National Laboratory, Roskilde, Denmark, 1990. ISBN 87-550-1639-1. 264 pp. Typeset with Bitstream fonts on a Canon Series III 300-dpi laser printer, with extensive graphics and data tables; the data files are available on IBM PC diskettes.

Troen:1990:AEE

[TP90b] Ib Troen and Erik Lundtang Petersen. El Atlas Eólico Europeo. Risø National Laboratory, Roskilde, Denmark, 1990. ISBN 87-550-1638-3. 260 pp. Typeset with Bitstream fonts on a Canon Series III 300-dpi laser printer, with extensive graphics and data tables; the data files are available on IBM PC diskettes.

Troen:1990:EW

[TP90c] Ib Troen and Erik Lundtang Petersen. Europäischer Windatlas. Risø National Laboratory, Roskilde, Denmark, 1990. ISBN 87-550-1636-7. 240 pp. Typeset with Bitstream fonts on a Canon Series III 300-dpi laser printer, with extensive graphics and data tables; the data files are available on IBM PC diskettes.

Troen:1991:AE

[TP91] Ib Troen and Erik Lundtang Petersen. Atlas Éolien Européen. Risø National Laboratory, Roskilde, Denmark, 1991. ISBN 87-550-1637-5. 278 pp. Typeset with Bitstream fonts on a Canon Series III 300-dpi laser printer, with extensive graphics and data tables; the data files are available on IBM PC diskettes.

Travers:1989:VRK

[Tra89] M. Travers. A visual representation for knowledge structures. In ACM Hypertext’89 Proc., Implementations and Interfaces, page 147. ACM Press, New York, NY, USA, 1989.

Trembly:1994:NPL

[Tre94] Ara Trembly. Nonimpact printers let you put your best font forward. Office systems: the magazine for small and medium offices, 11(6):48– ??, June 1994. ISSN 8750-3441.

Trickett:1969:DST

[Tri69] T. Trickett. Design of standard type font for optical character recognition. Honeywell Computer J, 3(1): 3–11, Winter 1969.

Touati:1985:PFE

[TS85] Herve Touati and Shin Sedai Konpyuta Gijutsu Kaihatsu Kiko (Japan). PSI font editor implementation notes. Technical report, Institute for New Generation Computer Technology, Tokyo, Japan, 1985. 114 pp.

Timbrell:1994:DR

[TS94] Nicola Timbrell and David Snowling. The dream of the rood. University of Oxford, Oxford University Computing Services, Oxford, UK, hypermedia edition edition, 1994. ISBN 1-85889-008-X. LCCN ???? 1 portfolio.

Turner:2000:DF

[Tur00a] David Turner. The design of FreeType 2. World-Wide Web document, 2000. URL http:

                                        //www.freetype.org/freetype2/
                                        docs/design/index.html.

Turner:2000:FGC

[Tur00b] David Turner. FreeType glyph conventions: Version 2.1. World-Wide Web document, 2000. URL http:

                                        //www.freetype.org/freetype2/
                                        docs/glyphs/index.html.

Turner:2003:FT

[Tur03] David Turner. FreeType 2 tutorial. World-Wide Web document, 2003. URL http://www.

freetype.org/freetype2/docs/
tutorial/step1.html;   http:
//www.freetype.org/freetype2/
docs/tutorial/step2.html.

Tutelaers:1992:FST

[Tut92] Piet Tutelaers. A font and a style for typesetting chess using LaTEX or TEX. TUGboat, 13(1):85–90, April 1992. ISSN 0896-3207.

Tang:1992:DAT

[TX92] T. Tang and J. Xu. The design of an alternate tool for making the character font. In Tou and Liang [TL92a], pages 197–202.

Uphaus:1995:ITA

[UB95] James A. Uphaus, Jr. and Richard Barry Bryant. Innovative tool for assessing the impact of cluster defects on dot-matrix display cockpit symbology. IEEE Proceedings of the National Aerospace and Electronics Conference, 1:558–565, 1995. CO-DEN NASEA9. IEEE catalog number 95CH35797.

Uphaus:1990:ICR

[UBR90] James A. Uphaus, Jr., Kristen Barthelemy, and John M. Reising. Improve character readability in spite of pixel failures — a better font. IEEE Proceedings of the National Aerospace and Electronics Conference, 1:278–283, 1990. CO-DEN NASEA9. IEEE catalog number 90CH2881-1.

Uehara:1990:CTP

[Ueh90] T. Uehara. Current technology and problems in computer fonts. Journal of Information Processing, 31(11):1570–1580, 1990. CODEN JOSHA4. ISSN 0447-8053.

Uchikata:1984:HCF

[UKKS84] Hideki Uchikata, Seiji Koyama, Takayuki Kageyama, and Masahiro Sasaki. High-quality character font generation system. National Technical Report (Matsushita Electric Industry Company), 30(3):124–132, June 1984. CODEN NTROAV. ISSN 0028-0291.

Umeda:1985:CMP

[UM85] Michio Umeda and Shinichi Meguro. Construction of a multifont printed Chinese character reader. Denki Tsushin Kenkyusho Kenkyu Jitsuyoka Hokoku, 34(1): 47–58, 1985. CODEN DTKKAA. ISSN 0415-3200.

Umeda:1982:RMP

[Ume82] Michio Umeda. Recognition of multi-font printed Chinese characters. Proceedings — International Conference on Pattern Recognition, pages 793–796, 1982. CODEN PI-CREG. IEEE Service Cent. Piscataway, NJ, USA.

Unger:2018:TTD

[Ung18] Gerard Unger. Theory of Type Design. nai010 publishers, Rotterdam, The Netherlands, 2018. ISBN 94-6208-440-8. 239 pp. LCCN Z250 .U48 2018.

UC:2000:SIUb

[Uni00] Unicode Consortium, editor. Seventeenth International Unicode Conference (IUC17) Unicode and the Web: the Global Connection, September 5–8, 2000, San Jose, California. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-inc@ unicode.org, 2000. ISBN ????

LCCN ???? URL http://www.
unicode.org/iuc/iuc17.

Updike:1922:PTT

[Upd22] Daniel Berkeley Updike. Printing Types; Their History, Forms, and Use; a Study in Survivals. Harvard University Press, Cambridge, MA, USA, 1922. ???? pp. LCCN Z250.A2 U6 1922. URL

http://catdir.loc.gov/catdir/
toc/fy1002/37032668.html.

Updike:1937:PTT

[Upd37] Daniel Berkeley Updike. Printing Types; Their History, Forms, and Use; a Study in Survivals. Harvard University Press, Cambridge, MA, USA, second edition, 1937. ISBN 0-486-23928-4, 0-486-23929-2. LCCN Z250.A2 U6 1937. URL

http://catdir.loc.gov/catdir/
toc/fy1002/37032668.html.

Updike:1962:PTT

[Upd62] Daniel Berkeley Updike. Printing Types, Their History, Forms, and Use: a Study in Survivals. Harvard University Press, Cambridge, MA, USA, third edition, 1962. ISBN 0-486-23929-2 (paperback). LCCN Z250.A2 U6 1962. Two volumes.

Updike:1980:PTT

[Upd80] Daniel Berkeley Updike. Printing Types: Their History, Forms and Use: a Study in Survivals. Dover, New York, NY, USA, second edition, 1980. ISBN 0-486-23928-4 (volume 1), 0-486-23929-2 (volume 2). xxxiv + 292 (volume 1), xvii + 326 (volume 2) pp. LCCN ????

USENIX:1987:FCG

[USE87] USENIX Association, editor. Fourth Computer Graphics Workshop Proceedings (Cambridge, MA, USA). USENIX Association, Berkeley, CA, USA, October 8–9, 1987. ISBN ???? LCCN ????

USENIX:1989:UFC

[USE89] USENIX Association, editor. USE-NIX Fifth Computer Graphics workshop (Monterey, CA, USA). USENIX Association, Berkeley, CA, USA, November 16–17, 1989. ISBN ???? LCCN ????

Ulrich:2017:PCN

[USK^{+}17] Ferdinand P. Ulrich, Jack Stauffacher, Donald Knuth, Hermann Zapf, and John Dreyfus. From punch cutters to number crunchers. Eye, 24(94):??, 2017. URL http:/

/www.eyemagazine.com/feature/
article/from-punch-cutters-to-
number-crunchers.

VanVliet:1986:TPD

[Van86] J. C. Van Vliet, editor. Text processing and document manipulation: Proceedings of the international conference, University of Nottingham, 14–16 April 1986. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-511-66313-7 (e-book), 0-521-11031-9 (paperback), 0-521-32592-7 (hardcover). LCCN QA76.9.T48 T48 1986.

Vanderburg:1988:SUV

[Van88] Glenn Vanderburg. Some useful variations of standard fonts. TUGboat, 9(2):125, August 1988. ISSN 0896-3207.

VanHorn:1992:CSA

[Van92a] G. Armour Van Horn. CorelDraw 3.0: a Swiss Army knife for illustrators. Byte Magazine, 17(10):223– 224, 226, October 1992. CODEN BYTEDJ. ISSN 0360-5280.

Vannatta:1992:BBB

[Van92b] R. Vannatta. Borland builds a better Quattro Pro. Byte Magazine, 17 (7):267–268, 270, July 1992. CO-DEN BYTEDJ. ISSN 0360-5280.

Horn:1993:CWI

[Van93a] G. Armour Van Horn. CorelDraw 4.0: The word is more: Corel releases an ambitious upgrade. Corel-Draw 4.0 features tighter integration across modules, multiple page illustrations, an underlying object database, an impressive bundle of fonts and clip art, path-based animation, and even OCR and forms generation. Van Horn finds out how it all comes together. Byte Magazine, 18(10):169–??, September 1993. CODEN BYTEDJ. ISSN 0360-5280.

VanHorn:1993:CWI

[Van93b] G. Armour Van Horn. CorelDraw 4.0: The word is more: Corel releases an ambitious upgrade. Corel-Draw 4.0 features tighter integration across modules, multiple page illustrations, an underlying object database, an impressive bundle of fonts and clip art, path-based animation, and even OCR and forms generation. Van Horn finds out how it all comes together. Byte Magazine, 18(10):169–??, September 1993. CODEN BYTEDJ. ISSN 0360-5280.

VonEhr:1987:FBF

[VCA87] Jim Von Ehr, Kevin Crowder, and Altsys Corporation. FONTastic bitmap font editor for the Macintosh, 1987. 1 computer disk.

VonEhr:1988:FPF

[VCA88] Jim Von Ehr, Kevin Crowder, and Altsys Corporation. Fontographer professional font editor, 1988. 1 computer disk.

VonEhr:1989:FPF

[VCG^{+}89] Jim Von Ehr, Kevin Crowder, Katharine Green, Earl Allen, and Altsys Corporation. Fontographer professional font editor, 1989. 2 computer disks + 1 ABCs tutorial (ii, 39, ii p.: ill.; 22 cm.) + 1 quick reference guide.

VonEhr:1990:FPF

[VCGA90] Jim Von Ehr, Kevin Crowder, Katharine Green, and Altsys Corporation. Fontographer professional font editor, 1990. 2 computer disks + 1 ABCs tutorial (iii, 73 p.: ill.; 22 cm.) + 1 addendum (Version 3.5, 1992; 76 p.: ill.; 22 cm.) + 1 quick reference guide.

Veron:1993:CFE

[VK93] H. Veron and J. M. Kistner. Comparative font evaluation based on observer performance and spectral analysis. In Morreale [Mor93], pages 518–520. ISSN 0097-966X (print), 2154-6738 (electronic), 2154-6746.

Bechtolsheim:1992:PPM

[von92a] Stephan von Bechtolsheim. TEX in Practice: Paragraphs, Math, and Fonts, volume 2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97596-9. 384 pp. LCCN Z253.4.T47 B4 1993. US$49.00.

vonBechtolsheim:1992:TPP

[von92b] Stephan von Bechtolsheim. TEX in Practice: Paragraphs, Math, and Fonts, volume 2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97596-9. 384 pp. LCCN Z253.4.T47 B4 1993. US$49.00.

vonBechtolsheim:1993:TPP

[von93] Stephan von Bechtolsheim. TEX in Practice: Paragraphs, Math, and Fonts, volume 2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-97596-9, 3-540-97596-9. LCCN Z253.4.T47 B4 1993.

Vincent:1994:DR

[VP94a] Luc M. Vincent and Theo Pavlidis, editors. Document Recognition, volume 2181. Society of Photo-optical Instrumentation Engineers (SPIE), Bellingham, WA, USA, 1994. CO-DEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TK7895.O6 D63 1994.

Vincent:1994:DRF

[VP94b] Luc M. Vincent and Theodosios Pavlidis, editors. Document recognition: 9–10 February 1994, San Jose, California, number 2181 in Proceedings / SPIE — The International Society for Optical Engineering. SPIE, ????, 1994. ISBN 0-8194-1476-X. ISSN 0361-0748. LCCN TK7895.O6 D63 1994.

Wade:1984:DFS

[Wad84] B. W. Wade. Digital font scaling method. IBM Technical Disclosure Bulletin, 26(9):4863–??, February 1984. CODEN IBMTAA. ISSN 0018-8689.

Waegel:1992:UT

[Wae92] Janet Waegel. Using type. Business publishing, 8(12):46–??, December 1992. ISSN 1060-2208.

Wagner:1985:DDM

[Wag85] Annette Wagner. Design Decisions Made in the Development of Type for Apple’s Lisa Computer. Natl Computer Graphics Assoc, Fairfax, VA, USA, 1985. ISBN 0-941514-09-

9. 491–501 pp. LCCN ????

Wallin:1905:SID

[Wal05] J. E. Wallace Wallin. The size illusion of the depressed letter P. Scientific American, 93(17): 315, October 21, 1905. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

                                        URL http://www.nature.com/
                                        scientificamerican/journal/v93/
                                        n17/pdf/scientificamerican10211905-
                                        315a.pdf.

Walters:2013:FTC

[Wal13] John L. Walters. Fifty Typefaces That Changed the World. Design Museum. Conran Octopus, London, UK, 2013. ISBN 1-84091-629-X (hardcover). 106 pp. LCCN Z250.A2 W38 2013. URL http://

www.loc.gov/catdir/enhancements/
fy1507/2014430507-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy1507/2014430507-d.html.

Wang:1980:FC

[Wan80] Diana Shu-Fan Wang. Font compression. Abstract ([1] leaf) bound with copy. thesis (m.s.), University of Kentucky, Lexington, KY, USA, 1980. iv + 33 pp.

Warnock:1980:DCU

[War80] J. E. Warnock. The display of characters using gray level sample arrays. Computer Graphics, 14 (3):302–307, July 1980. CODEN CGRADI. ISSN 0097-8930.

MicroWare:1984:FLE

[War84] Micro Ware. Font down loader editor library disk, 1984. 1 program file (Apple) on 1 computer disk.

Warde:2009:CGW

[War09] Beatrice Warde. The crystal goblet, or why printing should be invisible. In Helen Armstrong, editor, Graphic Design Theory: Readings from the Field, pages 39–43. Princeton Architectural Press, New York, NY, USA, 2009. ISBN 1-56898-772-2, 1-61689-123-8 (e-book). LCCN NC997 .G673 2009.

Wayzata:1992:FFH

[Way92a] Wayzata Technology. Font fun house, 1992. ISBN 1-56737-039-X. 1 computer laser optical disk TextWare instructions.

Wayzata:1992:FPE

[Way92b] Wayzata Technology, Inc. Font pro essentials, 1992. ISBN 1-56737-048-9. 1 computer laser optical disk.

Wayzata:1992:FPT

[Way92c] Wayzata Technology, Inc. Font pro types of distinction, 1992. ISBN 1-56737-049-7. 1 compact disk.

White:1995:PFS

[WB95] Ronald G. White and John Biard. A portable font specification. Dr. Dobb’s Journal of Software Tools, 20(3):28, 30, 32, 34, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Wong:1982:DAS

[WCW82] K. Y. Wong, R. G. Casey, and F. M. Wahl. Document Analysis System. IBM Journal of Research and Development, 26(6):647– 656, November 1982. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Wyatt:2018:TMR

[WD18] Christopher Scott Wyatt and Dànielle Nicole DeVoss, editors. Type Matters: the Rhetoricity of Letter Forms. Visual rhetoric. Parlor Press, Anderson, SC, USA, 2018. ISBN 1-60235-973-3 (paperback), 1-60235-977-6 (hardcover), 1-60235-978-4 (e-book), 1-60235-979-2 (e-pub), 1-60235-980-6 (iBook), 1-60235-981-4 (Kindle). xxi + 388 pp. LCCN Z250 .T935 2018.

Waldo:1987:GFF

[WDL87] James Waldo, Marcia Delaney, and John Laporta. A generalized font file format. In USENIX Association [USE87], pages 107–?? ISBN ???? LCCN ???? Abstract only.

Weber:2015:HZR

[Web15a] Andreas Weber. Hermann Zapf remembered: Mastery through precision and passion. or: In the beginning was the word, and the word found a new beginning. Web site., June 2015. URL http:

//www.linotype.com/645-34858/
hermannzapfsobituary.html.

Weber:2015:HZD

[Web15b] Bruce Weber. Hermann Zapf, 96, dies; designer whose letters are found everywhere. New York Times, ??(??):??, June 9, 2015. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http:

//www.nytimes.com/2015/06/10/
arts/design/hermann-zapf-96-
dies-designer-whose-letters-
are-found-everywhere.html.

Wendt:1967:RPR

[Wen67] Dirk Wendt. Research in progress:

Research on readability and congeniality of print. The Journal of Typographic Research, I (3):327, July 1967. ISSN 0449-

329X. URL https://s3-us-west-
2.amazonaws.com/visiblelanguage/
pdf/V1N3_1967_E.pdf.

Wendt:1968:SDT

[Wen68] Dirk Wendt. Semantic differentials of typefaces as a method of congeniality research. The Journal of Typographic Research, II(1): 3–25, January 1968. ISSN 0449-

                                        329X. URL https://s3-us-west-
                                        2.amazonaws.com/visiblelanguage/
                                        pdf/V2N1_1968_E.pdf.

Wolcott:1976:CCT

[WH76] Norman M. Wolcott and Joseph Hilsenrath. A contribution to computer typesetting techniques: tables of coordinates for Hershey’s repertory of occidental type fonts and graphic symbols. Technical Report Special Publication 424, U.S. National Bureau of Standards, Gaithersburg, MD, USA, April 1976. 169 pp.

Williams:1986:FSL

[WH86] R. Don Williams and R. D. Huchingson. Font style for large-area discrete-element displays. Digest of Technical Papers — SID International Symposium (Society for Information Display), 17:146–149, 1986. CODEN DTPSDS.

Wong:1995:DCT

[WH95] Paul Yiu Chung Wong and Siu Chi Hsu. Designing Chinese typeface using components. Proceedings — IEEE Computer Society’s International Computer Software & Applications Conference, pages 416–421, 1995. CODEN PSICD2. ISSN 0730-6512. IEEE catalog number 95CB35838.

Whitman:1927:FT

[Whi27] Walt Whitman. A font of type. The Laboratory Press, Pittsburgh, PA, USA, 1927. various pp.

Whitman:1947:FT

[Whi47] Walt Whitman. A font of type.

Diamant Typographic Service, New York, NY, USA, 1947. 4 pp.

Whitney:1981:FU

[Whi81] Ronald Whitney. Font update.

TUGboat, 2(3):35, November 1981. ISSN 0896-3207.

White:1987:HST

[Whi87] Alex White. How to SPEC Type.

Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 1987. ISBN 0-8230-2473-3. 128 pp. LCCN Z253.W46 1987. US$14.95.

Whitman:1991:FT

[Whi91] Walt Whitman. A font of type.

????, United States, 1991. 1 pp.

Williams:1980:FMS

[Wil80] Rodney Don Williams. Font and matrix size for distance viewed dotmatrix displays. “major subject: Industrial engineering.” thesis (ph. d.), Texas A and M University, College Station, TX, USA, 1980. x + 123 pp.

Wilcox:1989:MML

[Wil89] Patricia Wilcox. METAPLOT:

Machine-independent line graphics for TEX. TUGboat, 10(2):179–187, July 1989. ISSN 0896-3207.

Wilhelms:1991:SSF

[Wil91] Gerhard Wilhelms. Source of stroked fonts for Borland’s graphic interface. C Users Journal, 9(5): 136–??, May 1991. ISSN 0898-9788.

Wilson:1993:OS

[Wil93] Donald Wilson. Over our shoulder. Publish!: the how-to magazine of desktop publishing, 8(4):96– ??, April 1993. CODEN PBLIEJ. ISSN 0897-6007.

Williams:1994:HBY

[Wil94] Robin Williams. How to boss your fonts around: a primer on font technology and font management on the Macintosh. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1994. ISBN 1-56609-102-0. viii + 152 pp. LCCN ????

Williams:2004:NDD

[Wil04] Robin Williams. The non-designers design book: design and typographic principles for the visual novice. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, second edition, 2004. ISBN 0-321-19385-7. 191 pp. LCCN Z246 .W634 2004.

Williams:2006:VBG

[Wil06a] Paul Flo Williams. VT220 built-in glyphs. Web site, 2006. URL http:

//vt100.net/dec/vt220/glyphs.

Williams:2006:NDT

[Wil06b] Robin Williams. The nondesigner’s type book: insights and techniques for creating professionallevel type. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, second edition, 2006. ISBN 0-321-30336-9 (paperback). 239 pp. LCCN Z250.7 .W55 2006.

Winckler:1990:FIG

[Win90] Jörg Winckler. TEX fonts in image generation software. In Clark [Cla90]. ISBN 0-13-912296-6.

Witten:1988:FD

[Wit88] Ian H. Witten. Font design. Technical Report 88/322/34, University of Calgary, October 1988. URL

mailto::parin@cpsc.ucalgary.
ca.

Wolberg:1986:SOC

[Wol86] George Wolberg. A syntactic omnifont character recognition system. In CVPR 1986 [CVP86], pages 168– 173. ISBN 0-8186-0721-1. LCCN TA1632 .I36 1986. IEEE Service Cent. Piscataway, NJ, USA.

Wolberg:1987:SOC

[Wol87] G. Wolberg. A syntactic omni-font character recognition system. International Journal of Pattern Recognition and Artificial Intelligence, 1: 303–322, 1987. CODEN IPJIEI. ISSN 0218-0014.

Wong:1985:FSCa

[Won85a] Willard Lee Wong. Font storage, compression, and scaling for networked professional workstations. Thesis (m.s.), University of Illinois at Urbana-Champaign, Urbana-Champaign, IL 61801, USA, 1985. v + 80 pp.

Wong:1985:FSCb

[Won85b] Willard Lee Wong. Font storage, compression, and scaling for networked professional workstations. Report UIUCDCS-R-85-1201, Dept. of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, April 1985. v + 80 pp.

Wong:1993:TCP

[Won93] Wai Wong. Typesetting Chinese pinyin using virtual fonts. TUGboat, 14(1):8–11, April 1993. ISSN 0896-3207.

Irving:1987:FOF

[Woo87] William W. (William Wood). Foundation of open font character recognition system. Thesis (b.s.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, ????, 1987. 74 pp. Supervised by David H. Staelin.

Wood:1989:WWP

[Woo89] L. Wood. A window on word processing. Byte Magazine, 14(5):221– 223, May 1989. CODEN BYTEDJ. ISSN 0360-5280. WordStarInternational:1990:W

[Wor90] WordStar International. WordStar 2000, 1990. 17 computer disks.

Wrenholt:1987:PLU

[Wre87] John Wrenholt. PS lover’s utility set, 1987. 1 computer disk.

Wrenholt:1988:PLUa

[Wre88a] John Wrenholt. PS lover’s utility set, 1988. 1 computer disk.

Wrenholt:1988:PLUb

[Wre88b] John Wrenholt. PS lover’s utility set, 1988. 1 computer disk.

Wright:1983:UAW

[Wri83] Walter Wright. Using an artists’ workstation. Proceedings — Symposium on Small Computers in the Arts, pages 128–142, 1983. CO-DEN PSSAED. ISBN 0-8186-0499-9. IEEE Service Cent. Piscataway, NJ, USA.

Weisteen:1992:FSN

[WT92] K. Weisteen and T. Taxt. Font segmentation of non-rotated, printed symbols in the vector representation. In IEEE [IEE92], pages 669– 672. ISBN 0-8186-2930-4, 0-8186-2917-7, 0-8186-2915-0. ISSN 1051-4651. LCCN Q 327 I615 1992. Four volumes.

Williams:2006:NDC

[WT06a] Robin Williams and John Tollett. The non-designer’s collection. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 2006. ISBN 0-321-41276-1 (Set), 0-321-19385-7 (design book), 0-321-30336-9 (type book), 0-321-30337-7 (Web book). ???? pp. LCCN XJ.T06.

Williams:2006:NDW

[WT06b] Robin Williams and John Tollett.

The non-designer’s Web book: an easy guide to creating, designing, and posting your own Web site. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, third edition, 2006. ISBN 0-321-30337-7. 335 pp. LCCN TK5105.888 .W5536 2006.

Wujastyk:1988:MFS

[Wuj88] Dominik Wujastyk. The many faces of TEX: A survey of digital METAfonts. TUGboat, 9(2):131–151, August 1988. ISSN 0896-3207.

Wujastyk:1989:FN

[Wuj89] Dominik Wujastyk. Font news. TUGboat, 10(1):39, April 1989. ISSN 0896-3207.

Wilkes:1982:SCS

[WW82] A. J. Wilkes and N. E. Wiseman. A soft-edged character set and its derivation. The Computer Journal, 25(1):140–147, February 1982. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Wilkinson:1993:USR

[WW93] R. A. Wilkinson and Charles L. Wilson. Using self-organizing recognition as a mechanism for rejecting segmentation errors. Proceedings of SPIE — The International Society for Optical Engineering, 1825:378– 388, 1993. CODEN PSISDG. ISBN 0-8194-1026-8. ISSN 0277-786X (print), 1996-756X (electronic).

Xerox:1979:XEP

[Xer79] Xerox Corporation. Xerox 9700 Electronic Printing System: font user’s guide. Xerox, El Segundo, CA, USA, 1979. iv + 65 pp.

Xerox:1980:XEP

[Xer80] Xerox Corporation. Xerox Electronic Printing System: special font bulletin. Xerox, El Segundo, CA, USA, 1980. various pp.

Xerox:1982:XEPa

[Xer82a] Xerox Corporation. Xerox 8700/ 9700 electronic printing system font users guide. Xerox, El Segundo, CA, USA, 1982. various pp.

Xerox:1982:XEPb

[Xer82b] Xerox Corporation. Xerox 8700/ 9700 electronic printing system: standard font library font users guide. Xerox, El Segundo, CA, USA, 1982. various pp.

Yi:1992:ECN

[Yi92] S. Yi. EHE criterion in neural network approach for Chinese font restoration from out-of-focused image. In Tou and Liang [TL92a], pages 397–400.

Yuzuguchi:1989:DTP

[Yuz89] A. Yuzuguchi. Desk top publishing now in Japan. Journal of the Society of Instrument and Control Engineers, 28(3):199–203, March 1989. CODEN KESEA4. ISSN 0453-4662.

Youngen:1989:MCM

[YWL89] R. E. Youngen, W. B. Woolf, and D. C. Latterner. Migration from Computer Modern fonts to Times fonts. TUGboat, 10(4):513–519, December 1989. ISSN 0896-3207.

Zapf:1985:FTT

[Zap85] Hermann Zapf. Future tendencies in type design: The scientific approach to letterforms. Visible Language, XIX(1):23–34, Winter 1985. CODEN VSLGAO.

ISSN 0022-2224 (print), 2691-5529 (electronic). URL https:

//s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V19N1_1985_
E.pdf.

Zapf:2000:SDS

[Zap00] Hermann Zapf. Vom Stempelschnitt zur Digitalisierung von Schriftzeichen. Die technischen Veränderungen der Schriftherstellungen. (German) [From the die cut to digitization of characters. The technical changes of written productions]. Gutenberg-Gesellschaft, Mainz, Germany, 2000. ???? pp.

Zapf:2001:MCD

[Zap01] Hermann Zapf. My collaboration with Don Knuth and my font design work. TUGboat, 22(1/2):26– 30, March 2001. ISSN 0896-3207.

URL https://tug.org/TUGboat/
tb22-1-2/tb70zapf.pdf.

Zapf:2007:ASCa

[Zap07a] Hermann Zapf. Alphabet Stories: a Chronicle of Technical Developments. Mergenthaler Edition, Linotype GmbH, Bad Homburg, Germany, 2007. ISBN 3-9810319-6-2. 149 pp. LCCN ????

Zapf:2007:ASCb

[Zap07b] Hermann Zapf. Alphabet Stories: a Chronicle of Technical Developments. RIT Cary Art Graphics Press, Rochester, NY, USA, 2007. ISBN 1-933360-29-1. 149 pp. LCCN Z250.A2 Z37613

2007. URL http://hdl.library.
upenn.edu/1017.12/366299.

Zapf:2007:ACT

[Zap07c] Hermann Zapf. Alphabetgeschichten:

Eine Chronik technischer Entwicklungen. Mergenthaler Edition, Linotype GmbH, Bad Homburg, Germany, 2007. ISBN 3-9810319-5-4. 156 pp. LCCN ???? 29.00 EUR.

Zapf:20xx:LSH

[Zapxx] Hermann Zapf. The life story of Hermann Zapf: A new start in Darmstadta. Linotype Web site, 20xx. URL https:/

/download.linotype.com/free/
howtouse/ZapfBiography.pdf.

Zramdini:1993:OFR

[ZI93] Abdel Wahab Zramdini and Rolf Ingold. Optical font recognition from projection profiles. Electronic Publishing—Origination, Dissemination, and Design, 6(3):249– 260, September 1993. CODEN EPODEU. ISSN 0894-3982.

Zramdini:1995:APO

[ZI95] A. Zramdini and R. Ingold.

ApOFIS: an A priori Optical Font Identification System. In Braccini et al. [BDV95], pages 527– 532. ISBN 3-540-60298-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1632 .I552 1995.

Zable:1987:FDH

[ZL87] J. L. Zable and H. C. Lee. Font design for high-speed impact line printers. IBM Journal of Research and Development, 31(6):679– 683 (or 679–684??), December 1987. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Zlatuvska:1991:AGV

[Zla91] J. Zlatuška. Automatic generation of virtual fonts with accented letters for TEX. Cahiers GUTenberg, 10-11: 57–68, septembre 1991. ISSN 1140-9304.

Zhang:1984:FSB

[ZST84] X. Zhang, H. Sanada, and Y. Tezuka. Forming square-styled brushwritten Chinese characters with a computer (in Japanese). Transactions of the Institute of Electronics, Information and Communication Engineers D-II, J67D:599–606, May 1984. CODEN DTGDE7. ISSN 0915-1923.

vZubrinic:1996:CF

[Žub96] Darko Žubrinić. Croatian fonts. TUGboat, 17(1):29–33, March 1996. ISSN 0896-3207.

Znamenskaya:1996:REP

[ZZ96] L. N. Znamenskaya and S. V. Znamenskii. Russian encoding plurality problem and a new Cyrillic font set. TUGboat, 17(2):161–165, June 1996. ISSN 0896-3207.